### v1

# k-LLMmeans: Summaries as Centroids for Interpretable and Scalable LLM-Based Text Clustering

Abstract
We introduce k-LLMmeans, a novel modification of the k-means clustering algorithm that utilizes LLMs to generate textual summaries as cluster centroids, thereby capturing contextual and semantic nuances often lost when relying on purely numerical means of document embeddings. This modification preserves the properties of k-means while offering greater interpretability: the cluster centroid is represented by an LLM-generated summary, whose embedding guides cluster assignments. We also propose a mini-batch variant, enabling efficient online clustering for streaming text data and providing real-time interpretability of evolving cluster centroids. Through extensive simulations, we show that our methods outperform vanilla k-means on multiple metrics while incurring only modest LLM usage that does not scale with dataset size. Finally, We present a case study showcasing the interpretability of evolving cluster centroids in sequential text streams. As part of our evaluation, we compile a new dataset from StackExchange, offering a benchmark for text-stream clustering.
k-LLMmeans: Summaries as Centroids for Interpretable and Scalable LLM-Based Text Clustering
Jairo Diaz-Rodriguez Department of Mathematics and Statistics York University Toronto, Ontario M3J 1P3 jdiazrod@yorku.ca
1 Introduction
Text clustering is a fundamental task in natural language processing (NLP), widely applied in document organization, topic modeling, and information retrieval (Schütze et al., 2008; Steinbach, 2000). A common approach involves generating text embeddings (Devlin, 2018; Sanh, 2019; Mikolov, 2013; Pennington et al., 2014; Brown et al., 2020) for each document, which are then clustered using traditional algorithms (Petukhova et al., 2025). Among these, k-means (MacQueen, 1967) is one of the most widely used, iteratively refining cluster centroids based on the mean of assigned points. Several alternative methods exist for computing centroids (Jain and Dubes, 1988; Bradley et al., 1996; Kaufman and Rousseeuw, 2008), yet they often rely on purely numerical calculations. While effective, such approaches risk losing crucial contextual and semantic information, as averaging high-dimensional embeddings can obscure nuanced textual relationships (Reimers and Gurevych, 2019).
In this work, we propose a modification to the k-means algorithm by leveraging large language models (LLMs) to dynamically generate cluster centroids. We call this the k-LLMmeans. Instead of always computing centroids as the mean of embedding vectors, in spaced iterations we generate centroids based on textual summaries of the cluster contents. At each of these iterations, the centroid is represented as the embedding of the LLM-generated summary of the texts belonging to the cluster. Building on Jia and Diaz-Rodriguez (2025), the LLM-based centroids preserve key contextual and semantic aspects of the main documents in the cluster while filtering out secondary or lower-priority documents. This summary-based centroid provides a more interpretable and contextually relevant representation, mitigating the loss of meaning inherent in numerical averaging.
One key advantage of our approach is the interpretability of the clustering process. Traditional k-means lacks explicit explanations for how clusters evolve over iterations, whereas our method provides textual summaries that offer insights into the semantic shifts of cluster centroids. This can be practically beneficial for debugging and validating the expected behavior of the algorithm. Moreover, this transparency in current status of the clustering becomes a key advantage on sequential clustering where clusters might evolve on time and interpretability becomes important. In this area we provide a modified version of the mini-batch k-Means (Sculley, 2010) by utilizing our k-LLMmeans sequentially.
Several studies have explored unsupervised text clustering using LLMs (Zhang et al., 2023; Feng et al., 2024; De Raedt et al., 2023; Viswanathan et al., 2024; Shi and Sakai, 2023; Tarekegn et al., 2024; Nakshatri et al., 2023), demonstrating state-of-the-art performance across various datasets and benchmarks. These approaches consistently outperform traditional clustering algorithms such as k-means. However, a key limitation is their complexity and their reliance on fine-tuning or iterative querying of LLMs that scales with the dataset. While this improves clustering quality, it introduces instability, might requires extensive parameter/prompt tuning, and limits scalability for big data.
Our approach is not designed to surpass complex state-of-the-art LLM-based text clustering methods but rather to provide a scalable and transparent LLM-enhanced modification to the well-established k-means algorithm, maintaining its key behavior. Fundamentally, a centroid is a numerical abstraction that represents a cluster, and we posit that an LLM-generated textual summary can serve an analogous role. Crucially, our approach preserves most of the mathematical properties of k-means, ensuring that core theoretical guarantees—such as convergence behavior, cluster compactness, and complexity— remain intact. Unlike alternative strategies that introduce complex pre-processing or post-processing steps, often altering the underlying mathematical framework, our method seamlessly integrates LLMs while maintaining k-means’ well-defined optimization landscape. This balance allows for enhanced interpretability and adaptability while preserving the efficiency that have made k-means a cornerstone of clustering algorithms.
In summary our contributions are as follows:
  • We introduce k-LLMmeans (Section 4), that leverages LLMs to enhance centroid estimation in k-means for text clustering.
  • We propose mini-batch k-LLMmeans (Section 5), designed for sequential, scalable, and interpretable text-stream clustering.
  • Through extensive simulations (Section 6), we demonstrate that both methods outperform k-means while maintaining low LLM usage that does not scale with the dataset size.
  • We present a case study (Section 7) demonstrating the interpretability of evolving cluster centroids in sequential text streams.
As a by-product of our experiments, we compile a dataset from StackExchange (2024) suitable for benchmarking text-streaming analysis methods (Section 6.2).
2 Related work
Clustering techniques are central to natural language processing and machine learning. Hierarchical methods (Johnson, 1967; Blashfield and Aldenderfer, 1978) build tree-structured representations of nested document relationships. Density-based approaches like DBSCAN (Ester et al., 1996) and graph-based methods detect clusters of arbitrary shapes, while spectral clustering (Ng et al., 2001) leverages eigen-decomposition to uncover complex structures. Model-based techniques—including Gaussian mixture models (Dempster et al., 1977) and recent neural network frameworks (Zhou et al., 2019; Huang et al., 2014; Yang et al., 2016; Zhang et al., 2021; Xie et al., 2016)—provide probabilistic clustering formulations. Additionally, topic modeling methods, from probabilistic latent semantic analysis (Hofmann, 2001) to latent Dirichlet allocation (Blei et al., 2003), capture word co-occurrence patterns and latent topics. However, these approaches diverge from our objective of enhancing k-means with LLMs, and are therefore not directly comparable.
Recent studies have integrated LLMs into clustering pipelines to reduce expert supervision and enhance performance. For instance, Viswanathan et al. (2024) employ LLMs to augment document representations, generate pseudo pairwise constraints, and post-correct low-confidence assignments for query-efficient, few-shot semi-supervised clustering. Similarly, Zhang et al. (2023) propose ClusterLLM, which uses instruction-tuned LLMs via interactive triplet and pairwise feedback to cost-effectively refine clustering granularity. Complementary approaches (Tipirneni et al., 2024; Petukhova et al., 2025) show that context-derived representations capture subtle semantic nuances beyond traditional embeddings. Additionally, Wang et al. (2023) introduce a goal-driven, explainable clustering method that employs natural language descriptions to clarify cluster boundaries, while De Raedt et al. (2023) present IDAS for intent discovery using abstractive summarization. Moreover, Feng et al. (2024) propose an iterative refinement mechanism that forms super-points to mitigate outliers and reassign ambiguous edge points, resulting in clusters with higher coherence and robustness. In contrast, our approach directly enhances the core k-means algorithm in an LLM-scalable manner.
input : $\ D=\{d_{1},\dots,d_{n}\},k,I,m,l,T$
for $i\leftarrow 1$ to $n$ do
       $\mathbf{x}_{i}=\text{Embedding}(d_{i})$;
end for
for $t\leftarrow 1$ to $T$ do
       if $t=1$ then
             // Initialize centroids using k-means++
             // This step can be omitted if initial centroids are provided
             $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\texttt{k-means++}(\{d_{1},\dots,d_{n}\},k)$;
       end if
      else if $t\bmod l=0$ then
             // Update centroids with LLM every $l$ iterations
             for $j\leftarrow 1$ to $k$ do
                   $m_{j}\leftarrow\min(m,|C_{j}|)$;
                  $\{d_{z_{1}},\dots,d_{z_{m_{j}}}\}\leftarrow\texttt{k-means++}(\{d_{i}\mid i\in[C_{j}]\},m_{j})$;
                  $p_{j}\leftarrow\text{Prompt}(I,\{d_{z_{1}},\dots,d_{z_{m_{j}}}\})$;
                   $s_{j}\leftarrow f_{\text{LLM}}(p_{j})$;
                   $\boldsymbol{\mu}_{j}\leftarrow\text{Embedding}(s_{j})$;
             end for
       end if
      else
             // Update centroids using standard averaging
             for $j\leftarrow 1$ to $k$ do
                   $\boldsymbol{\mu}_{j}\leftarrow\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}$;
             end for
       end if
      for $j\leftarrow 1$ to $k$ do
             $C_{j}=\{\}$;
       end for
      for $i\leftarrow 1$ to $n$ do
             $j^{*}\leftarrow\arg\min_{j\in\{1,\dots,k\}}d(x_{i},\boldsymbol{\mu}_{j})$;
             // Assign $x_{i}$ to cluster $C_{j^{*}}$
             $C_{j^{*}}\leftarrow C_{j^{*}}\cup\{x_{i}\}$;
       end for
end for
return $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\},\{s_{1},\dots,s_{k}\}$
Algorithm 1 k-LLMmeans
3 Preliminaries: k-Means for text clustering
We can start formalizing the text clustering problem. Given a corpus of $n$ text documents $D=\{d_{1},\cdots d_{n}\}$. Each document $d_{i}$ is represented as a $d$-dimensional embedding vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$ such that:
$\mathbf{x}_{i}=\text{Embedding}(d_{i}).$
We assume that all embeddings are normalized such that $\|\mathbf{x}_{i}\|=1$, equivalently using the euclidian distance or cosine similarity as the distance metric between embeddings. The goal of K-Means clustering is to partition these $n$ document embeddings into $k$ clusters, minimizing the intra-cluster variance. Formally, we define the clustering objective as:
$\arg\min_{C_{1},C_{2},\dots,C_{k}}\sum_{j=1}^{k}\sum_{i\in[C_{j}]}\|\mathbf{x}_{i}-\boldsymbol{\mu}_{j}\|^{2},$ (1)
where $C_{j}$ denotes the set of embeddings assigned to cluster $j$, $[C_{j}]=\{i|\mathbf{x}_{i}\in C_{j}\}$ denotes the set of embedding indices assigned to cluster $j$ and $\boldsymbol{\mu}_{j}$ is the cluster centroid, computed as the mean of the assigned embeddings:
$\boldsymbol{\mu}_{j}=\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}.$ (2)
The k-means algorithm assigns each document embedding $\mathbf{x}_{i}$ to the closest centroid based on the smallest distance and updates centroids accordingly until convergence after $T$ iterations. The primary objective of the k-means algorithm is to iteratively adjust cluster centroids to minimize the within-cluster variance, effectively guiding the algorithm toward an optimal set of centroids. However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima (MacQueen, 1967; Lloyd, 1982). Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results (Arthur and Vassilvitskii, 2006).
4 k-LLMmeans
To enhance k-means for text clustering, we introduce k-LLMmeans, a novel variant that integrates LLM-based centroids at intermediate stages of the clustering process. The formal procedure is outlined in Algorithm 1. The key distinction between k-LLMmeans and standard k-means lies in the centroid update mechanism: every $l$ iterations, the traditional update step in Equation (2) is replaced by
$\boldsymbol{\mu}_{j}=\text{Embedding}(f_{\text{LLM}}(p_{j}))\$ (3)
where $p_{j}=\text{Prompt}\left(I,\{d_{z_{i}}|z_{i}\sim[C_{j}]\}_{i=1}^{m_{j}}\right)$, and $m_{j}=\min(m,|C_{j}|)$. Here, $z_{i}\sim[C_{j}]$ denotes a sampled index of the embeddings assigned to cluster $C_{j}$ (without repetitions) and $m$ is a parameter that represents the maximum number of sampled indices used to compute the cluster centroid $\boldsymbol{\mu}_{j}$. In simple terms, we update a cluster’s centroid by using the embedding of the response generated by an LLM when queried with a prompt containing a summarization instruction $I$ and a representative sample of documents from the cluster. Rather than providing all documents within the cluster as input, the LLM processes a representative sample as a context prompt. While incorporating the entire cluster is theoretically possible, it poses practical challenges due to prompt length limitations. Therefore, we propose selecting the sample cluster documents using a k-means++ sampling of the cluster embeddings. Our experiments demonstrate that this sampling process facilitates a more effective synthesis of the cluster’s content, leading to improved summaries and, consequently, more refined centroid updates. The instruction $I$ varies depending on the clustering task, but standard summarization prompts are generally sufficient.
This novel approach can help mitigate the tendency of k-means to get stuck in local optima by dynamically adjusting centroids using semantic insights from the data. Unlike standard k-means, which relies solely on Euclidean updates, an LLM can refine centroids based on contextual meaning and high-dimensional representations, allowing for better adaptation to complex structures in text data. By periodically re-centering clusters with LLM-informed embeddings, the algorithm can escape poor local minima and achieve more coherent and semantically meaningful clusters, even when the initial k-means++ seeding is suboptimal. Apart from this modification, our algorithm adheres to the core principles of k-means, preserving its well-established properties and ensuring practical robustness. We demonstrate in Section 6.4 that k-LLMmeans consistently outperforms k-means across extensive simulations.
Figure 1: An illustration of the dynamic evolution of an LLM-generated centroid during the k-LLMmeans clustering process.
4.1 Scalability and transparency
Our approach offers two key advantages over more complex LLM-based clustering methods: scalability and transparency. Unlike most state-of-the-art methods, whose LLM usage complexity grows with sample size (Feng et al., 2024), our algorithm depends only on $k$ and $l$, with even small $l$ values yielding performance gains over k-means. Additionally, our method enhances interpretability by producing LLM-updated centroids that meaningfully represent clusters, allowing practitioners to track and validate the algorithm evolution without requiring post-processing. While similar insights can be extracted indirectly in other clustering algorithms, our approach integrates this interpretability directly into the clustering process, improving both usability and practical applicability (See Figure 1 for an example of such evolution of summaries).
5 Mini-batch k-LLMmeans
input : $\ \{D_{1},\cdots,D_{b}\},k,I,m,l,T$
// $b$ batches of documents
for $j\leftarrow 1$ to $k$ do
       $C_{j}=\{\}$;
end for
$\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\{{\bf 0},\dots,{\bf 0}\}$;
for $i\leftarrow 1$ to $b$ do
// Compute k-LLMmeans with documents in batch11      1Here k-LLMmeans is initialized with the final centroids of the previous batch
       $\{\boldsymbol{\mu}^{*}_{1},\dots,\boldsymbol{\mu}^{*}_{k}\},\{C^{*}_{1},\dots,C^{*}_{k}\},S_{b}\leftarrow\texttt{k-LLMmeans}(D_{i},k,I,m,l,T)$;
       // Update centroids proportional to current cluster sizes and batch cluster sizes
       for $j\leftarrow 1$ to $k$ do
             $\eta\leftarrow\frac{|C^{*}_{j}|}{|C_{j}|+|C^{*}_{j}|}$;
             $\boldsymbol{\mu}_{j}\leftarrow\boldsymbol{\mu}_{j}(1-\eta)+\eta\boldsymbol{\mu}^{*}_{j}$;
       end for
end for
return $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\},\{S_{1},\dots,S_{b}\}$
Algorithm 2 Mini-batch k-LLMmeans
Mini-batch k-means (Sculley, 2010) is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset. This approach substantially reduces memory usage and computational cost, making it well suited for continuously generated text streams—such as those from social media, news, or customer feedback—where data must be clustered incrementally without full dataset access. Mini-batch k-means exhibits convergence properties comparable to standard k-means while offering superior scalability.
Although numerous streaming clustering methods that do not rely on LLMs have been studied (Silva et al., 2013; Aggarwal, 2018; Ribeiro et al., 2017; Aggarwal et al., 2003; Ackermann et al., 2012; Ordonez, 2003), only a few have incorporated LLMs (Tarekegn et al., 2024; Nakshatri et al., 2023). Moreover, existing offline LLM-based clustering approaches face scalability issues, highlighting the need for scalable LLM-driven clustering in an online setting. To address this gap, we introduce mini-batch k-LLMmeans, which directly extends mini-batch k-means by incorporating minimal LLM usage during centroid updates. Algorithm 2 details how mini-batch k-LLMmeans sequentially receives $b$ batches of documents $D_{1},\dots D_{b}$ where each batch contains a set of documents (these batches can either be random samples from a large corpus or represent sequential data). It processes each batch sequentially with k-LLMmeans and updates centroids incrementally using a weighted rule like mini-batch k-means. Our mini-batch k-LLMmeans algorithm preserves the desirable properties of mini-batch k-means, with low memory and LLM usage. Section  6 shows that it also outperforms in simulations.
6 Experiments
Dataset/Method Bank77 CLINC GoEmo Massive (I) Massive (D)
NMI dist NMI dist NMI dist NMI dist NMI dist
distilbert k-means 64.4 0.335 77.2 0.34 18.3 0.364 58.1 0.366 45.1 0.309
k-medoids 56.0 0.591 66.8 0.69 14.0 0.769 42.6 0.847 27.2 0.81
k-LLMmeans-1 64.8 0.351 77.4 0.352 18.3 0.358 58.6 0.362 46.0 0.294
k-LLMmeans-5 64.8 0.358 77.7 0.356 18.3 0.354 59.0 0.363 45.6 0.306
k-LLMmeans-FS1 64.8 0.352 78.1 0.346 18.9 0.357 58.8 0.365 45.7 0.299
k-LLMmeans-FS5 64.9 0.363 78.7 0.343 18.8 0.351 59.0 0.362 46.4 0.295
openai k-means 83.0 0.225 92.0 0.2 20.5 0.287 72.4 0.32 67.9 0.246
k-medoids 69.5 0.639 77.7 0.694 15.9 0.852 52.4 0.86 38.8 0.797
k-LLMmeans-1 83.6 0.221 92.5 0.194 22.3 0.278 73.0 0.314 69.6 0.232
k-LLMmeans-5 83.6 0.226 92.8 0.186 22.1 0.275 73.5 0.307 69.5 0.238
k-LLMmeans-FS1 83.8 0.219 92.8 0.186 21.9 0.283 73.6 0.314 69.4 0.233
k-LLMmeans-FS5 84.1 0.22 93.1 0.179 22.3 0.278 73.9 0.302 70.6 0.227
e5-large k-means 76.7 0.142 90.8 0.131 22.8 0.176 70.9 0.175 63.7 0.138
k-medoids 64.5 0.331 73.4 0.371 15.7 0.447 50.3 0.437 36.3 0.405
k-LLMmeans-1 78.6 0.138 91.6 0.127 23.4 0.176 72.0 0.17 64.8 0.134
k-LLMmeans-5 79.0 0.136 92.0 0.127 23.7 0.174 72.1 0.166 66.0 0.131
k-LLMmeans-FS1 78.8 0.136 91.7 0.127 23.7 0.174 71.6 0.173 64.9 0.136
k-LLMmeans-FS5 79.5 0.134 92.5 0.119 24.3 0.168 72.4 0.166 65.9 0.133
sbert k-means 80.9 0.255 91.0 0.215 13.3 0.355 70.7 0.344 64.6 0.271
k-medoids 69.0 0.64 78.5 0.68 12.0 0.899 54.6 0.817 44.6 0.8
k-LLMmeans-1 81.7 0.253 91.6 0.215 13.6 0.351 71.2 0.334 65.6 0.248
k-LLMmeans-5 81.8 0.253 91.9 0.21 13.7 0.348 71.5 0.325 65.2 0.25
k-LLMmeans-FS1 82.0 0.246 91.8 0.208 13.9 0.348 71.4 0.333 66.2 0.243
k-LLMmeans-FS5 82.2 0.247 92.5 0.198 13.9 0.346 71.9 0.337 65.6 0.254
Table 1: Average Normalized Mutual Information (NMI) and distance between final and true centroids (dist) for k-means, k-medoids, and four k-LLMmeans variants across 10 random seeds on four benchmark datasets (including both domain and intent from MASSIVE), using four different embedding models.
6.1 Datasets
We evaluate our clustering approach on four benchmark datasets:
  • Bank77 (Casanueva et al., 2020): Consists of 3,080 customer queries related to banking services, categorized into 77 distinct intents.
  • CLINC (Larson et al., 2019): A diverse set of 4,500 queries spanning 150 intent classes across multiple domains, designed for open-domain intent classification.
  • GoEmo (Demszky et al., 2020): Contains 2,984 social media posts annotated with 27 fine-grained emotion categories. We removed the neutral expressions to address data imbalance and retained only entries with a single, unique emotion.
  • MASSIVE (FitzGerald et al., 2023): Comprises 2,974 English-language virtual assistant utterances grouped into 18 domains and 59 intent categories.
These datasets provide a robust evaluation setting for text clustering across different domains and classification granularities.
6.2 New compiled dataset for testing text-Streaming Clustering Algorithms
We extract and unify a challenging data stream comprising unique archive posts collected from 84 Stack Exchange sites (StackExchange, 2024). Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods. Our raw dataset spans 84 domains, each containing at least 20 posts per year from 2018 to 2023 (with post lengths ranging from 20 to 1000 characters), totaling 499,359 posts. For our experiments, we focus on posts from 2020 to 2023 and further filter out labels that do not exceed 500 posts in 2023. The resulting subset comprises 35 distinct groups and 69,147 posts. Both the raw and clean data are provided with this paper.
6.3 Methods
Year/Method 2020 2021 2022 2023
(69147 posts) (54322 posts) (43521 posts) (38953 posts)
NMI dist NMI dist NMI dist NMI dist
k-means 80.6 0.14 79.0 0.167 79.0 0.154 79.6 0.138
mini-batch k-means 78.2 0.181 77.4 0.166 77.6 0.173 77.0 0.175
seq. mini-batch k-means 76.6 0.187 75.2 0.185 75.6 0.187 74.8 0.184
mini-batch k-LLMmeans-1 81.2 0.141 80.3 0.138 79.4 0.145 80.0 0.136
mini-batch k-LLMmeans-5 80.9 0.144 80.5 0.127 80.5 0.125 80.3 0.129
mini-batch k-LLMmeans-FS1 81.1 0.136 79.8 0.141 79.3 0.147 79.8 0.138
mini-batch k-LLMmeans-FS5 81.6 0.126 80.2 0.126 80.1 0.129 80.1 0.133
Table 2: Average Normalized Mutual Information (NMI) and distance from final and true centroids (dist) for k-means, mini-batch k-means, sequential mini-batch k-means and four sequential mini-batch k-LLMmeans variants with across 5 random seeds on StackExchange data.
We evaluate our k-LLMmeans algorithm on each of the four static (non-streaming) dataset using the known number of clusters and performing 120 centroid-update iterations and 10 different seeds. To demonstrate the robustness of our approach, we compute embeddings with four different pretrained models: distilbert (Sanh, 2019), e5-large (Wang et al., 2022), s-bert (Reimers and Gurevych, 2019), and OpenAI’s text-embedding-3-small (OpenAI, 2023). For the LLM component, we only use OpenAI’s gpt-4o to ensure any observed differences in performance arise from the effectiveness of our clustering method rather than the inherent strengths or weaknesses of different LLMs.
For the instruction task $I$, we employ a simple summarization prompt that adapted to each dataset. For example, for Bank77 we use the prompt: “The following is a cluster of online banking questions. Write a single question that represents the cluster concisely.” We examine four variations of k-LLMmeans based on different numbers of summarization steps and the size of the prompts:
  • k-LLMmeans-1: A single summarization step ($l=60$) using all documents in the cluster as input.
  • k-LLMmeans-5: Five summarization steps ($l=20$) with all documents in the cluster as input at each step.
  • k-LLMmeans-FS1: A few-shot variant with 10 randomly selected documents in a single summarization step.
  • k-LLMmeans-FS5: A few-shot variant with 10 randomly selected documents in five summarization steps.
Our baselines are the standard centroid-based algorithms k-means and k-medoids. While alternative clustering methods may achieve stronger performance, our primary goal is to demonstrate improvement specifically over widely used centroid-based approaches.
Methods for streaming dataset. We partition the StackExchange data into four subsets, each corresponding to a single year from 2020 to 2023 (See Table 2 for yearly dataset sizes). For each yearly subset, we split the data into $b=\lceil\tfrac{n}{10000}\rceil$ equal-sized batches ${D_{1},\dots,D_{b}}$ in chronological order, where $n$ is the number of documents for that year. We then use the ground-truth clusters and run the mini-batch k-LLMmeans algorithm in its four previously described variants (for practical reasons for the full cluster method we use 50 samples). For comparison, we evaluate three baselines: mini-batch k-means with standard random sampling across the entire year (until convergence), sequential mini-batch k-means with $b$ chronological batches, and standard k-means on the full dataset. In all experiments, we use OpenAI’s models for both embeddings and the LLM.
6.4 Results
Table 1 reports the average Normalized Mutual Information (NMI) across all method–dataset–embedding combinations, as well as the average distance (dist) between the centroids produced by each method and the true centroids calculated with the ground truth clusters. A lower distance indicates greater proximity to the optimal solution. This distance aligns well with the objectives of our approach. We also measure Accuracy (not shown here due to space constraints), observing results consistent with the reported metrics. Overall, all k-LLMmeans variants outperform k-means except when using the distilbert embedding. However, the performance of every method using this embedding is clearly inferior compared to those employing any other embedding. Meanwhile, k-medoids emerges as the worst performer. Comparing the four k-LLMmeans configurations reveals only minor differences, although running five summarization steps ($l=20$) tends to offer better performance. Interestingly, using few-shot summarization generally leads to better results than prompting on the entire cluster. These observations suggest k-LLMmeans improves over k-means while remaining efficient, as few-shot summarization appears sufficient without needing extensive LLM usage.
We present the same evaluation metrics for sequential data in Table 2. The mini-batch k-LLMmeans methods consistently outperform all three baselines across all scenarios. Notably, they even surpass standard k-means, which operates on the full dataset rather than mini-batches. This highlights the scalability and efficiency of our approach for sequential data, achieving superior performance despite processing data sequentially. As observed previously, few-shot summarization proves effective, mitigating the need for extensive LLM usage.
7 Case study
Figure 2: Sequential evolution of the LLM-generated centroids for four primary clusters during the three batches of the sequential mini-batch k-LLMmeans process applied to 2021 posts from the AI Stack Exchange site (StackExchange, 2024). Main aspects are manually highlighted in different color on each cluster.
To demonstrate the interpretability of our method in capturing the evolution of clusters within sequential data, we present a case study using posts from the AI site in the 2021 Stack Exchange dataset (StackExchange, 2024). We apply our mini-batch k-LLMmeans algorithm with three equal-length batches and a total of ten clusters. We use the instruction “The following is a cluster of questions from the AI community. Write a single question that represents the cluster”.
Clustering results. The resulting LLM-based centroids span key areas in AI and ML, including neural network training optimization, computer vision tasks, and broader topics such as small datasets, class imbalance, and interpretability. Other clusters cover advanced themes like symbolic-neural integration for AGI, deep reinforcement learning, theoretical debates (e.g., Bayesian vs. frequentist methods), as well as NLP and various architectural choices.
Interpretation. The key insight from our algorithm is the evolution of clusters over batches, reflecting their dynamic nature over time. Figure 2 illustrates this progression for four major themes: Image Model Optimization, AI Evolution and Challenges, AI Mathematics, and Advanced NLP Techniques. The Image Model Optimization cluster refines its focus from broad improvements in image classification and object detection to specific deep learning optimizations, such as facial expression recognition and anomaly detection. Over time, it emphasizes practical challenges, including handling imbalanced datasets, model uncertainty, and performance enhancement through transfer learning and specialized loss functions. The AI Evolution and Challenges cluster transitions from early rule-based AI and symbolic systems to advanced deep learning and generative models, highlighting both progress and limitations. The shift toward hybrid and neuro-symbolic AI reflects the growing need to integrate diverse techniques for AGI, balancing efficiency, adaptability, and safety. The AI Mathematics cluster starts with unconventional symbol usage and parameter definitions in machine learning, later expanding to matrix rank analysis, Bayesian and frequentist methods, adversarial attacks, and the mathematical foundations necessary for advanced AI concepts. The Advanced NLP Techniques cluster progresses from foundational models (e.g., Word2Vec, BERT) and basic tasks (e.g., sentiment analysis, text classification) to more complex challenges, including model evaluation, domain-specific vocabulary, and multilingual alignment. The shift highlights an increasing focus on contextual understanding, transfer learning, and alternative evaluation metrics for robust NLP.
The result of this case study could enhance post categorization, searchability, answer relevance, and trend detection, making AI discussions more efficient and insightful. More broadly, this demonstrates how our interpretable mini-batch k-LLMmeans clustering algorithm, applied to sequential text streaming, can help practitioners track topic evolution, improve decision-making, and enhance transparency in dynamic information flows.
8 Conclusions
We introduced k-LLMmeans and mini-batch k-LLMmeans, unsupervised clustering algorithms that refine k-means by integrating a lightweight LLM-based summarization into centroid updates. This modification enriches the contextual representation while preserving k-means’ efficiency and scalability, yielding interpretable clusters for both static corpus of documents and streaming text data.
Limitations
Our method relies on both text embeddings and LLM queries, making it sensitive to the quality and biases of the underlying language models. Any inaccuracies in these models may propagate into the clustering results. While we demonstrate that a simple instruction-based approach is effective, the design of these instructions can still influence outcomes. Additionally, the few-shot variant of our algorithm assumes that a small number of samples can sufficiently capture the overall structure of the clusters. While this is generally practical, it may become a limitation when dealing with highly complex or heterogeneous clusters. Finally, similar to k-means, our k-LLMmeans method requires specifying the number of clusters in advance, which may not always align with the true structure of the data.

### v2

@@
-k-LLMmeans: Summaries as Centroids for Interpretable and Scalable LLM-Based Text Clustering
+k-LLMmeans: Scalable, Stable, and Interpretable Text Clustering via LLM-based Centroids
 Abstract
-We introduce k-LLMmeans, a novel modification of the k-means clustering algorithm that utilizes LLMs to generate textual summaries as cluster centroids, thereby capturing contextual and semantic nuances often lost when relying on purely numerical means of document embeddings.
-This modification preserves the properties of k-means while offering greater interpretability: the cluster centroid is represented by an LLM-generated summary, whose embedding guides cluster assignments.
-We also propose a mini-batch variant, enabling efficient online clustering for streaming text data and providing real-time interpretability of evolving cluster centroids.
-Through extensive simulations, we show that our methods outperform vanilla k-means on multiple metrics while incurring only modest LLM usage that does not scale with dataset size.
-Finally, We present a case study showcasing the interpretability of evolving cluster centroids in sequential text streams.
-As part of our evaluation, we compile a new dataset from StackExchange, offering a benchmark for text-stream clustering.
-k-LLMmeans: Summaries as Centroids for Interpretable and Scalable LLM-Based Text Clustering
-Jairo Diaz-Rodriguez Department of Mathematics and Statistics York University Toronto, Ontario M3J 1P3 jdiazrod@yorku.ca
+We introduce k-LLMmeans, a novel modification of the k-means algorithm for text clustering that leverages LLM-generated summaries as cluster centroids, capturing semantic nuances often missed by purely numerical averages.
+This design preserves the core optimization properties of k-means while enhancing semantic interpretability and avoiding the scalability and instability issues typical of modern LLM-based clustering.
+Unlike existing methods, our approach does not increase LLM usage with dataset size and produces transparent intermediate outputs.
+We further extend it with a mini-batch variant for efficient, real-time clustering of streaming text.
+Extensive experiments across multiple datasets, embeddings, and LLMs show that k-LLMmeans consistently outperforms k-means and other traditional baselines and achieves results comparable to state-of-the-art LLM-based clustering—with a fraction of the LLM calls.
+Finally, we present a case study on sequential text streams and introduce a new benchmark dataset constructed from StackExchange to evaluate text-stream clustering methods.
 1 Introduction
-Text clustering is a fundamental task in natural language processing (NLP), widely applied in document organization, topic modeling, and information retrieval (Schütze et al., 2008; Steinbach, 2000).
-A common approach involves generating text embeddings (Devlin, 2018; Sanh, 2019; Mikolov, 2013; Pennington et al., 2014; Brown et al., 2020) for each document, which are then clustered using traditional algorithms (Petukhova et al., 2025).
-Among these, k-means (MacQueen, 1967) is one of the most widely used, iteratively refining cluster centroids based on the mean of assigned points.
-Several alternative methods exist for computing centroids (Jain and Dubes, 1988; Bradley et al., 1996; Kaufman and Rousseeuw, 2008), yet they often rely on purely numerical calculations.
-While effective, such approaches risk losing crucial contextual and semantic information, as averaging high-dimensional embeddings can obscure nuanced textual relationships (Reimers and Gurevych, 2019).
-In this work, we propose a modification to the k-means algorithm by leveraging large language models (LLMs) to dynamically generate cluster centroids.
-We call this the k-LLMmeans.
-Instead of always computing centroids as the mean of embedding vectors, in spaced iterations we generate centroids based on textual summaries of the cluster contents.
-At each of these iterations, the centroid is represented as the embedding of the LLM-generated summary of the texts belonging to the cluster.
-Building on Jia and Diaz-Rodriguez (2025), the LLM-based centroids preserve key contextual and semantic aspects of the main documents in the cluster while filtering out secondary or lower-priority documents.
-This summary-based centroid provides a more interpretable and contextually relevant representation, mitigating the loss of meaning inherent in numerical averaging.
-One key advantage of our approach is the interpretability of the clustering process.
-Traditional k-means lacks explicit explanations for how clusters evolve over iterations, whereas our method provides textual summaries that offer insights into the semantic shifts of cluster centroids.
-This can be practically beneficial for debugging and validating the expected behavior of the algorithm.
-Moreover, this transparency in current status of the clustering becomes a key advantage on sequential clustering where clusters might evolve on time and interpretability becomes important.
-In this area we provide a modified version of the mini-batch k-Means (Sculley, 2010) by utilizing our k-LLMmeans sequentially.
-Several studies have explored unsupervised text clustering using LLMs (Zhang et al., 2023; Feng et al., 2024; De Raedt et al., 2023; Viswanathan et al., 2024; Shi and Sakai, 2023; Tarekegn et al., 2024; Nakshatri et al., 2023), demonstrating state-of-the-art performance across various datasets and benchmarks.
-These approaches consistently outperform traditional clustering algorithms such as k-means.
-However, a key limitation is their complexity and their reliance on fine-tuning or iterative querying of LLMs that scales with the dataset.
-While this improves clustering quality, it introduces instability, might requires extensive parameter/prompt tuning, and limits scalability for big data.
-Our approach is not designed to surpass complex state-of-the-art LLM-based text clustering methods but rather to provide a scalable and transparent LLM-enhanced modification to the well-established k-means algorithm, maintaining its key behavior.
+Text clustering is a fundamental task in natural language processing (NLP), widely applied in document organization, topic modeling, and information retrieval [37, 42].
+A common approach involves generating text embeddings [13, 36, 27, 32, 8, 21] for each document, which are then clustered using traditional algorithms [33].
+Among these, k-means [26] is one of the most widely used, iteratively refining cluster centroids based on the mean of assigned points.
+Several alternative methods exist for computing centroids [19, 7, 23], yet they often rely on purely numerical calculations.
+While effective, such approaches risk losing crucial contextual and semantic information, as averaging high-dimensional embeddings can obscure nuanced textual relationships [34].
+We propose k-LLMmeans, a modification to the k-means algorithm by leveraging large language models (LLMs) to dynamically generate cluster centroids.
+Instead of always computing centroids as the mean of embedding vectors (k-means), in spaced iterations we generate centroids based on textual summaries of the cluster contents.
+At each of these iterations, the centroid is represented as the embedding of the LLM-generated summary of the texts belonging to the cluster (summarization step).
+LLM-based centroids preserve key contextual and semantic aspects of the main documents in the cluster while filtering out secondary or lower-priority documents [20].
+Figure 1 illustrates the k-LLMmeans algorithm with a single summarization step.
+The example demonstrates how this step enables the algorithm to escape local minima (common in k-means) and converge to the optimal solution.
+Unlike k-means, k-LLMmeans generates a concise textual summary for every centroid at each summarization step, exposing the semantic drift of clusters in real time.
+This built-in transparency simplifies debugging and validation and is especially valuable in sequential or streaming scenarios, where clusters evolve over time.
+Figure 2 illustrates one centroid through five summarization updates.
+To capitalize on this property for large-scale or online data, we adapt mini-batch k-means [38] by inserting our summarization step into a mini-batch k-LLMmeans.
+Several studies have explored unsupervised text clustering using LLMs [51, 15, 10, 45, 39, 43, 28], demonstrating state-of-the-art performance across various datasets and benchmarks.
+These approaches consistently outperform traditional clustering algorithms.
+Yet despite their impressive empirical gains, existing LLM-based clustering pipelines still face two practical obstacles that limit widespread adoption.
+First, lack of scalability.
+Most methods rely on fine-tuning [51], semi-supervised approaches [45], or extensive LLM queries that grow at least linearly with sample size [15, 10], quickly becoming prohibitive for real-world applications.
+Second, unclear optimization landscape.
+Most methods braid together bespoke heuristics—hard prompts, similarity thresholds, or greedy merges—without a single, well-defined loss function, leaving convergence behaviour opaque and theoretical guarantees out of reach.
+We address these two shortcomings with k-LLMmeans: First, by summarizing each cluster in batches we cap LLM usage independent of dataset size; and second, by inheriting the classic k-means objective we retain a provable descent property and a transparent optimization landscape.
+Regardless of summarization quality, the algorithm will converge to a local minimum.
+Our approach is not intended to outperform complex state-of-the-art LLM-based clustering methods, but rather to provide a robust, scalable and LLM-efficient enhancement to the well-established k-means algorithm, preserving its core behavior while using significantly fewer LLM calls than contemporary LLM-based clustering approaches.
 Fundamentally, a centroid is a numerical abstraction that represents a cluster, and we posit that an LLM-generated textual summary can serve an analogous role.
 Crucially, our approach preserves most of the mathematical properties of k-means, ensuring that core theoretical guarantees—such as convergence behavior, cluster compactness, and complexity— remain intact.
-Unlike alternative strategies that introduce complex pre-processing or post-processing steps, often altering the underlying mathematical framework, our method seamlessly integrates LLMs while maintaining k-means’ well-defined optimization landscape.
-This balance allows for enhanced interpretability and adaptability while preserving the efficiency that have made k-means a cornerstone of clustering algorithms.
+Figure 1: Illustration of a single-summarization k-LLMmeans algorithm.
+The first panel displays the text embeddings projected onto the first two principal components along with sample texts, with stars indicating initial random centroids.
+The second panel shows the local minimum reached by standard k-means.
+The third panel illustrates the summarization step, where centroids are updated based on LLM-generated summaries of the clustered texts.
+The final panel presents the resulting clusters after the final k-means step, showing convergence to a more optimal k-LLMmeans solution.
 In summary our contributions are as follows:
-  • We introduce k-LLMmeans (Section 4), that leverages LLMs to enhance centroid estimation in k-means for text clustering.
-  • We propose mini-batch k-LLMmeans (Section 5), designed for sequential, scalable, and interpretable text-stream clustering.
-  • Through extensive simulations (Section 6), we demonstrate that both methods outperform k-means while maintaining low LLM usage that does not scale with the dataset size.
-  • We present a case study (Section 7) demonstrating the interpretability of evolving cluster centroids in sequential text streams.
-As a by-product of our experiments, we compile a dataset from StackExchange (2024) suitable for benchmarking text-streaming analysis methods (Section 6.2).
-2 Related work
-Clustering techniques are central to natural language processing and machine learning.
-Hierarchical methods (Johnson, 1967; Blashfield and Aldenderfer, 1978) build tree-structured representations of nested document relationships.
-Density-based approaches like DBSCAN (Ester et al., 1996) and graph-based methods detect clusters of arbitrary shapes, while spectral clustering (Ng et al., 2001) leverages eigen-decomposition to uncover complex structures.
-Model-based techniques—including Gaussian mixture models (Dempster et al., 1977) and recent neural network frameworks (Zhou et al., 2019; Huang et al., 2014; Yang et al., 2016; Zhang et al., 2021; Xie et al., 2016)—provide probabilistic clustering formulations.
-Additionally, topic modeling methods, from probabilistic latent semantic analysis (Hofmann, 2001) to latent Dirichlet allocation (Blei et al., 2003), capture word co-occurrence patterns and latent topics.
-However, these approaches diverge from our objective of enhancing k-means with LLMs, and are therefore not directly comparable.
-Recent studies have integrated LLMs into clustering pipelines to reduce expert supervision and enhance performance.
-For instance, Viswanathan et al.
-(2024) employ LLMs to augment document representations, generate pseudo pairwise constraints, and post-correct low-confidence assignments for query-efficient, few-shot semi-supervised clustering.
-Similarly, Zhang et al.
-(2023) propose ClusterLLM, which uses instruction-tuned LLMs via interactive triplet and pairwise feedback to cost-effectively refine clustering granularity.
-Complementary approaches (Tipirneni et al., 2024; Petukhova et al., 2025) show that context-derived representations capture subtle semantic nuances beyond traditional embeddings.
-Additionally, Wang et al.
-(2023) introduce a goal-driven, explainable clustering method that employs natural language descriptions to clarify cluster boundaries, while De Raedt et al.
-(2023) present IDAS for intent discovery using abstractive summarization.
-Moreover, Feng et al.
-(2024) propose an iterative refinement mechanism that forms super-points to mitigate outliers and reassign ambiguous edge points, resulting in clusters with higher coherence and robustness.
-In contrast, our approach directly enhances the core k-means algorithm in an LLM-scalable manner.
+  • In Section 3 we introduce k-LLMmeans, which enhances the classical k-means algorithm with LLM-powered centroid updates for robust text clustering; Section 4 then builds on this foundation with a mini-batch variant that brings sequential, scalable, and intrinsically interpretable clustering to streaming text data.
+  • Section 5 presents extensive simulations demonstrating that k-LLMmeans consistently surpasses k-means and other traditional clustering baselines across a wide range of embeddings and LLMs, while obtaining comparable results to recent LLM-based clustering methods with substantially fewer LLM calls.
+  • Section 6 shows that mini-batch k-LLMmeans decisively outperforms streaming variants of k-means, and the case study in Section 7 illustrates how its evolving centroids provide clear, human-readable insight into real-time text streams.
+As a by-product of our experiments, we compile a dataset from StackExchange [41] suitable for benchmarking text-streaming analysis methods (Appendix A.2).
+Figure 2: An illustration of the dynamic evolution of an LLM-generated centroid during the k-LLMmeans clustering process.
+2 Preliminaries: k-means for text clustering
+We can formalizae the text clustering problem.
+Given a corpus of $n$ text documents $D=\{d_{1},\cdots d_{n}\}$.
+Each document $d_{i}$ is represented as a $d$-dimensional embedding vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$ such that:
+$\mathbf{x}_{i}=\text{Embedding}(d_{i}).$
+The goal of k-means clustering is to partition these $n$ document embeddings into $k$ clusters, minimizing the intra-cluster variance. Formally, we define the clustering objective as:
+$\min_{C_{1},C_{2},\dots,C_{k}}\sum_{j=1}^{k}\sum_{i\in[C_{j}]}\|\mathbf{x}_{i}-\boldsymbol{\mu}_{j}\|^{2},$ (1)
+where $C_{j}$ denotes the set of embeddings assigned to cluster $j$, $[C_{j}]=\{i|\mathbf{x}_{i}\in C_{j}\}$ denotes the set of embedding indices assigned to cluster $j$ and $\boldsymbol{\mu}_{j}$ is the cluster centroid, computed as the mean of the assigned embeddings:
+$\boldsymbol{\mu}_{j}=\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}.$ (2)
+The k-means algorithm assigns each document embedding $\mathbf{x}_{i}$ to the closest centroid based on the smallest distance and updates centroids accordingly until convergence after $T$ iterations.
+The primary objective of the k-means algorithm is to iteratively adjust cluster centroids to minimize the within-cluster variance, effectively guiding the algorithm toward an optimal set of centroids.
+However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima [26, 25].
+Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results [4].
+3 k-LLMmeans
 input : $\ D=\{d_{1},\dots,d_{n}\},k,I,m,l,T$
@@
        if $t=1$ then
-             // Initialize centroids using k-means++
-             // This step can be omitted if initial centroids are provided
+             // Initialize using k-means++ (omitted if already initialized)
              $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\texttt{k-means++}(\{d_{1},\dots,d_{n}\},k)$;
@@
       else if $t\bmod l=0$ then
-             // Update centroids with LLM every $l$ iterations
+             // Update with LLM every $l$ iterations (summarization step)
              for $j\leftarrow 1$ to $k$ do
@@
       else
-             // Update centroids using standard averaging
+             // Update centroids using standard averaging (k-means step)
              for $j\leftarrow 1$ to $k$ do
@@
 Algorithm 1 k-LLMmeans
-3 Preliminaries: k-Means for text clustering
-We can start formalizing the text clustering problem.
-Given a corpus of $n$ text documents $D=\{d_{1},\cdots d_{n}\}$.
-Each document $d_{i}$ is represented as a $d$-dimensional embedding vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$ such that:
-$\mathbf{x}_{i}=\text{Embedding}(d_{i}).$
-We assume that all embeddings are normalized such that $\|\mathbf{x}_{i}\|=1$, equivalently using the euclidian distance or cosine similarity as the distance metric between embeddings.
-The goal of K-Means clustering is to partition these $n$ document embeddings into $k$ clusters, minimizing the intra-cluster variance.
-Formally, we define the clustering objective as:
-$\arg\min_{C_{1},C_{2},\dots,C_{k}}\sum_{j=1}^{k}\sum_{i\in[C_{j}]}\|\mathbf{x}_{i}-\boldsymbol{\mu}_{j}\|^{2},$ (1)
-where $C_{j}$ denotes the set of embeddings assigned to cluster $j$, $[C_{j}]=\{i|\mathbf{x}_{i}\in C_{j}\}$ denotes the set of embedding indices assigned to cluster $j$ and $\boldsymbol{\mu}_{j}$ is the cluster centroid, computed as the mean of the assigned embeddings:
-$\boldsymbol{\mu}_{j}=\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}.$ (2)
-The k-means algorithm assigns each document embedding $\mathbf{x}_{i}$ to the closest centroid based on the smallest distance and updates centroids accordingly until convergence after $T$ iterations.
-The primary objective of the k-means algorithm is to iteratively adjust cluster centroids to minimize the within-cluster variance, effectively guiding the algorithm toward an optimal set of centroids.
-However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima (MacQueen, 1967; Lloyd, 1982).
-Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results (Arthur and Vassilvitskii, 2006).
-4 k-LLMmeans
 To enhance k-means for text clustering, we introduce k-LLMmeans, a novel variant that integrates LLM-based centroids at intermediate stages of the clustering process.
@@
 The instruction $I$ varies depending on the clustering task, but standard summarization prompts are generally sufficient.
+Figure 1 illustrates how k-LLMmeans augments the standard k-means algorithm by introducing few summarization steps between clustering iterations.
+Advantages over k-means.
 This novel approach can help mitigate the tendency of k-means to get stuck in local optima by dynamically adjusting centroids using semantic insights from the data.
@@
 We demonstrate in Section 5.1 that k-LLMmeans consistently outperforms k-means across extensive simulations.
-Figure 1: An illustration of the dynamic evolution of an LLM-generated centroid during the k-LLMmeans clustering process.
-4.1 Scalability and transparency
-Our approach offers two key advantages over more complex LLM-based clustering methods: scalability and transparency.
-Unlike most state-of-the-art methods, whose LLM usage complexity grows with sample size (Feng et al., 2024), our algorithm depends only on $k$ and $l$, with even small $l$ values yielding performance gains over k-means.
-Additionally, our method enhances interpretability by producing LLM-updated centroids that meaningfully represent clusters, allowing practitioners to track and validate the algorithm evolution without requiring post-processing.
-While similar insights can be extracted indirectly in other clustering algorithms, our approach integrates this interpretability directly into the clustering process, improving both usability and practical applicability (See Figure 1 for an example of such evolution of summaries).
+Advantages over advanced LLM-based clustering methods.
+Our approach offers three key advantages over more complex LLM-based clustering methods: (1) Optimization landscape.
+Grounded in the standard k-means objective, k-LLMmeans inherits provable convergence without relying on the fragile heuristics common in newer LLM-driven methods.
+A faulty summary merely returns the algorithm to a regular k-means local optimum, whereas competing approaches hinge on stable, format-specific LLM outputs.
+(2) Scalability.
+Unlike most state-of-the-art methods, whose LLM usage complexity grows with sample size [15, 10] or require fine-tuning [51], our algorithm only performs $O(kl)$ LLM calls, with even small $l$ values yielding substancial performance gains.
+(See Section 5.1).
+(3) Interpretability.
+By replacing numerical centroids with LLM-generated textual summaries, k-LLMmeans turns each centroid into a concise, human-readable synopsis of its cluster.
+Practitioners can therefore monitor how cluster semantics evolve across iterations without any post-hoc labeling or visualization.
+While comparable insights can be extracted indirectly from other clustering algorithms, k-LLMmeans embeds this interpretability natively.
+Figure 2 illustrates how these summaries progressively sharpen as the algorithm converges.
 4 Mini-batch k-LLMmeans
-input : $\ \{D_{1},\cdots,D_{b}\},k,I,m,l,T$
-// $b$ batches of documents
+input : $\ \{D_{1},\cdots,D_{b}\},k,I,m,l,T$ // $b$ batches of documents
 for $j\leftarrow 1$ to $k$ do
@@
        $\{\boldsymbol{\mu}^{*}_{1},\dots,\boldsymbol{\mu}^{*}_{k}\},\{C^{*}_{1},\dots,C^{*}_{k}\},S_{b}\leftarrow\texttt{k-LLMmeans}(D_{i},k,I,m,l,T)$;
-       // Update centroids proportional to current cluster sizes and batch cluster sizes
+       // Update centroids proportional to cluster and batch sizes
        for $j\leftarrow 1$ to $k$ do
@@
 Algorithm 2 Mini-batch k-LLMmeans
-Mini-batch k-means (Sculley, 2010) is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset.
+Mini-batch k-means [38] is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset.
 This approach substantially reduces memory usage and computational cost, making it well suited for continuously generated text streams—such as those from social media, news, or customer feedback—where data must be clustered incrementally without full dataset access.
 Mini-batch k-means exhibits convergence properties comparable to standard k-means while offering superior scalability.
-Although numerous streaming clustering methods that do not rely on LLMs have been studied (Silva et al., 2013; Aggarwal, 2018; Ribeiro et al., 2017; Aggarwal et al., 2003; Ackermann et al., 2012; Ordonez, 2003), only a few have incorporated LLMs (Tarekegn et al., 2024; Nakshatri et al., 2023).
+Although numerous streaming clustering methods that do not rely on LLMs have been studied [40, 2, 35, 3, 1, 31], only a few have incorporated LLMs [43, 28].
 Moreover, existing offline LLM-based clustering approaches face scalability issues, highlighting the need for scalable LLM-driven clustering in an online setting.
@@
 Section  6 shows that it also outperforms in simulations.
-6 Experiments
-Dataset/Method Bank77 CLINC GoEmo Massive (I) Massive (D)
-NMI dist NMI dist NMI dist NMI dist NMI dist
-distilbert k-means 64.4 0.335 77.2 0.34 18.3 0.364 58.1 0.366 45.1 0.309
-k-medoids 56.0 0.591 66.8 0.69 14.0 0.769 42.6 0.847 27.2 0.81
-k-LLMmeans-1 64.8 0.351 77.4 0.352 18.3 0.358 58.6 0.362 46.0 0.294
-k-LLMmeans-5 64.8 0.358 77.7 0.356 18.3 0.354 59.0 0.363 45.6 0.306
-k-LLMmeans-FS1 64.8 0.352 78.1 0.346 18.9 0.357 58.8 0.365 45.7 0.299
-k-LLMmeans-FS5 64.9 0.363 78.7 0.343 18.8 0.351 59.0 0.362 46.4 0.295
-openai k-means 83.0 0.225 92.0 0.2 20.5 0.287 72.4 0.32 67.9 0.246
-k-medoids 69.5 0.639 77.7 0.694 15.9 0.852 52.4 0.86 38.8 0.797
-k-LLMmeans-1 83.6 0.221 92.5 0.194 22.3 0.278 73.0 0.314 69.6 0.232
-k-LLMmeans-5 83.6 0.226 92.8 0.186 22.1 0.275 73.5 0.307 69.5 0.238
-k-LLMmeans-FS1 83.8 0.219 92.8 0.186 21.9 0.283 73.6 0.314 69.4 0.233
-k-LLMmeans-FS5 84.1 0.22 93.1 0.179 22.3 0.278 73.9 0.302 70.6 0.227
-e5-large k-means 76.7 0.142 90.8 0.131 22.8 0.176 70.9 0.175 63.7 0.138
-k-medoids 64.5 0.331 73.4 0.371 15.7 0.447 50.3 0.437 36.3 0.405
-k-LLMmeans-1 78.6 0.138 91.6 0.127 23.4 0.176 72.0 0.17 64.8 0.134
-k-LLMmeans-5 79.0 0.136 92.0 0.127 23.7 0.174 72.1 0.166 66.0 0.131
-k-LLMmeans-FS1 78.8 0.136 91.7 0.127 23.7 0.174 71.6 0.173 64.9 0.136
-k-LLMmeans-FS5 79.5 0.134 92.5 0.119 24.3 0.168 72.4 0.166 65.9 0.133
-sbert k-means 80.9 0.255 91.0 0.215 13.3 0.355 70.7 0.344 64.6 0.271
-k-medoids 69.0 0.64 78.5 0.68 12.0 0.899 54.6 0.817 44.6 0.8
-k-LLMmeans-1 81.7 0.253 91.6 0.215 13.6 0.351 71.2 0.334 65.6 0.248
-k-LLMmeans-5 81.8 0.253 91.9 0.21 13.7 0.348 71.5 0.325 65.2 0.25
-k-LLMmeans-FS1 82.0 0.246 91.8 0.208 13.9 0.348 71.4 0.333 66.2 0.243
-k-LLMmeans-FS5 82.2 0.247 92.5 0.198 13.9 0.346 71.9 0.337 65.6 0.254
-Table 1: Average Normalized Mutual Information (NMI) and distance between final and true centroids (dist) for k-means, k-medoids, and four k-LLMmeans variants across 10 random seeds on four benchmark datasets (including both domain and intent from MASSIVE), using four different embedding models.
-6.1 Datasets
-We evaluate our clustering approach on four benchmark datasets:
-  • Bank77 (Casanueva et al., 2020): Consists of 3,080 customer queries related to banking services, categorized into 77 distinct intents.
-  • CLINC (Larson et al., 2019): A diverse set of 4,500 queries spanning 150 intent classes across multiple domains, designed for open-domain intent classification.
-  • GoEmo (Demszky et al., 2020): Contains 2,984 social media posts annotated with 27 fine-grained emotion categories.
-We removed the neutral expressions to address data imbalance and retained only entries with a single, unique emotion.
-  • MASSIVE (FitzGerald et al., 2023): Comprises 2,974 English-language virtual assistant utterances grouped into 18 domains and 59 intent categories.
-These datasets provide a robust evaluation setting for text clustering across different domains and classification granularities.
-6.2 New compiled dataset for testing text-Streaming Clustering Algorithms
-We extract and unify a challenging data stream comprising unique archive posts collected from 84 Stack Exchange sites (StackExchange, 2024).
-Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods.
-Our raw dataset spans 84 domains, each containing at least 20 posts per year from 2018 to 2023 (with post lengths ranging from 20 to 1000 characters), totaling 499,359 posts.
-For our experiments, we focus on posts from 2020 to 2023 and further filter out labels that do not exceed 500 posts in 2023.
-The resulting subset comprises 35 distinct groups and 69,147 posts.
-Both the raw and clean data are provided with this paper.
-6.3 Methods
+5 Static experiments
+We use four benchmark datasets that span diverse domains and classification granularities: Bank77 [9], CLINC [24], GoEmo [12] and MASSIVE (domain and intent) [16].
+See Appendix A.1 for detailed descriptions of the datasets.
+We evaluate our k-LLMmeans algorithm on each of the four static (non-streaming) dataset using the known number of clusters and performing 120 centroid-update iterations.
+To demonstrate the robustness of our approach, we compute embeddings with four different pretrained models: DistilBERT [36], e5-large [46], S-BERT [34], and text-embedding-3-small [30].
+For the LLM component, we use GPT-3.5-turbo, GPT-4o, Llama-3.3, Claude-3.7 and DeepSeek-V3.
+For the instruction task $I$, we employ a simple summarization prompt that adapted to each dataset.
+For example, for Bank77 we use the prompt: “The following is a cluster of online banking questions.
+Write a single question that represents the cluster concisely.” We evaluate four variants of k-LLMmeans, differing in the number of summarization steps and the prompt configuration.
+The single variant uses a single summarization step with all cluster documents as input ($l=60$), while the multiple variant performs five summarization steps with shorter prompts ($l=20$), again using all cluster documents each time.
+Additionally, we explore few-shot variants (FS-single and FS-multiple), where each prompt includes $m=10$ randomly selected documents instead of the full cluster.
+As traditional clustering baselines, we consider k-means, k-medoids [23], spectral clustering [29], agglomerative clustering [22], and Gaussian Mixture Models (GMM) [11].
+We use ground-truth number of clusters and ten different seeds.
+For advanced LLM-based methods, we compare with results obtained by Feng et al.
+[15] on ClusterLLM [51], IDAS [10], and LLMEdgeRefine [15].
+5.1 Results with static data
+Table 1: Average Accuracy (ACC) and Normalized Mutual Information (NMI) for four k-LLMmeans variants using GPT-4o, compared against traditional clustering baselines, with text-embedding-3-small embeddings, evaluated on four benchmark datasets.
+Dataset/Method BANK77 CLINC GoEmo Massive (D) Massive (I)
+ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI
+k-LLMmeans
+   single 67.1 83.6 78.1 92.5 24.0 22.3 61.1 69.6 54.0 73.0
+   multiple 66.7 83.6 79.1 92.8 24.1 22.1 60.6 69.5 55.8 73.5
+   FS-single 67.5 83.8 79.2 92.8 23.0 21.9 60.8 69.4 55.3 73.6
+   FS-multiple 67.9 84.1 80.2 93.1 24.2 22.3 63.2 70.6 56.4 73.9
+k-means 66.2 83.0 77.3 92.0 20.7 20.5 59.4 67.9 52.9 72.4
+k-medoids 41.7 69.5 49.3 77.7 15.7 15.9 37.6 38.8 35.9 52.4
+GMM 67.7 83.0 78.9 92.7 21.5 20.6 56.2 68.6 53.9 73.2
+Agglomerative 69.9 83.7 81.0 92.5 15.8 14.1 62.8 67.1 56.6 70.5
+Spectral 68.2 83.3 76.3 90.9 17.6 15.2 61.5 67.0 56.7 71.4
+Table 2: Average ACC, NMI, and distance between final and true centroids ($dist$) for k-means and k-LLMmeans FS-multiple, evaluated on three datasets using four different embedding models.
+Dataset CLINC GoEmo Massive (D)
+embedding/Method ACC NMI $dist$ ACC NMI $dist$ ACC NMI $dist$
+DistilBERT
+   k-means 53.9 77.2 0.340 17.8 18.3 0.364 44.4 45.1 0.309
+   k-LLMmeans 55.3 78.7 0.343 18.2 18.8 0.351 46.2 46.4 0.295
+text-embedding-3-small
+   k-means 77.3 92.0 0.200 20.7 20.5 0.287 59.4 67.9 0.246
+   k–LLMmeans 80.2 93.1 0.179 24.2 22.3 0.278 63.2 70.6 0.227
+e5-large
+   k-means 73.8 90.8 0.131 22.8 22.8 0.176 58.4 63.7 0.138
+   k-LLMmeans 77.2 92.5 0.119 24.2 24.3 0.168 62.3 65.9 0.133
+S-BERT
+   k-means 76.9 91.0 0.215 13.7 13.3 0.355 58.2 64.6 0.271
+   k-LLMmeans 79.7 92.5 0.198 14.7 13.9 0.346 59.7 65.6 0.254
+Table 3: Number of LLM calls (prompts), average ACC, and average NMI for k-LLMmeans FS-multiple using various LLMs and e5-large embeddings, compared against other LLM-based text clustering methods on three benchmark datasets.
+Dataset/Method CLINC GoEmo Massive (D)
+prompts ACC NMI prompts ACC NMI prompts ACC NMI
+k-LLMmeans
+   GPT-3.5 750 76.0 92.1 135 23.9 24.1 90 63.3 66.2
+   GPT-4o 750 77.2 92.5 135 24.2 24.3 90 62.3 65.9
+   Llama-3.3 750 77.2 92.3 135 23.8 23.1 90 62.0 66.3
+   DeepSeek-V3 750 69.7 90.8 135 22.8 22.7 90 62.6 66.3
+   Claude-3.7 750 76.9 92.5 135 24.2 23.7 90 61.8 66.0
+ClusterLLM 1618 83.8 94.0 1618 26.8 23.9 1618 60.9 68.8
+IDAS 4650 81.4 92.4 3011 30.6 25.6 2992 53.5 63.9
+LLMEdgeRefine 1350 86.8 94.9 895 34.8 29.7 892 63.1 68.7
+Comparing k-LLMmeans variants and traditional clustering methods.
+Table 1 reports Average Accuracy (ACC) and Normalized Mutual Information (NMI) for four k-LLMmeans variants using GPT-4o, compared against traditional clustering baselines, with text-embedding-3-small embeddings, evaluated on four benchmark datasets.
+Overall, all k-LLMmeans variants outperform traditional clustering methods in terms of NMI and in general achieve higher average accuracy across datasets.
+Within k-LLMmeans variants there are minor differences, although running multiple summarization steps tends to offer better performance.
+Interestingly, using few-shot summarization generally leads to better results than prompting on the entire cluster.
+These observations suggest k-LLMmeans improves over k-means while remaining efficient, as few-shot summarization appears sufficient without needing extensive prompt length (See Appendix B for standard errors of all simulations).
+[…]
+Comparing k-LLMmeans with different LLMs and state-of-the-art LLM-based clustering methods.
+Table 2 reports number of LLM calls (prompts) and Average Accuracy (ACC) and Normalized Mutual Information (NMI) for the k-LLMmeans FS-multiple variant on three benchmark datasets, pairing e5-large embeddings with several backbone LLMs.
+Performance is stable across LLM choices, indicating that k-LLMmeans is robust to the specific language model.
+For reference, Table 2 also reports results from Feng et al.
+[15] for ClusterLLM, IDAS, and LLMEdgeRefine.
+All three baselines rely on GPT-3.5: ClusterLLM (with fine-tuning) and IDAS use the same e5-large embeddings as we do, whereas LLMEdgeRefine benefits from the stronger Instructor embeddings [21].
+Even so, our GPT-3.5 configuration attains ACC and NMI that are only marginally lower—and often comparable—while requiring orders of magnitude fewer LLM calls and no fine-tuning.
+This substantial reduction in model prompts translates directly into improved scalability and cost-efficiency.
+6 Sequential (mini-batch) experiments
+We use yearly posts from 2020 to 2023 from 35 StackExchange sites [41].
+Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods (See Appendix A.2 for details).
+For each yearly subset, we split the data into $b=\lceil\tfrac{n}{10000}\rceil$ equal-sized batches ${D_{1},\dots,D_{b}}$ in chronological order, where $n$ is the number of documents for that year.
+We then use the ground-truth clusters and run the mini-batch k-LLMmeans algorithm in its four variants described in Section 5 (for practical reasons for the full cluster method we use 50 samples).
+For comparison, we evaluate three baselines: mini-batch k-means with standard random sampling across the entire year (until convergence), sequential mini-batch k-means with $b$ chronological batches, and standard k-means on the full dataset.
+In all experiments, we use text-embedding-3-small and GPT-4o for the summarization step.
+6.1 Results with streaming data
+Table 4: Average ACC, and average NMI for four sequential mini-batch variants, k-means, mini-batch k-means, sequential mini-batch k-means on the yearly StackExchange data.
 Year/Method 2020 2021 2022 2023
 (69147 posts) (54322 posts) (43521 posts) (38953 posts)
-NMI dist NMI dist NMI dist NMI dist
-k-means 80.6 0.14 79.0 0.167 79.0 0.154 79.6 0.138
-mini-batch k-means 78.2 0.181 77.4 0.166 77.6 0.173 77.0 0.175
-seq. mini-batch k-means 76.6 0.187 75.2 0.185 75.6 0.187 74.8 0.184
-mini-batch k-LLMmeans-1 81.2 0.141 80.3 0.138 79.4 0.145 80.0 0.136
-mini-batch k-LLMmeans-5 80.9 0.144 80.5 0.127 80.5 0.125 80.3 0.129
-mini-batch k-LLMmeans-FS1 81.1 0.136 79.8 0.141 79.3 0.147 79.8 0.138
-mini-batch k-LLMmeans-FS5 81.6 0.126 80.2 0.126 80.1 0.129 80.1 0.133
-Table 2: Average Normalized Mutual Information (NMI) and distance from final and true centroids (dist) for k-means, mini-batch k-means, sequential mini-batch k-means and four sequential mini-batch k-LLMmeans variants with across 5 random seeds on StackExchange data.
-We evaluate our k-LLMmeans algorithm on each of the four static (non-streaming) dataset using the known number of clusters and performing 120 centroid-update iterations and 10 different seeds.
-To demonstrate the robustness of our approach, we compute embeddings with four different pretrained models: distilbert (Sanh, 2019), e5-large (Wang et al., 2022), s-bert (Reimers and Gurevych, 2019), and OpenAI’s text-embedding-3-small (OpenAI, 2023).
-For the LLM component, we only use OpenAI’s gpt-4o to ensure any observed differences in performance arise from the effectiveness of our clustering method rather than the inherent strengths or weaknesses of different LLMs.
-For the instruction task $I$, we employ a simple summarization prompt that adapted to each dataset.
-For example, for Bank77 we use the prompt: “The following is a cluster of online banking questions.
-Write a single question that represents the cluster concisely.” We examine four variations of k-LLMmeans based on different numbers of summarization steps and the size of the prompts:
-  • k-LLMmeans-1: A single summarization step ($l=60$) using all documents in the cluster as input.
-  • k-LLMmeans-5: Five summarization steps ($l=20$) with all documents in the cluster as input at each step.
-  • k-LLMmeans-FS1: A few-shot variant with 10 randomly selected documents in a single summarization step.
-  • k-LLMmeans-FS5: A few-shot variant with 10 randomly selected documents in five summarization steps.
-Our baselines are the standard centroid-based algorithms k-means and k-medoids.
-While alternative clustering methods may achieve stronger performance, our primary goal is to demonstrate improvement specifically over widely used centroid-based approaches.
-Methods for streaming dataset.
-We partition the StackExchange data into four subsets, each corresponding to a single year from 2020 to 2023 (See Table 2 for yearly dataset sizes).
-For each yearly subset, we split the data into $b=\lceil\tfrac{n}{10000}\rceil$ equal-sized batches ${D_{1},\dots,D_{b}}$ in chronological order, where $n$ is the number of documents for that year.
-We then use the ground-truth clusters and run the mini-batch k-LLMmeans algorithm in its four previously described variants (for practical reasons for the full cluster method we use 50 samples).
-For comparison, we evaluate three baselines: mini-batch k-means with standard random sampling across the entire year (until convergence), sequential mini-batch k-means with $b$ chronological batches, and standard k-means on the full dataset.
-In all experiments, we use OpenAI’s models for both embeddings and the LLM.
-6.4 Results
-Table 1 reports the average Normalized Mutual Information (NMI) across all method–dataset–embedding combinations, as well as the average distance (dist) between the centroids produced by each method and the true centroids calculated with the ground truth clusters.
-A lower distance indicates greater proximity to the optimal solution.
-This distance aligns well with the objectives of our approach.
-We also measure Accuracy (not shown here due to space constraints), observing results consistent with the reported metrics.
-Overall, all k-LLMmeans variants outperform k-means except when using the distilbert embedding.
-However, the performance of every method using this embedding is clearly inferior compared to those employing any other embedding.
-Meanwhile, k-medoids emerges as the worst performer.
-Comparing the four k-LLMmeans configurations reveals only minor differences, although running five summarization steps ($l=20$) tends to offer better performance.
-Interestingly, using few-shot summarization generally leads to better results than prompting on the entire cluster.
-These observations suggest k-LLMmeans improves over k-means while remaining efficient, as few-shot summarization appears sufficient without needing extensive LLM usage.
-We present the same evaluation metrics for sequential data in Table 2.
+ACC NMI ACC NMI ACC NMI ACC NMI
+mini-batch k-LLMmeans
+   single 73.1 81.2 73.2 80.3 70.8 79.4 72.3 80.0
+   multiple 72.5 80.9 73.6 80.5 74.4 80.5 73.4 80.3
+   FS + single 74.0 81.1 72.2 79.8 70.2 79.3 71.6 79.8
+   FS + multiple 75.4 81.6 73.5 80.2 72.8 80.1 72.7 80.1
+k-means 73.4 80.6 67.7 79.0 68.6 79.0 72.0 79.6
+mini-batch k-means 67.0 78.2 67.7 77.4 67.5 77.6 67.2 77.0
+seq. mini-batch k-means 67.0 76.6 66.7 75.2 65.6 75.6 65.8 74.8
+Table 4 presents ACC and NMI on the yearly StackExchange corpus, comparing four sequential mini-batch variants of k-LLMmeans against standard k-means, mini-batch k-means, and sequential mini-batch k-means across five random seeds.
 The mini-batch k-LLMmeans methods consistently outperform all three baselines across all scenarios.
@@
 7 Case study
-Figure 2: Sequential evolution of the LLM-generated centroids for four primary clusters during the three batches of the sequential mini-batch k-LLMmeans process applied to 2021 posts from the AI Stack Exchange site (StackExchange, 2024).
-Main aspects are manually highlighted in different color on each cluster.
-To demonstrate the interpretability of our method in capturing the evolution of clusters within sequential data, we present a case study using posts from the AI site in the 2021 Stack Exchange dataset (StackExchange, 2024).
+To demonstrate the interpretability of our method in capturing the evolution of clusters within sequential data, we present a case study using posts from the AI site in the 2021 Stack Exchange dataset [41].
 We apply our mini-batch k-LLMmeans algorithm with three equal-length batches and a total of ten clusters.
@@
 More broadly, this demonstrates how our interpretable mini-batch k-LLMmeans clustering algorithm, applied to sequential text streaming, can help practitioners track topic evolution, improve decision-making, and enhance transparency in dynamic information flows.
-8 Conclusions
-We introduced k-LLMmeans and mini-batch k-LLMmeans, unsupervised clustering algorithms that refine k-means by integrating a lightweight LLM-based summarization into centroid updates.
-This modification enriches the contextual representation while preserving k-means’ efficiency and scalability, yielding interpretable clusters for both static corpus of documents and streaming text data.
-Limitations
-Our method relies on both text embeddings and LLM queries, making it sensitive to the quality and biases of the underlying language models.
-Any inaccuracies in these models may propagate into the clustering results.
+Figure 3: Sequential evolution of the LLM-generated centroids for four primary clusters during the three batches of the sequential mini-batch k-LLMmeans process applied to 2021 posts from the AI Stack Exchange site [41].
+Main aspects are manually highlighted in different color on each cluster.
+8 Related work
+Traditional clustering.
+Hierarchical methods [22, 5] build tree-structured representations of nested document relationships.
+Density-based approaches like DBSCAN [14] and graph-based methods detect clusters of arbitrary shapes, while spectral clustering [29] leverages eigen-decomposition to uncover complex structures.
+Model-based techniques—including Gaussian mixture models [11] and recent neural network frameworks [52, 18, 49, 50, 48]—provide probabilistic clustering formulations.
+Additionally, topic modeling methods, from probabilistic latent semantic analysis [17] to latent Dirichlet allocation [6], capture word co-occurrence patterns and latent topics.
+LLM-based clustering.
+Viswanathan et al.
+[45] employ LLMs to augment document representations, generate pseudo pairwise constraints, and post-correct low-confidence assignments for query-efficient, few-shot semi-supervised clustering.
+Zhang et al.
+[51] propose ClusterLLM, which uses instruction-tuned LLMs via interactive triplet and pairwise feedback to cost-effectively refine clustering granularity.
+Complementary approaches [44, 33] show that context-derived representations capture subtle semantic nuances beyond traditional embeddings.
+Wang et al.
+[47] introduce a goal-driven, explainable method that employs natural language descriptions to clarify cluster boundaries, while [10] present IDAS for intent discovery using abstractive summarization.
+Feng et al.
+[15] propose LLMEdgeRefine an iterative mechanism that forms super-points to mitigate outliers and reassign ambiguous edge points, resulting in clusters with higher coherence and robustness.
+9 Discussion
+We introduced k-LLMmeans and mini-batch k-LLMmeans, unsupervised text clustering algorithms that refine k-means by integrating a lightweight LLM-based summarization into centroid updates.
+This modification enriches the contextual representation while preserving k-means’ stability and scalability, yielding interpretable clusters for both static corpus of documents and streaming text data, with low LLM usage.
+Limitations.
+Our approach depends on pre-computed embeddings and LLM-generated summaries, so any bias or error in the underlying language models can propagate to the final clusters.
 While we demonstrate that a simple instruction-based approach is effective, the design of these instructions can still influence outcomes.
@@
 While this is generally practical, it may become a limitation when dealing with highly complex or heterogeneous clusters.
-Finally, similar to k-means, our k-LLMmeans method requires specifying the number of clusters in advance, which may not always align with the true structure of the data.
+Appendix A Datasets
+A.1 Non-streaming datasets
+We evaluate our clustering approach on four benchmark datasets.
+Bank77 [9]: Consists of 3,080 customer queries related to banking services, categorized into 77 distinct intents.
+CLINC [24]: A diverse set of 4,500 queries spanning 150 intent classes across multiple domains, designed for open-domain intent classification.
+GoEmo [12]: Contains 2,984 social media posts annotated with 27 fine-grained emotion categories.
+We removed the neutral expressions to address data imbalance and retained only entries with a single, unique emotion.
+MASSIVE [16]: Comprises 2,974 English-language virtual assistant utterances grouped into 18 domains and 59 intent categories.
+These datasets provide a robust evaluation setting for text clustering across different domains and classification granularities.
+A.2 New compiled dataset for testing text-Streaming Clustering Algorithms
+We extract and unify a challenging data stream comprising unique archive posts collected from 84 Stack Exchange sites [41].
+Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods.
+Our raw dataset spans 84 domains, each containing at least 20 posts per year from 2018 to 2023 (with post lengths ranging from 20 to 1000 characters), totaling 499,359 posts.
+For our experiments, we focus on posts from 2020 to 2023 and further filter out labels that do not exceed 500 posts in 2023.
+The resulting subset comprises 35 distinct groups and 69,147 posts.
+Both the raw and clean data are provided with this paper.
+Stack Exchange content is licensed under the Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) license.
+Appendix B Extended simulation results
+B.1 Static experiments
+Tables 59 present extended results from the simulations described in Section 5.
+For each dataset, we report the average Accuracy (ACC), Normalized Mutual Information (NMI), and distance between final and true centroids ($dist$), along with standard deviations computed over 10 random seeds.
+Results are provided for all k-LLMmeans variants using GPT-4o, as well as for traditional clustering algorithms, evaluated under four different embedding types.
+Table 10 show extended results for all benchmark datasets and multiple LLMs.
+B.2 Streaming (mini-batch) experiments
+Table 11 report the standard deviation for results presented on Table 4.
+Appendix C Computational resources for experiments
+Our simulations span approximately one week on a standard laptop without specialized hardware, with the primary computational bottleneck arising from LLM usage.
+We conducted API-based experiments using GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3.
+Most experiments relied on OpenAI’s API (GPT-4o and text-embedding-3-small), incurring a cost of roughly \$700 and consuming around 120 million tokens.
+Experiments involving the other LLMs collectively cost under \$30.
+Appendix D Broader Impacts
+While k-LLMmeans is primarily a foundational method aimed at improving scalability and interpretability in LLM-based text clustering, potential societal risks arise from the misuse of generated cluster summaries.
+For example, if used in sensitive domains such as legal, medical, or financial text analysis, misleading or biased summaries could propagate incorrect conclusions.
+Additionally, because the method relies on pretrained language models, any biases or factual inaccuracies embedded in those models may influence clustering outcomes.
+To mitigate these risks, we recommend cautious deployment in high-stakes settings, transparency about the LLMs used, and regular human-in-the-loop validation when interpretability affects downstream decisions.
+Table 5: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids (dist) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on Bank77 using four different embedding models (standard deviations in parenthesis).
+embedding/Method ACC NMI $dist$
+DistilBERT
+k-LLMmeans
+   single 43.8(1.12) 64.8(0.493) 0.351(0.013)
+   multiple 43.5(0.991) 64.8(0.516) 0.358(0.006)
+   FS-single 43.7(0.75) 64.8(0.48) 0.352(0.007)
+   FS-multiple 43.8(1.032) 64.9(0.362) 0.363(0.011)
+k-means 43.653(0.714) 64.389(0.456) 0.335(0.007)
+k-medoids 32.89(0.0) 55.991(0.0) 0.591(0.0)
+GMM 44.701(1.163) 64.481(0.63) 0.318(0.005)
+Agglomerative 45.974(0.0) 67.064(0.0) 0.33(0.0)
+Spectral 45.925(0.201) 66.057(0.241) 0.365(0.001)
+text-embedding-3-small
+k-LLMmeans
+   single 67.1(1.216) 83.6(0.459) 0.221(0.012)
+   multiple 66.7(0.778) 83.6(0.2) 0.226(0.009)
+   FS-single 67.5(1.072) 83.8(0.358) 0.219(0.011)
+   FS-multiple 67.9(1.546) 84.1(0.412) 0.22(0.014)
+k-means 66.25(1.815) 83.005(0.559) 0.225(0.014)
+k-medoids 41.688(0.0) 69.457(0.0) 0.639(0.0)
+GMM 68.792(1.784) 83.048(0.543) 0.198(0.013)
+Agglomerative 69.87(0.0) 83.741(0.0) 0.2(0.0)
+Spectral 68.562(0.562) 83.746(0.38) 0.227(0.003)
+e5-large
+k-LLMmeans
+   single 61.1(1.041) 78.6(0.412) 0.138(0.006)
+   multiple 61.6(0.783) 79.0(0.355) 0.136(0.005)
+   FS-single 61.5(1.218) 78.8(0.484) 0.136(0.004)
+   FS-multiple 62.7(1.706) 79.5(0.646) 0.134(0.007)
+k-means 58.097(1.732) 76.703(0.795) 0.142(0.007)
+k-medoids 37.175(0.0) 64.487(0.0) 0.331(0.0)
+GMM 56.977(1.384) 75.801(0.502) 0.142(0.006)
+Agglomerative 61.623(0.0) 79.259(0.0) 0.138(0.0)
+Spectral 63.195(0.69) 79.684(0.312) 0.144(0.003)
+S-BERT
+k-LLMmeans
+   single 64.4(1.83) 81.7(0.563) 0.253(0.014)
+   multiple 64.6(1.393) 81.8(0.389) 0.253(0.013)
+   FS-single 65.4(1.693) 82.0(0.515) 0.246(0.013)
+   FS-multiple 65.7(1.841) 82.2(0.648) 0.247(0.013)
+k-means 63.643(2.272) 80.937(0.744) 0.255(0.016)
+k-medoids 41.299(0.0) 68.985(0.0) 0.64(0.0)
+GMM 63.565(2.014) 80.504(0.579) 0.248(0.017)
+Agglomerative 64.643(0.0) 81.601(0.0) 0.255(0.0)
+Spectral 65.737(0.652) 83.247(0.395) 0.268(0.006)
+Table 6: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids ($dist$) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on CLINC using four different embedding models (standard deviations in parenthesis).
+embedding/Method ACC NMI $dist$
+DistilBERT
+k-LLMmeans
+   single 53.3(1.468) 77.4(0.547) 0.352(0.01)
+   multiple 53.5(1.141) 77.7(0.437) 0.356(0.01)
+   FS-single 54.7(1.117) 78.1(0.552) 0.346(0.008)
+   FS-multiple 55.3(1.104) 78.7(0.531) 0.343(0.009)
+k-means 53.889(0.956) 77.188(0.494) 0.34(0.007)
+k-medoids 36.889(0.0) 66.765(0.0) 0.69(0.0)
+GMM 53.42(1.362) 76.534(0.442) 0.334(0.01)
+Agglomerative 55.333(0.0) 78.805(0.0) 0.343(0.0)
+Spectral 52.289(0.395) 76.215(0.268) 0.364(0.004)
+text-embedding-3-small
+k-LLMmeans
+   single 78.1(1.67) 92.5(0.474) 0.194(0.015)
+   multiple 79.1(1.675) 92.8(0.48) 0.186(0.015)
+   FS-single 79.2(1.725) 92.8(0.427) 0.186(0.015)
+   FS-multiple 80.2(1.523) 93.1(0.323) 0.179(0.015)
+k-means 77.253(1.828) 91.965(0.351) 0.2(0.014)
+k-medoids 49.289(0.0) 77.718(0.0) 0.694(0.0)
+GMM 80.484(1.181) 92.492(0.222) 0.168(0.011)
+Agglomerative 81.0(0.0) 92.544(0.0) 0.169(0.0)
+Spectral 76.02(0.401) 90.812(0.139) 0.214(0.003)
+e5-large
+k-LLMmeans
+   single 75.2(1.917) 91.6(0.48) 0.127(0.008)
+   multiple 75.4(1.858) 92.0(0.472) 0.127(0.009)
+   FS-single 75.4(1.746) 91.7(0.46) 0.127(0.008)
+   FS-multiple 77.2(1.604) 92.5(0.399) 0.119(0.007)
+k-means 73.831(2.338) 90.77(0.739) 0.131(0.01)
+k-medoids 41.911(0.0) 73.371(0.0) 0.371(0.0)
+GMM 77.647(1.286) 91.41(0.279) 0.106(0.006)
+Agglomerative 80.333(0.0) 92.544(0.0) 0.101(0.0)
+Spectral 72.853(0.252) 89.828(0.092) 0.128(0.002)
+S-BERT
+k-LLMmeans
+   single 77.3(1.475) 91.6(0.381) 0.215(0.013)
+   multiple 77.9(1.373) 91.9(0.436) 0.21(0.012)
+   FS-single 78.0(1.115) 91.8(0.332) 0.208(0.01)
+   FS-multiple 79.7(0.734) 92.5(0.247) 0.198(0.008)
+k-means 76.853(1.63) 91.036(0.496) 0.215(0.013)
+k-medoids 51.156(0.0) 78.506(0.0) 0.68(0.0)
+GMM 77.333(0.93) 91.025(0.254) 0.203(0.009)
+Agglomerative 82.622(0.0) 92.78(0.0) 0.158(0.0)
+Spectral 77.167(0.36) 90.794(0.105) 0.215(0.002)
+Table 7: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids ($dist$) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on GoEmo using four different embedding models (standard deviations in parenthesis).
+embedding/Method ACC NMI $dist$
+DistilBERT
+k-LLMmeans
+   single 17.5(0.467) 18.3(0.447) 0.358(0.009)
+   multiple 17.9(0.861) 18.3(0.62) 0.354(0.006)
+   FS-single 18.0(0.681) 18.9(0.407) 0.357(0.006)
+   FS-multiple 18.2(0.52) 18.8(0.733) 0.351(0.004)
+k-means 17.755(0.857) 18.267(0.667) 0.364(0.005)
+k-medoids 15.851(0.0) 13.964(0.0) 0.769(0.0)
+GMM 18.07(0.819) 18.729(0.471) 0.36(0.005)
+Agglomerative 16.857(0.0) 15.402(0.0) 0.352(0.0)
+Spectral 17.198(0.21) 17.23(0.189) 0.42(0.002)
+text-embedding-3-small
+k-LLMmeans
+   single 24.0(1.346) 22.3(0.572) 0.278(0.007)
+   multiple 24.1(1.326) 22.1(0.651) 0.275(0.004)
+   FS-single 23.0(0.937) 21.9(0.638) 0.283(0.011)
+   FS-multiple 24.2(1.156) 22.3(0.67) 0.278(0.011)
+k-means 20.734(0.997) 20.466(0.672) 0.287(0.005)
+k-medoids 15.65(0.0) 15.943(0.0) 0.852(0.0)
+GMM 20.831(1.418) 20.123(0.524) 0.296(0.007)
+Agglomerative 15.818(0.0) 14.098(0.0) 0.286(0.0)
+Spectral 17.661(0.328) 15.045(0.319) 0.341(0.004)
+e5-large
+k-LLMmeans
+   single 23.4(0.875) 23.4(0.694) 0.176(0.007)
+   multiple 23.4(0.9) 23.7(0.521) 0.174(0.004)
+   FS-single 23.4(1.089) 23.7(0.356) 0.174(0.003)
+   FS-multiple 24.2(0.761) 24.3(0.702) 0.168(0.003)
+k-means 22.778(0.917) 22.761(0.729) 0.176(0.003)
+k-medoids 14.812(0.0) 15.667(0.0) 0.447(0.0)
+GMM 21.367(0.986) 22.197(0.69) 0.181(0.003)
+Agglomerative 19.571(0.0) 18.544(0.0) 0.177(0.0)
+Spectral 23.388(0.524) 21.362(0.257) 0.192(0.001)
+S-BERT
+k-LLMmeans
+   single 14.4(0.739) 13.6(0.547) 0.351(0.006)
+   multiple 14.4(0.75) 13.7(0.59) 0.348(0.005)
+   FS-single 14.5(0.79) 13.9(0.413) 0.348(0.006)
+   FS-multiple 14.7(0.551) 13.9(0.501) 0.346(0.004)
+k-means 13.737(0.617) 13.269(0.574) 0.355(0.005)
+k-medoids 13.941(0.0) 11.974(0.0) 0.899(0.0)
+GMM 14.31(0.744) 13.628(0.482) 0.353(0.003)
+Agglomerative 12.399(0.0) 10.766(0.0) 0.355(0.0)
+Spectral 13.05(0.211) 11.777(0.123) 0.416(0.003)
+Table 8: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids ($dist$) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on MASSIVE (Domain) using four different embedding models (standard deviations in parenthesis).
+embedding/Method ACC NMI $dist$
+DistilBERT
+k-LLMmeans
+   single 45.8(1.508) 46.0(1.164) 0.294(0.016)
+   multiple 45.2(2.264) 45.6(1.199) 0.306(0.017)
+   FS-single 45.3(1.789) 45.7(1.099) 0.299(0.021)
+   FS-multiple 46.2(1.348) 46.4(1.119) 0.295(0.015)
+k-means 44.368(1.528) 45.065(0.988) 0.309(0.021)
+k-medoids 33.457(0.0) 27.223(0.0) 0.81(0.0)
+GMM 45.451(1.892) 46.966(0.74) 0.293(0.02)
+Agglomerative 35.777(0.0) 39.931(0.0) 0.355(0.0)
+Spectral 40.212(2.693) 44.476(1.514) 0.473(0.009)
+text-embedding-3-small
+k-LLMmeans
+   single 61.1(3.393) 69.6(1.604) 0.232(0.018)
+   multiple 60.6(2.822) 69.5(1.661) 0.238(0.016)
+   FS-single 60.8(3.206) 69.4(1.091) 0.233(0.02)
+   FS-multiple 63.2(2.843) 70.6(1.379) 0.227(0.022)
+k-means 59.391(4.026) 67.869(1.671) 0.246(0.015)
+k-medoids 37.559(0.0) 38.788(0.0) 0.797(0.0)
+GMM 59.391(2.883) 68.0(1.01) 0.253(0.014)
+Agglomerative 62.845(0.0) 67.07(0.0) 0.222(0.0)
+Spectral 61.51(0.022) 66.987(0.023) 0.331(0.0)
+e5-large
+k-LLMmeans
+   single 60.6(2.856) 64.8(1.611) 0.134(0.014)
+   multiple 61.5(4.447) 66.0(2.105) 0.131(0.016)
+   FS-single 61.0(3.441) 64.9(1.42) 0.136(0.019)
+   FS-multiple 62.3(1.358) 65.9(0.776) 0.133(0.007)
+k-means 58.426(3.889) 63.663(1.748) 0.138(0.02)
+k-medoids 41.19(0.0) 36.343(0.0) 0.405(0.0)
+GMM 63.073(3.473) 66.018(1.588) 0.132(0.008)
+Agglomerative 66.073(0.0) 66.279(0.0) 0.132(0.0)
+Spectral 57.001(0.02) 63.327(0.018) 0.202(0.0)
+S-BERT
+k-LLMmeans
+   single 61.0(3.797) 65.6(1.36) 0.248(0.026)
+   multiple 60.1(2.016) 65.2(0.735) 0.25(0.019)
+   FS-single 61.7(2.529) 66.2(1.134) 0.243(0.026)
+   FS-multiple 59.7(2.543) 65.6(0.795) 0.254(0.019)
+k-means 58.248(2.029) 64.608(1.09) 0.271(0.023)
+k-medoids 45.763(0.0) 44.605(0.0) 0.8(0.0)
+GMM 57.935(3.645) 64.404(1.094) 0.27(0.032)
+Agglomerative 63.416(0.0) 64.72(0.0) 0.222(0.0)
+Spectral 58.487(0.85) 63.148(0.49) 0.367(0.005)
+Table 9: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids ($dist$) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on MASSIVE (Intention) using four different embedding models (standard deviations in parenthesis).
+embedding/Method ACC NMI $dist$
+DistilBERT
+k-LLMmeans
+   single 44.1(1.099) 58.6(0.511) 0.362(0.014)
+   multiple 43.8(1.494) 59.0(0.46) 0.363(0.013)
+   FS-single 43.5(1.196) 58.8(0.59) 0.365(0.01)
+   FS-multiple 43.6(1.343) 59.0(0.594) 0.362(0.012)
+k-means 43.124(1.465) 58.102(0.516) 0.366(0.014)
+k-medoids 25.084(0.0) 42.582(0.0) 0.847(0.0)
+GMM 41.379(1.249) 57.375(0.452) 0.367(0.013)
+Agglomerative 39.139(0.0) 55.378(0.0) 0.386(0.0)
+Spectral 46.022(0.482) 58.237(0.234) 0.378(0.004)
+text-embedding-3-small
+k-LLMmeans
+   single 54.0(1.513) 73.0(0.582) 0.314(0.013)
+   multiple 55.8(1.396) 73.5(0.762) 0.307(0.013)
+   FS-single 55.3(1.759) 73.6(0.774) 0.314(0.014)
+   FS-multiple 56.4(1.32) 73.9(0.7) 0.302(0.012)
+k-means 52.878(1.072) 72.373(0.541) 0.32(0.011)
+k-medoids 35.911(0.0) 52.442(0.0) 0.86(0.0)
+GMM 54.56(1.918) 73.105(0.798) 0.306(0.015)
+Agglomerative 56.59(0.0) 70.461(0.0) 0.29(0.0)
+Spectral 58.453(0.837) 71.797(0.23) 0.28(0.002)
+e5-large
+k-LLMmeans
+   single 55.2(1.625) 72.0(0.515) 0.17(0.01)
+   multiple 55.6(2.161) 72.1(0.725) 0.166(0.012)
+   FS-single 54.3(2.395) 71.6(0.833) 0.173(0.012)
+   FS-multiple 55.8(1.995) 72.4(0.912) 0.166(0.01)
+k-means 52.885(1.264) 70.917(0.708) 0.175(0.01)
+k-medoids 32.112(0.0) 50.313(0.0) 0.437(0.0)
+GMM 50.595(2.28) 70.15(1.225) 0.183(0.007)
+Agglomerative 55.447(0.0) 70.315(0.0) 0.17(0.0)
+Spectral 56.594(0.585) 69.034(0.417) 0.152(0.003)
+S-BERT
+k-LLMmeans
+   single 52.2(2.754) 71.2(0.785) 0.334(0.022)
+   multiple 52.9(2.005) 71.5(0.675) 0.325(0.015)
+   FS-single 52.0(2.552) 71.4(0.507) 0.333(0.019)
+   FS-multiple 53.3(1.768) 71.9(0.466) 0.337(0.011)
+k-means 50.908(1.747) 70.708(0.296) 0.344(0.019)
+k-medoids 35.978(0.0) 54.587(0.0) 0.817(0.0)
+GMM 50.494(1.22) 70.559(0.488) 0.338(0.012)
+Agglomerative 52.186(0.0) 69.544(0.0) 0.313(0.0)
+Spectral 56.126(1.395) 70.227(0.463) 0.304(0.005)
+Table 10: Average accuracy (ACC) and normalized mutual information (NMI) for k-LLMmeans FS-multiple using various LLMs and e5-large embeddings on all benchmark datasets (standard deviations in parenthesis).
+Dataset/Method Bank77 CLINC GoEmo Massive (D) Massive (I)
+ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI
+k-LLMmeans
+   GPT-3.5 56.6 77.6 76.0 92.1 23.9 24.1 63.3 66.2 56.9 72.5
+(1.32) (0.67) (1.16) (0.28) (0.65) (0.39) (1.47) (0.62) (1.62) (0.86)
+   GPT-4o 62.7 79.5 77.2 92.5 24.2 24.3 62.3 65.9 55.8 72.4
+(1.71) (0.65) (1.6) (0.4) (0.76) (0.7) (1.36) (0.78) (1.99) (0.91)
+   Llama-3.3 61.6 79.2 77.2 92.3 23.8 23.1 62.0 66.3 56.8 72.9
+(1.7) (0.63) (1.6) (0.44) (0.66) (0.62) (2.97) (1.56) (1.26) (0.5)
+   DeepSeek-V3 63.1 79.9 69.7 90.8 22.8 22.7 62.6 66.3 57.6 72.0
+(1.61) (0.75) (1.02) (0.29) (1.68) (0.88) (2.24) (0.71) (1.43) (0.61)
+   Claude-3.7 61.0 78.8 76.9 92.5 24.2 23.7 61.8 66.0 57.3 72.9
+(1.52) (0.53) (1.83) (0.52) (1.0) (0.65) (4.83) (1.99) (1.47) (0.85)
+Table 11: Standard deviation of the accuracy (ACC) and Normalized Mutual Information (NMI) for four sequential mini-batch variants, k-means, mini-batch k-means, sequential mini-batch k-means on the yearly StackExchange data.
+Year/Method 2020 2021 2022 2023
+(69147 posts) (54322 posts) (43521 posts) (38953 posts)
+ACC NMI ACC NMI ACC NMI ACC NMI
+mini-batch k-LLMmeans
+   single (3.19) (0.57) (2.48) (0.42) (2.72) (0.86) (1.59) (0.46)
+   multiple (3.02) (0.65) (2.18) (0.46) (2.48) (0.87) (2.24) (0.81)
+   FS + single (2.86) (0.39) (2.15) (0.35) (2.68) (0.9) (1.68) (0.59)
+   FS + multiple (2.02) (0.48) (2.39) (0.7) (2.28) (0.91) (1.38) (0.45)
+k-means (3.79) (0.75) (2.8) (0.25) (4.03) (0.69) (3.79) (0.56)
+mini-batch k-means (2.09) (0.9) (2.12) (0.84) (4.02) (0.98) (2.76) (0.97)
+seq. mini-batch k-means (2.39) (0.57) (4.61) (1.54) (1.23) (0.78) (2.55) (0.9)

### v3

@@
-k-LLMmeans: Scalable, Stable, and Interpretable Text Clustering via LLM-based Centroids
+Summaries as Centroids for Interpretable and Scalable Text Clustering
 Abstract
-We introduce k-LLMmeans, a novel modification of the k-means algorithm for text clustering that leverages LLM-generated summaries as cluster centroids, capturing semantic nuances often missed by purely numerical averages.
-This design preserves the core optimization properties of k-means while enhancing semantic interpretability and avoiding the scalability and instability issues typical of modern LLM-based clustering.
-Unlike existing methods, our approach does not increase LLM usage with dataset size and produces transparent intermediate outputs.
-We further extend it with a mini-batch variant for efficient, real-time clustering of streaming text.
-Extensive experiments across multiple datasets, embeddings, and LLMs show that k-LLMmeans consistently outperforms k-means and other traditional baselines and achieves results comparable to state-of-the-art LLM-based clustering—with a fraction of the LLM calls.
-Finally, we present a case study on sequential text streams and introduce a new benchmark dataset constructed from StackExchange to evaluate text-stream clustering methods.
+We introduce k-NLPmeans and k-LLMmeans, text-clustering variants of k-means that periodically replace numeric centroids with textual summaries.
+The key idea—summary-as-centroid—retains k-means assignments in embedding space while producing human-readable, auditable cluster prototypes.
+The method is LLM-optional: k-NLPmeans uses lightweight, deterministic summarizers, enabling offline, low-cost, and stable operation; k-LLMmeans is a drop-in upgrade that uses an LLM for summaries under a fixed per-iteration budget whose cost does not grow with dataset size.
+We also present a mini-batch extension for real-time clustering of streaming text.
+Across diverse datasets, embedding models, and summarization strategies, our approach consistently outperforms classical baselines and approaches the accuracy of recent LLM-based clustering—without extensive LLM calls.
+Finally, we provide a case study on sequential text streams and release a StackExchange-derived benchmark for evaluating streaming text clustering.
 1 Introduction
-Text clustering is a fundamental task in natural language processing (NLP), widely applied in document organization, topic modeling, and information retrieval [37, 42].
-A common approach involves generating text embeddings [13, 36, 27, 32, 8, 21] for each document, which are then clustered using traditional algorithms [33].
-Among these, k-means [26] is one of the most widely used, iteratively refining cluster centroids based on the mean of assigned points.
-Several alternative methods exist for computing centroids [19, 7, 23], yet they often rely on purely numerical calculations.
-While effective, such approaches risk losing crucial contextual and semantic information, as averaging high-dimensional embeddings can obscure nuanced textual relationships [34].
-We propose k-LLMmeans, a modification to the k-means algorithm by leveraging large language models (LLMs) to dynamically generate cluster centroids.
-Instead of always computing centroids as the mean of embedding vectors (k-means), in spaced iterations we generate centroids based on textual summaries of the cluster contents.
-At each of these iterations, the centroid is represented as the embedding of the LLM-generated summary of the texts belonging to the cluster (summarization step).
-LLM-based centroids preserve key contextual and semantic aspects of the main documents in the cluster while filtering out secondary or lower-priority documents [20].
-Figure 1 illustrates the k-LLMmeans algorithm with a single summarization step.
-The example demonstrates how this step enables the algorithm to escape local minima (common in k-means) and converge to the optimal solution.
-Unlike k-means, k-LLMmeans generates a concise textual summary for every centroid at each summarization step, exposing the semantic drift of clusters in real time.
-This built-in transparency simplifies debugging and validation and is especially valuable in sequential or streaming scenarios, where clusters evolve over time.
-Figure 2 illustrates one centroid through five summarization updates.
-To capitalize on this property for large-scale or online data, we adapt mini-batch k-means [38] by inserting our summarization step into a mini-batch k-LLMmeans.
-Several studies have explored unsupervised text clustering using LLMs [51, 15, 10, 45, 39, 43, 28], demonstrating state-of-the-art performance across various datasets and benchmarks.
-These approaches consistently outperform traditional clustering algorithms.
-Yet despite their impressive empirical gains, existing LLM-based clustering pipelines still face two practical obstacles that limit widespread adoption.
-First, lack of scalability.
-Most methods rely on fine-tuning [51], semi-supervised approaches [45], or extensive LLM queries that grow at least linearly with sample size [15, 10], quickly becoming prohibitive for real-world applications.
-Second, unclear optimization landscape.
-Most methods braid together bespoke heuristics—hard prompts, similarity thresholds, or greedy merges—without a single, well-defined loss function, leaving convergence behaviour opaque and theoretical guarantees out of reach.
-We address these two shortcomings with k-LLMmeans: First, by summarizing each cluster in batches we cap LLM usage independent of dataset size; and second, by inheriting the classic k-means objective we retain a provable descent property and a transparent optimization landscape.
-Regardless of summarization quality, the algorithm will converge to a local minimum.
-Our approach is not intended to outperform complex state-of-the-art LLM-based clustering methods, but rather to provide a robust, scalable and LLM-efficient enhancement to the well-established k-means algorithm, preserving its core behavior while using significantly fewer LLM calls than contemporary LLM-based clustering approaches.
-Fundamentally, a centroid is a numerical abstraction that represents a cluster, and we posit that an LLM-generated textual summary can serve an analogous role.
-Crucially, our approach preserves most of the mathematical properties of k-means, ensuring that core theoretical guarantees—such as convergence behavior, cluster compactness, and complexity— remain intact.
-Figure 1: Illustration of a single-summarization k-LLMmeans algorithm.
-The first panel displays the text embeddings projected onto the first two principal components along with sample texts, with stars indicating initial random centroids.
-The second panel shows the local minimum reached by standard k-means.
-The third panel illustrates the summarization step, where centroids are updated based on LLM-generated summaries of the clustered texts.
-The final panel presents the resulting clusters after the final k-means step, showing convergence to a more optimal k-LLMmeans solution.
-In summary our contributions are as follows:
-  • In Section 3 we introduce k-LLMmeans, which enhances the classical k-means algorithm with LLM-powered centroid updates for robust text clustering; Section 4 then builds on this foundation with a mini-batch variant that brings sequential, scalable, and intrinsically interpretable clustering to streaming text data.
-  • Section 5 presents extensive simulations demonstrating that k-LLMmeans consistently surpasses k-means and other traditional clustering baselines across a wide range of embeddings and LLMs, while obtaining comparable results to recent LLM-based clustering methods with substantially fewer LLM calls.
-  • Section 6 shows that mini-batch k-LLMmeans decisively outperforms streaming variants of k-means, and the case study in Section 7 illustrates how its evolving centroids provide clear, human-readable insight into real-time text streams.
-As a by-product of our experiments, we compile a dataset from StackExchange [41] suitable for benchmarking text-streaming analysis methods (Appendix A.2).
-Figure 2: An illustration of the dynamic evolution of an LLM-generated centroid during the k-LLMmeans clustering process.
+Text clustering is a core problem in natural language processing (NLP), with applications in document organization, topic exploration, and information retrieval [55, 61].
+A standard pipeline embeds documents into vectors [17, 54, 38, 46, 10, 29] and then groups them with a clustering algorithm [47].
+Among these algorithms, k-means [36] remains ubiquitous, iteratively updating each centroid as the mean of its assigned points.
+While effective, this purely numerical averaging can blur contextual nuance present in the original texts [52].
+Prior work has explored alternative centroid definitions and related objectives [27, 9, 31], yet these approaches remain anchored in vector space, which limits interpretability and can induce semantic drift between centroids and their underlying documents.
+This motivates centroids that are explicitly textual, aligning prototypes with human-interpretable summaries.
+Our proposal.
+We introduce a simple modification to k-means: periodically replace numerical centroid updates with summarization steps.
+Rather than averaging embeddings at every iteration, we compute, at spaced iterations, a textual prototype that summarizes each cluster and then re-embed it with the same encoder to obtain the centroid used for assignments.
+These summary-as-centroid updates remain inside the standard k-means loop but capture richer contextual meaning.
+Alternating numerical and summary-based updates yields clusters that are more interpretable and often more semantically coherent.
+Figure 1 illustrates our proposal and shows how even a single summarization step can redirect k-means toward a qualitatively improved solution.
+Figure 1: Illustration of k-NLPmeans/k-LLMmeans with a single summarization step.
+First panel shows the text embeddings with stars marking the initial centroids; second shows the partition reached after k-means iterations (a local minimum); third performs the summarization step, each previous cluster is summarized into a textual prototype and re-embedded; final panel runs one more k-means iteration using these summaries as centroids, yielding a qualitatively improved partition.
+Summarization step.
+We consider two families of summarizers: (1) Classical NLP—centroid-based summarization [48], graph-based methods such as TextRank [37], and LSA-style techniques [14]—which are fast and deterministic (yielding k-NLPmeans); and (2) LLM-based summaries that can capture richer context and filter secondary or noisy content [71, 50, 28] (yielding k-LLMmeans).
+In both variants, summaries are recomputed only at scheduled iterations, so the algorithms share the same summary-as-centroid mechanism and differ solely in the summarizer.
+While methods from (1) and (2) are standard in summarization, our periodic use of summary-derived centroids within the k-means loop is, to our knowledge, novel and central to the method’s interpretability.
+Interpretability and scalability.
+Each summarization step produces a concise, human-readable centroid that exposes evolving cluster semantics, simplifying debugging, validation, and labeling.
+The summarizer is modular: it can be LLM-free (k-NLPmeans) or LLM-assisted under a fixed per-iteration budget that does not grow with dataset size (k-LLMmeans).
+For large-scale and streaming settings, we adapt mini-batch k-means [56] by inserting summarization steps into the mini-batch update rule, enabling efficient online clustering while retaining interpretability.
+Relation to existing LLM-based clustering.
+Recent work shows that LLMs can deliver strong unsupervised clustering [73, 20, 13, 64, 58, 62, 39].
+However, many pipelines face two practical issues: (i) scalability, as they may rely on semi-supervision, iterative labeling, or a number of LLM calls that grows with dataset size; and (ii) opaque optimization, since they often combine prompts, greedy merges, and similarity thresholds without an explicit objective, making convergence behavior hard to analyze.
+Our k-LLMmeans addresses both: by summarizing per cluster at spaced iterations, we cap LLM usage independently of dataset size, and by keeping assignments and numeric updates in embedding space, we preserve the standard k-means objective between summary steps.
+In both k-NLPmeans and k-LLMmeans, if the summaries are weak, the procedure gracefully degrades to vanilla k-means with a suboptimal initialization, retaining its standard local-convergence behavior.
+Note that our method is LLM-optional: k-NLPmeans replaces the summarizer with classical NLP techniques, incurring no LLM usage.
+Contributions.
+In summary, we (i) propose k-NLPmeans and k-LLMmeans, a summary-as-centroid variant of k-means that periodically replaces numeric centroids with textual prototypes re-embedded by the same encoder—an LLM-optional design that, by construction, preserves the standard k-means objective between summary steps; interpretability follows from human-readable centroids and transparent intermediate outputs; (ii) extend the approach to mini-batch k-means for efficient, online clustering of streams; (iii) present a comprehensive empirical study across datasets, embeddings, and summarizers, showing consistent gains over classical baselines and competitiveness with recent LLM-based clustering under a dataset-size–independent LLM budget; and (iv) provide a case study on sequential text streams and release a StackExchange-derived benchmark for evaluating streaming text clustering.
 2 Preliminaries: k-means for text clustering
-We can formalizae the text clustering problem.
-Given a corpus of $n$ text documents $D=\{d_{1},\cdots d_{n}\}$.
-Each document $d_{i}$ is represented as a $d$-dimensional embedding vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$ such that:
+Given a corpus of $n$ text documents $D=\{d_{1},\cdots d_{n}\}$. Each document $d_{i}$ is represented as a $d$-dimensional embedding vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$ such that:
 $\mathbf{x}_{i}=\text{Embedding}(d_{i}).$
@@
 $\min_{C_{1},C_{2},\dots,C_{k}}\sum_{j=1}^{k}\sum_{i\in[C_{j}]}\|\mathbf{x}_{i}-\boldsymbol{\mu}_{j}\|^{2},$ (1)
-where $C_{j}$ denotes the set of embeddings assigned to cluster $j$, $[C_{j}]=\{i|\mathbf{x}_{i}\in C_{j}\}$ denotes the set of embedding indices assigned to cluster $j$ and $\boldsymbol{\mu}_{j}$ is the cluster centroid, computed as the mean of the assigned embeddings:
+where $C_{j}$ denotes the set of embeddings assigned to cluster $j$, $C_{j}]=\{i|\mathbf{x}_{i}\in C_{j}\}$ denotes the set of embedding indices assigned to cluster $j$ and $\boldsymbol{\mu}_{j}$ is the cluster centroid, computed as the mean of the assigned embeddings:
 $\boldsymbol{\mu}_{j}=\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}.$ (2)
-The k-means algorithm assigns each document embedding $\mathbf{x}_{i}$ to the closest centroid based on the smallest distance and updates centroids accordingly until convergence after $T$ iterations.
-The primary objective of the k-means algorithm is to iteratively adjust cluster centroids to minimize the within-cluster variance, effectively guiding the algorithm toward an optimal set of centroids.
+Lloyd’s algorithm [35]—the standard heuristic for k-means—alternates between assigning each document embedding $\mathbf{x}_{i}$ to its nearest centroid and recomputing each centroid as the mean of its assigned points.
+Repeating these two steps for $T$ iterations monotonically reduces the within‑cluster sum of squared distances, steering the procedure toward a (locally) optimal set of centroids.
 However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima [36, 35].
 Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results [5].
-3 k-LLMmeans
-input : $\ D=\{d_{1},\dots,d_{n}\},k,I,m,l,T$
-for $i\leftarrow 1$ to $n$ do
-       $\mathbf{x}_{i}=\text{Embedding}(d_{i})$;
-end for
-for $t\leftarrow 1$ to $T$ do
-       if $t=1$ then
-             // Initialize using k-means++ (omitted if already initialized)
-             $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\texttt{k-means++}(\{d_{1},\dots,d_{n}\},k)$;
-       end if
-      else if $t\bmod l=0$ then
-             // Update with LLM every $l$ iterations (summarization step)
-             for $j\leftarrow 1$ to $k$ do
-                   $m_{j}\leftarrow\min(m,|C_{j}|)$;
-                  $\{d_{z_{1}},\dots,d_{z_{m_{j}}}\}\leftarrow\texttt{k-means++}(\{d_{i}\mid i\in[C_{j}]\},m_{j})$;
-                  $p_{j}\leftarrow\text{Prompt}(I,\{d_{z_{1}},\dots,d_{z_{m_{j}}}\})$;
-                   $s_{j}\leftarrow f_{\text{LLM}}(p_{j})$;
-                   $\boldsymbol{\mu}_{j}\leftarrow\text{Embedding}(s_{j})$;
-             end for
-       end if
-      else
-             // Update centroids using standard averaging (k-means step)
-             for $j\leftarrow 1$ to $k$ do
-                   $\boldsymbol{\mu}_{j}\leftarrow\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}$;
-             end for
-       end if
-      for $j\leftarrow 1$ to $k$ do
-             $C_{j}=\{\}$;
-       end for
-      for $i\leftarrow 1$ to $n$ do
-             $j^{*}\leftarrow\arg\min_{j\in\{1,\dots,k\}}d(x_{i},\boldsymbol{\mu}_{j})$;
-             // Assign $x_{i}$ to cluster $C_{j^{*}}$
-             $C_{j^{*}}\leftarrow C_{j^{*}}\cup\{x_{i}\}$;
-       end for
-end for
-return $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\},\{s_{1},\dots,s_{k}\}$
-Algorithm 1 k-LLMmeans
-To enhance k-means for text clustering, we introduce k-LLMmeans, a novel variant that integrates LLM-based centroids at intermediate stages of the clustering process.
-The formal procedure is outlined in Algorithm 1.
-The key distinction between k-LLMmeans and standard k-means lies in the centroid update mechanism: every $l$ iterations, the traditional update step in Equation (2) is replaced by
-$\boldsymbol{\mu}_{j}=\text{Embedding}(f_{\text{LLM}}(p_{j}))\$ (3)
+3 k-means with summarization steps
+We enhance k-means for text clustering by periodically replacing the numerical centroid update with summarization steps that yield textual-summary–based centroids.
+The procedure (Algorithm 1 in Appendix C) is identical to k-means algorithm except that every $l$ iterations the mean update in (2) is replaced by the embedding of a cluster textual summary.
+At all other iterations, the standard update is used (see Figure 1 for an illustration with one summarization step).
+The summarizer can be instantiated with classical, deterministic methods (k-NLPmeans) or with LLM-based summaries (k-LLMmeans), and is not restricted to the specific techniques we discuss—any textual summarization operator can be dropped in.
+3.1 k-NLPmeans.
+Our first variant uses classical extractive summarization to compute a textual prototype in place of the standard centroid update. Formally, we replace (2) with
+$\boldsymbol{\mu}_{j}=\text{Embedding}\!\left(f^{(q)}_{\text{NLP}}(S_{j})\right),$ (3)
+where $S_{j}$ is the collection (multiset) of sentences obtained by tokenizing all documents assigned to cluster $j$.
+Unless otherwise noted, sentence–sentence similarities are computed as cosine similarity between sentence embeddings produced by the same encoder used for documents.
+The operator $f^{(q)}_{\text{NLP}}$ returns a short summary of $q$ sentences; typical instantiations include:
+  • Centroid-based summarization [48]: compute the centroid of the sentence embeddings for $S_{j}$; rank sentences by cosine similarity to this centroid; concatenate the top $q$ sentences (by rank) to form $f^{(q)}_{\text{NLP}}(S_{j})$.
+  • Graph-based (TextRank) [37]: build a graph whose nodes are sentences in $S_{j}$ with edge weights given by pairwise cosine similarity of their embeddings; run a PageRank-style algorithm to score sentences; select and concatenate the top $q$.
+  • LSA-style SVD in embedding space [14]: stack the sentence embeddings for $S_{j}$, apply singular value decomposition, score sentences by contribution to the leading components, and concatenate the top $q$.
+The resulting summary text $f^{(q)}_{\text{NLP}}(S_{j})$ is then embedded in the same vector space as the documents to yield the new centroid $\boldsymbol{\mu}_{j}$.
+3.2 k-LLMmeans.
+If instead we generate the summary with an LLM, we have k-LLMmeans. Formally, we replace (2) by
+$\boldsymbol{\mu}_{j}=\text{Embedding}(f_{\text{LLM}}(p_{j}))\\$ (4)
 where $p_{j}=\text{Prompt}\left(I,\{d_{z_{i}}|z_{i}\sim[C_{j}]\}_{i=1}^{m_{j}}\right)$, and $m_{j}=\min(m,|C_{j}|)$.
@@
 The instruction $I$ varies depending on the clustering task, but standard summarization prompts are generally sufficient.
-Figure 1 illustrates how k-LLMmeans augments the standard k-means algorithm by introducing few summarization steps between clustering iterations.
-Advantages over k-means.
-This novel approach can help mitigate the tendency of k-means to get stuck in local optima by dynamically adjusting centroids using semantic insights from the data.
-Unlike standard k-means, which relies solely on Euclidean updates, an LLM can refine centroids based on contextual meaning and high-dimensional representations, allowing for better adaptation to complex structures in text data.
-By periodically re-centering clusters with LLM-informed embeddings, the algorithm can escape poor local minima and achieve more coherent and semantically meaningful clusters, even when the initial k-means++ seeding is suboptimal.
-Apart from this modification, our algorithm adheres to the core principles of k-means, preserving its well-established properties and ensuring practical robustness.
-We demonstrate in Section 5.1 that k-LLMmeans consistently outperforms k-means across extensive simulations.
-Advantages over advanced LLM-based clustering methods.
+Figure 1 illustrates how k-LLMmeans with a single summarization step enhances the standard k-means algorithm.
+3.3 Advantages of our approaches
+Over k-means.
+Periodic summarization steps act as a semantic prototype update that can redirect the search trajectory and reduce sensitivity to initialization.
+Instead of relying solely on Euclidean means, a textual summary captures contextual cues present in the underlying documents; re-embedding this summary yields centroids that better reflect cluster semantics.
+In practice, summary-based centroids produce more interpretable and often more semantically coherent—partitions, even when k-means++ seeding is suboptimal.
+Apart from the summary steps, the procedure follows standard k-means: assignments and numeric updates are unchanged and the usual objective is preserved between summary iterations.
+Section 5.1 shows that our methods often outperform vanilla k-means across diverse settings.
+Over advanced LLM-based clustering methods.
 Our approach offers three key advantages over more complex LLM-based clustering methods: (1) Optimization landscape.
-Grounded in the standard k-means objective, k-LLMmeans inherits provable convergence without relying on the fragile heuristics common in newer LLM-driven methods.
-A faulty summary merely returns the algorithm to a regular k-means local optimum, whereas competing approaches hinge on stable, format-specific LLM outputs.
+Grounded in the standard k-means objective, we inherit algorithmic convergence without relying on the fragile heuristics common in newer LLM‑driven methods.
+A poor summary simply leads the procedure back toward a regular k-means local optimum, whereas competing approaches hinge on stable, format‑specific LLM outputs.
 (2) Scalability.
-Unlike most state-of-the-art methods, whose LLM usage complexity grows with sample size [15, 10] or require fine-tuning [51], our algorithm only performs $O(kl)$ LLM calls, with even small $l$ values yielding substancial performance gains.
+Unlike most state-of-the-art methods, whose LLM usage complexity grows with sample size [20, 13] or require fine-tuning [73].
+k-NLPmeans requires no LLM calls and k-LLMmeans only performs $k$ LLM calls per summarization step, with even few summarization steps yielding substantial performance gains.
 (See Section 5.1).
 (3) Interpretability.
-By replacing numerical centroids with LLM-generated textual summaries, k-LLMmeans turns each centroid into a concise, human-readable synopsis of its cluster.
-Practitioners can therefore monitor how cluster semantics evolve across iterations without any post-hoc labeling or visualization.
-While comparable insights can be extracted indirectly from other clustering algorithms, k-LLMmeans embeds this interpretability natively.
-Figure 2 illustrates how these summaries progressively sharpen as the algorithm converges.
-4 Mini-batch k-LLMmeans
-input : $\ \{D_{1},\cdots,D_{b}\},k,I,m,l,T$ // $b$ batches of documents
-for $j\leftarrow 1$ to $k$ do
-       $C_{j}=\{\}$;
-end for
-$\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\{{\bf 0},\dots,{\bf 0}\}$;
-for $i\leftarrow 1$ to $b$ do
-// Compute k-LLMmeans with documents in batch11      1Here k-LLMmeans is initialized with the final centroids of the previous batch
-       $\{\boldsymbol{\mu}^{*}_{1},\dots,\boldsymbol{\mu}^{*}_{k}\},\{C^{*}_{1},\dots,C^{*}_{k}\},S_{b}\leftarrow\texttt{k-LLMmeans}(D_{i},k,I,m,l,T)$;
-       // Update centroids proportional to cluster and batch sizes
-       for $j\leftarrow 1$ to $k$ do
-             $\eta\leftarrow\frac{|C^{*}_{j}|}{|C_{j}|+|C^{*}_{j}|}$;
-             $\boldsymbol{\mu}_{j}\leftarrow\boldsymbol{\mu}_{j}(1-\eta)+\eta\boldsymbol{\mu}^{*}_{j}$;
-       end for
-end for
-return $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\},\{S_{1},\dots,S_{b}\}$
-Algorithm 2 Mini-batch k-LLMmeans
+Replacing numeric centroids with textual summaries turns each prototype into a concise, human-readable synopsis; practitioners can track how cluster semantics evolve over time without post-hoc labeling.
+This transparency extends naturally to our mini-batch variant, enabling real-time monitoring in streaming scenarios (see Figure 2 and Section 5.1).
+4 Mini-batch k-NLPmeans and k-LLMmeans
 Mini-batch k-means [56] is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset.
@@
 Mini-batch k-means exhibits convergence properties comparable to standard k-means while offering superior scalability.
-Although numerous streaming clustering methods that do not rely on LLMs have been studied [40, 2, 35, 3, 1, 31], only a few have incorporated LLMs [43, 28].
-Moreover, existing offline LLM-based clustering approaches face scalability issues, highlighting the need for scalable LLM-driven clustering in an online setting.
-To address this gap, we introduce mini-batch k-LLMmeans, which directly extends mini-batch k-means by incorporating minimal LLM usage during centroid updates.
-Algorithm 2 details how mini-batch k-LLMmeans sequentially receives $b$ batches of documents $D_{1},\dots D_{b}$ where each batch contains a set of documents (these batches can either be random samples from a large corpus or represent sequential data).
-It processes each batch sequentially with k-LLMmeans and updates centroids incrementally using a weighted rule like mini-batch k-means.
-Our mini-batch k-LLMmeans algorithm preserves the desirable properties of mini-batch k-means, with low memory and LLM usage.
+Although numerous streaming clustering methods that do not rely on LLMs have been studied [59, 2, 53, 3, 1, 45], only a few incorporate LLMs [62, 39].
+Moreover, existing offline LLM-based clustering approaches face scalability issues, highlighting the need for scalable summary-based clustering in an online setting.
+To address this, we introduce mini-batch k-NLPmeans and k-LLMmeans, which directly extend mini-batch k-means by inserting summarization steps into the mini-batch update.
+Algorithm 2 details how our approaches sequentially receive $b$ batches of documents $D_{1},\dots D_{b}$ where each batch contains a set of documents (these batches can either be random samples from a large corpus or represent sequential data).
+It processes each batch sequentially with k-NLPmeans/k-LLMmeans and updates centroids incrementally using a weighted rule like mini-batch k-means.
+Our algorithm preserves the desirable properties of mini-batch k-means, with low memory and none or low LLM usage.
 Section  6 shows that it also outperforms in simulations.
@@
 See Appendix A.1 for detailed descriptions of the datasets.
-We evaluate our k-LLMmeans algorithm on each of the four static (non-streaming) dataset using the known number of clusters and performing 120 centroid-update iterations.
-To demonstrate the robustness of our approach, we compute embeddings with four different pretrained models: DistilBERT [36], e5-large [46], S-BERT [34], and text-embedding-3-small [30].
-For the LLM component, we use GPT-3.5-turbo, GPT-4o, Llama-3.3, Claude-3.7 and DeepSeek-V3.
-For the instruction task $I$, we employ a simple summarization prompt that adapted to each dataset.
+We evaluate our algorithms on each of the four datasets using the known number of clusters and performing 120 centroid-update iterations.
+We calculate two variants of our algorithms differing in the number of summarization steps.
+The single variant uses a single summarization step ($l=60$), while the multiple variant performs five summarization steps ($l=20$).
+To demonstrate the robustness of our approach, we compute embeddings with models: DistilBERT[54], e5-large[65], S-BERT[52], and text-embedding-3-small[43].
+For k-NLPmeans we evaluate using TextRank, Centroid and LSA summarization methods mentioned in Section 3.1, with $q=5$.
+For the LLM component of k-LLMmeans, we use GPT-3.5-turbo[44], GPT-4o[26], Llama-3.3[22], Claude-3.7[4] and DeepSeek-V3[34].
+For the instruction task $I$, we employ a simple summarization prompt depending on the task.
 For example, for Bank77 we use the prompt: “The following is a cluster of online banking questions.
-Write a single question that represents the cluster concisely.” We evaluate four variants of k-LLMmeans, differing in the number of summarization steps and the prompt configuration.
-The single variant uses a single summarization step with all cluster documents as input ($l=60$), while the multiple variant performs five summarization steps with shorter prompts ($l=20$), again using all cluster documents each time.
-Additionally, we explore few-shot variants (FS-single and FS-multiple), where each prompt includes $m=10$ randomly selected documents instead of the full cluster.
+Write a single question that represents the cluster concisely.” We explore the effect of prompt size by summarizing using all cluster documents as input, and a few-shot (FS) variant where each prompt includes only $m=10$ randomly selected documents instead of the full cluster.
+For each setting we run five different seeds.
 As traditional clustering baselines, we consider k-means, k-medoids [31], spectral clustering [41], agglomerative clustering [30], and Gaussian Mixture Models (GMM) [15].
@@
 [20] on ClusterLLM [73], IDAS [13], and LLMEdgeRefine [20].
+All k-means based methods are initialized with k-means++.
+We now provide a summary of the results.
 5.1 Results with static data
-Table 1: Average Accuracy (ACC) and Normalized Mutual Information (NMI) for four k-LLMmeans variants using GPT-4o, compared against traditional clustering baselines, with text-embedding-3-small embeddings, evaluated on four benchmark datasets.
+Table 1: Average ACC and NMI for four k-LLMmeans variants using GPT-4o, compared against traditional baselines, with text-embedding-3-small embeddings, evaluated on benchmark datasets.
 Dataset/Method BANK77 CLINC GoEmo Massive (D) Massive (I)
 ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI
+k-NLPmeans
+     TextRank-single 66.4 83.6 78.7 92.4 21.5 20.9 60.6 68.9 53.0 72.5
+     TextRank-multiple 66.8 83.9 79.6 92.8 21.6 21.0 60.5 69.3 53.7 72.8
+     Centroid-single 67.2 84.0 78.5 92.2 22.1 21.2 60.9 68.3 53.5 73.0
+     Centroid-multiple 67.5 84.0 79.0 92.4 21.4 21.1 60.8 69.1 54.3 73.5
+     LSA-single 66.7 83.7 78.8 92.5 21.9 20.5 61.2 68.7 54.1 73.0
+     LSA-multiple 67.1 84.0 80.2 92.9 22.3 20.2 63.3 70.0 55.3 73.4
 k-LLMmeans
@@
      FS-single 67.5 83.8 79.2 92.8 23.0 21.9 60.8 69.4 55.3 73.6
-   FS-multiple 67.9 84.1 80.2 93.1 24.2 22.3 63.2 70.6 56.4 73.9
+     FS-multiple 67.9 84.1 80.2 93.1 24.2 22.3 63.2 70.6 56.7 73.9
 k-means 66.2 83.0 77.3 92.0 20.7 20.5 59.4 67.9 52.9 72.4
@@
 Agglomerative 69.9 83.7 81.0 92.5 15.8 14.1 62.8 67.1 56.6 70.5
-Spectral 68.2 83.3 76.3 90.9 17.6 15.2 61.5 67.0 56.7 71.4
-Table 2: Average ACC, NMI, and distance between final and true centroids ($dist$) for k-means and k-LLMmeans FS-multiple, evaluated on three datasets using four different embedding models.
+Spectral 68.2 83.3 76.3 90.9 17.6 15.2 61.5 67.0 56.6 71.4
+Table 2: Average ACC, NMI, and distdist for k-means and k-LLMmeans FS-multiple, evaluated on three datasets using four different embedding models.
 Dataset CLINC GoEmo Massive (D)
-embedding/Method ACC NMI $dist$ ACC NMI $dist$ ACC NMI $dist$
+embedding/Method ACC NMI distdist ACC NMI distdist ACC NMI distdist
 DistilBERT
-   k-means 53.9 77.2 0.340 17.8 18.3 0.364 44.4 45.1 0.309
+     k-means 53.9 77.2 0.34 17.8 18.3 0.364 44.4 45.1 0.309
+     k-NLPmeans 52.9 77.3 0.353 18.5 18.2 0.365 44.7 45.3 0.311
      k-LLMmeans 55.3 78.7 0.343 18.2 18.8 0.351 46.2 46.4 0.295
 text-embedding-3-small
-   k-means 77.3 92.0 0.200 20.7 20.5 0.287 59.4 67.9 0.246
-   k–LLMmeans 80.2 93.1 0.179 24.2 22.3 0.278 63.2 70.6 0.227
+     k-means 77.3 92.0 0.2 20.7 20.5 0.287 59.4 67.9 0.246
+     k-NLPmeans 80.0 92.9 0.173 22.3 20.2 0.29 63.3 70.0 0.227
+     k-LLMmeans 80.2 93.1 0.179 24.2 22.3 0.278 63.2 70.6 0.227
 e5-large
      k-means 73.8 90.8 0.131 22.8 22.8 0.176 58.4 63.7 0.138
+     k-NLPmeans 76.5 92.0 0.119 22.6 22.6 0.179 60.4 65.7 0.137
      k-LLMmeans 77.2 92.5 0.119 24.2 24.3 0.168 62.3 65.9 0.133
@@
      k-means 76.9 91.0 0.215 13.7 13.3 0.355 58.2 64.6 0.271
+     k-NLPmeans 79.0 91.9 0.201 14.1 12.8 0.363 58.5 64.6 0.272
      k-LLMmeans 79.7 92.5 0.198 14.7 13.9 0.346 59.7 65.6 0.254
@@
 prompts ACC NMI prompts ACC NMI prompts ACC NMI
+k-NLPmeans 0 76.5 92.0 0 22.6 22.6 0 60.4 65.7
 k-LLMmeans
-   GPT-3.5 750 76.0 92.1 135 23.9 24.1 90 63.3 66.2
+     GPT-3.5-turbo 750 76.0 92.1 135 23.9 24.1 90 63.3 66.2
      GPT-4o 750 77.2 92.5 135 24.2 24.3 90 62.3 65.9
@@
 LLMEdgeRefine 1350 86.8 94.9 895 34.8 29.7 892 63.1 68.7
-Comparing k-LLMmeans variants and traditional clustering methods.
-Table 1 reports Average Accuracy (ACC) and Normalized Mutual Information (NMI) for four k-LLMmeans variants using GPT-4o, compared against traditional clustering baselines, with text-embedding-3-small embeddings, evaluated on four benchmark datasets.
-Overall, all k-LLMmeans variants outperform traditional clustering methods in terms of NMI and in general achieve higher average accuracy across datasets.
-Within k-LLMmeans variants there are minor differences, although running multiple summarization steps tends to offer better performance.
-Interestingly, using few-shot summarization generally leads to better results than prompting on the entire cluster.
-These observations suggest k-LLMmeans improves over k-means while remaining efficient, as few-shot summarization appears sufficient without needing extensive prompt length (See Appendix B for standard errors of all simulations).
-[…]
-Comparing k-LLMmeans with different LLMs and state-of-the-art LLM-based clustering methods.
-Table 2 reports number of LLM calls (prompts) and Average Accuracy (ACC) and Normalized Mutual Information (NMI) for the k-LLMmeans FS-multiple variant on three benchmark datasets, pairing e5-large embeddings with several backbone LLMs.
-Performance is stable across LLM choices, indicating that k-LLMmeans is robust to the specific language model.
-For reference, Table 2 also reports results from Feng et al.
-[15] for ClusterLLM, IDAS, and LLMEdgeRefine.
-All three baselines rely on GPT-3.5: ClusterLLM (with fine-tuning) and IDAS use the same e5-large embeddings as we do, whereas LLMEdgeRefine benefits from the stronger Instructor embeddings [21].
-Even so, our GPT-3.5 configuration attains ACC and NMI that are only marginally lower—and often comparable—while requiring orders of magnitude fewer LLM calls and no fine-tuning.
-This substantial reduction in model prompts translates directly into improved scalability and cost-efficiency.
-6 Sequential (mini-batch) experiments
-We use yearly posts from 2020 to 2023 from 35 StackExchange sites [41].
-Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods (See Appendix A.2 for details).
-For each yearly subset, we split the data into $b=\lceil\tfrac{n}{10000}\rceil$ equal-sized batches ${D_{1},\dots,D_{b}}$ in chronological order, where $n$ is the number of documents for that year.
-We then use the ground-truth clusters and run the mini-batch k-LLMmeans algorithm in its four variants described in Section 5 (for practical reasons for the full cluster method we use 50 samples).
-For comparison, we evaluate three baselines: mini-batch k-means with standard random sampling across the entire year (until convergence), sequential mini-batch k-means with $b$ chronological batches, and standard k-means on the full dataset.
-In all experiments, we use text-embedding-3-small and GPT-4o for the summarization step.
-6.1 Results with streaming data
 Table 4: Average ACC, and average NMI for four sequential mini-batch variants, k-means, mini-batch k-means, sequential mini-batch k-means on the yearly StackExchange data.
@@
 ACC NMI ACC NMI ACC NMI ACC NMI
+mini-batch k-NLPmeans 68.0 79.5 67.9 78.5 69.0 78.9 71.6 78.8
 mini-batch k-LLMmeans
-   single 73.1 81.2 73.2 80.3 70.8 79.4 72.3 80.0
      multiple 72.5 80.9 73.6 80.5 74.4 80.5 73.4 80.3
-   FS + single 74.0 81.1 72.2 79.8 70.2 79.3 71.6 79.8
      FS + multiple 75.4 81.6 73.5 80.2 72.8 80.1 72.7 80.1
@@
 seq. mini-batch k-means 67.0 76.6 66.7 75.2 65.6 75.6 65.8 74.8
-Table 4 presents ACC and NMI on the yearly StackExchange corpus, comparing four sequential mini-batch variants of k-LLMmeans against standard k-means, mini-batch k-means, and sequential mini-batch k-means across five random seeds.
-The mini-batch k-LLMmeans methods consistently outperform all three baselines across all scenarios.
-Notably, they even surpass standard k-means, which operates on the full dataset rather than mini-batches.
-This highlights the scalability and efficiency of our approach for sequential data, achieving superior performance despite processing data sequentially.
-As observed previously, few-shot summarization proves effective, mitigating the need for extensive LLM usage.
+Comparing summarization variants and traditional clustering methods.
+Table 1 reports mean accuracy (ACC) and normalized mutual information (NMI) for multiple k-NLPmeans and k-LLMmeans variants (LLM: GPT-4o), using text-embedding-3-small embeddings on four datasets.
+Across all datasets, our methods outperform traditional baselines on NMI and generally achieve higher ACC.
+Within k-NLPmeans, differences are modest, with LSA yielding the best overall scores among extractive summarizers.
+k-LLMmeans achieves the strongest results, attributable to higher-quality abstractive summaries; interestingly few-shot variants show better performance.
+Multiple summarization steps tend to offer additional gains over a single step.
+Overall, both k-NLPmeans and k-LLMmeans improve upon traditional algorithms, with k-LLMmeans delivering the best results, and remaining efficient, as few-shot summarization appears sufficient without needing extensive prompt length.
+Comparing our approaches with k-means using different embeddings.
+Table 2 compares k-NLPmeans (LSA-multiple) and k-LLMmeans (FS‑multiple) with k-means with different embeddings on three benchmark datasets, reporting average ACC, average NMI, and average Euclidean distance between the learned and ground‑truth centroids ($\mathit{dist}$).
+The $\mathit{dist}$ metric directly gauges how closely each algorithm recovers the true centroids, an especially meaningful criterion for centroid‑based methods.
+Across all tested embeddings, our approaches achieve higher average ACC and NMI than k-means producing smaller average $\mathit{dist}$ values.
+Interestingly k-LLMmeans stands as the best, demonstrating that its LLM‑guided centroid updates converge to solutions that are both more accurate and more faithfully aligned with the underlying cluster structure.
+Comparing our approaches with different LLMs and state‑of‑the‑art LLM‑based clustering methods.
+Table 3 reports the number of LLM calls (prompts) and mean ACC/NMI for k-NLPmeans (LSA-multiple) and k-LLMmeans (FS-multiple) on three benchmarks with e5-large embeddings.
+For k-LLMmeans we vary the backbone generator—including GPT-3.5—and observe stable performance across LLMs, indicating robustness to the specific model.
+By design, k-NLPmeans uses zero LLM calls.
+For context, we also include results from Feng et al.
+[20]: ClusterLLM (fine-tuned) and IDAS use GPT-3.5 with the same e5-large embeddings, while LLMEdgeRefine uses the stronger Instructor embeddings [29].
+Our GPT-3.5 configuration of k-LLMmeans achieves comparable, sometimes slightly lower ACC/NMI, but with much fewer LLM calls (independent of dataset size) and no fine-tuning.
+k-NLPmeans trails k-LLMmeans slightly but remains competitive without any LLM usage.
+Overall, our framework offers a favorable quality–cost trade-off: k-NLPmeans for zero-inference-cost deployments, and k-LLMmeans for higher accuracy under a small, fixed summarization budget.
+Results with standard deviations are reported in Tables 5, 6, and 7 of Appendix B.1.
+We also assess sensitivity to the instruction prompt II in k-LLMmeans and to parameter $q$ in k-NLPmeans.
+As shown in Table 9 (Appendix B.2), performance is generally stable across prompt choices and $q$ values.
+6 Sequential (mini-batch) experiments
+We use yearly posts from 2020 to 2023 from 35 StackExchange sites [60].
+Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods.
+We release this dataset alongside our submission (see Appendix A.2 for details).
+For each yearly subset, we split the data into $b=\lceil\tfrac{n}{10000}\rceil$ equal-sized batches $D_{1},\dots,D_{b}$} in chronological order, where $n$ is the number of documents for that year.
+We run the mini-batch k-LLMmeans algorithm in its four variants described in Section 5 (for practical reasons we set $m=50$ for the full cluster variant).
+We compare with three baselines: mini-batch k-means with standard random sampling, sequential mini-batch k-means with $b$ chronological batches, and standard k-means on the full dataset.
+We use ground-truth clusters, text-embedding-3-small for embeddings, GPT-4o for the summarization step, and five different seeds.
+6.1 Results with streaming data
+Table 4 reports average ACC and NMI on the yearly StackExchange corpus, comparing mini‑batch k-NLPmeans (LSA-multiple) and mini‑batch k-LLMmeans variants against standard k-means, mini‑batch k-means, and sequential mini‑batch k-means.
+Mini-batch k-LLMmeans variants achieve the highest ACC and NMI—consistently outperforming baselines, and even surpassing full-dataset k-means.
+Despite operating in a streaming regime, our approach clusters the entire corpus of 205,943 posts using no more than 3,850 LLM calls, demonstrating a favorable accuracy-per-LLM-token trade-off.
+The few-shot summarization design keeps prompts short, sidestepping context-window limits while preserving cluster quality.
+Mini-batch k-NLPmeans also improves over all mini-batch baselines, though it trails the k-LLMmeans variants—likely because its extractive summaries are more sensitive to the noisiness and heterogeneity of community posts than LLM-generated centroids.
+Overall, these results highlight that our interpretable, mini-batch formulation scales to long sequential streams with strong accuracy and limited LLM usage.
+Results with standard deviations are reported in Table 8 (Appendix B.1).
 7 Case study
@@
 Write a single question that represents the cluster”.
-Clustering results.
-The resulting LLM-based centroids span key areas in AI and ML, including neural network training optimization, computer vision tasks, and broader topics such as small datasets, class imbalance, and interpretability.
-Other clusters cover advanced themes like symbolic-neural integration for AGI, deep reinforcement learning, theoretical debates (e.g., Bayesian vs.
-frequentist methods), as well as NLP and various architectural choices.
 Interpretation.
-The key insight from our algorithm is the evolution of clusters over batches, reflecting their dynamic nature over time.
-Figure 3 illustrates this progression for four major themes: Image Model Optimization, AI Evolution and Challenges, AI Mathematics, and Advanced NLP Techniques.
-The Image Model Optimization cluster refines its focus from broad improvements in image classification and object detection to specific deep learning optimizations, such as facial expression recognition and anomaly detection.
-Over time, it emphasizes practical challenges, including handling imbalanced datasets, model uncertainty, and performance enhancement through transfer learning and specialized loss functions.
-The AI Evolution and Challenges cluster transitions from early rule-based AI and symbolic systems to advanced deep learning and generative models, highlighting both progress and limitations.
-The shift toward hybrid and neuro-symbolic AI reflects the growing need to integrate diverse techniques for AGI, balancing efficiency, adaptability, and safety.
-The AI Mathematics cluster starts with unconventional symbol usage and parameter definitions in machine learning, later expanding to matrix rank analysis, Bayesian and frequentist methods, adversarial attacks, and the mathematical foundations necessary for advanced AI concepts.
-The Advanced NLP Techniques cluster progresses from foundational models (e.g., Word2Vec, BERT) and basic tasks (e.g., sentiment analysis, text classification) to more complex challenges, including model evaluation, domain-specific vocabulary, and multilingual alignment.
-The shift highlights an increasing focus on contextual understanding, transfer learning, and alternative evaluation metrics for robust NLP.
-The result of this case study could enhance post categorization, searchability, answer relevance, and trend detection, making AI discussions more efficient and insightful.
-More broadly, this demonstrates how our interpretable mini-batch k-LLMmeans clustering algorithm, applied to sequential text streaming, can help practitioners track topic evolution, improve decision-making, and enhance transparency in dynamic information flows.
-Figure 3: Sequential evolution of the LLM-generated centroids for four primary clusters during the three batches of the sequential mini-batch k-LLMmeans process applied to 2021 posts from the AI Stack Exchange site [41].
-Main aspects are manually highlighted in different color on each cluster.
+Our interpretable clustering enables temporal analysis of topic change, since each centroid is a human-readable summary rather than a latent vector.
+Figure 2 shows this for three major clusters: Image Model Optimization, AI Evolution and Challenges, and Advanced NLP Techniques.
+Across 2021, these clusters drift from fundamentals toward integration-at-scale, mirroring how tools, models, and deployment pressures matured.
+Image Model Optimization begins with robustness anxieties (adversaries, occlusion, scale) around mainstream detectors/segmenters, pivots mid-year to “how do I implement this well?” (transfer learning, augmentation, preprocessing), and ends with production hardening (class imbalance, specialized losses, edge efficiency) as strong pretrained models and libraries made setup routine and pushed bottlenecks to data and deployment [8, 32].
+AI Evolution and Challenges moves from historical “what is AI becoming?” to curiosity about large generative/multimodal systems—probably sparked by public releases like CLIP/DALL·E and their failure modes—and finishes with system-level synthesis (neuro-symbolic integration, RL control, compute efficiency, safety) as enterprise adoption and policy attention grow [49, 51, 40, 42, 19].
+Advanced NLP Techniques tracks the transformer boom: early questions focus on which models and preprocessing to use; mid-year attention shifts to measuring semantics beyond BLEU (e.g., BERTScore, BLEURT) as long-context and domain-shift issues surface; late-year concerns are product-driven—sequence limits, domain vocab, multilingual/cross-modal alignment—reflecting rapid uptake of pretrained encoders/decoders in industry [10, 72, 57, 68].
+Practically, these trajectories can power better post categorization, searchability, answer routing, and trend detection.
+More broadly, they showcase how our mini-batch k-LLMmeans exposes readable centroids enabling end-to-end interpretability for sequential text streams, letting practitioners track topic evolution, attribute causes, and make transparent, timely decisions in dynamic corpora.
+Figure 2: Sequential evolution of the LLM-generated centroids for three primary clusters during the three batches of the sequential mini-batch k-LLMmeans process applied to 2021 posts from the AI Stack Exchange site [60].
+Main aspects are manually highlighted.
 8 Related work
@@
 Additionally, topic modeling methods, from probabilistic latent semantic analysis [24] to latent Dirichlet allocation [7], capture word co-occurrence patterns and latent topics.
+BERTopic [23] offers an embedding–reduction–density pipeline for interpretable topics, while DeepCluster [11] uses clustering to iteratively train encoders in a self-supervised manner (orthogonal to our fixed-embedding setting).
 LLM-based clustering.
@@
 9 Discussion
-We introduced k-LLMmeans and mini-batch k-LLMmeans, unsupervised text clustering algorithms that refine k-means by integrating a lightweight LLM-based summarization into centroid updates.
-This modification enriches the contextual representation while preserving k-means’ stability and scalability, yielding interpretable clusters for both static corpus of documents and streaming text data, with low LLM usage.
+We presented k-NLPmeans and k-LLMmeans, which replace the numeric centroid in k-means with textual summaries, yielding interpretable prototypes while preserving scalability.
+The LLM-free k-NLPmeans achieves competitive accuracy with minimal additional complexity; k-LLMmeans augments clusters with LLM summaries under a fixed, iteration-bounded LLM budget, offering a favorable accuracy–efficiency trade-off without fine-tuning.
+A mini-batch extension enables streaming operation.
+Overall, summary-as-centroid is a simple yet powerful and novel modification that unifies interpretability and efficiency and broadens the applicability of k-means to modern text streams.
+Beyond text, summary-as-centroid could be generalized by replacing numeric centroids with re-embedded, domain-specific prototypes (e.g., image collages, synthetic utterances, DNA consensus motifs, representative subgraphs), preserving interpretability without altering the clustering.
 Limitations.
-Our approach depends on pre-computed embeddings and LLM-generated summaries, so any bias or error in the underlying language models can propagate to the final clusters.
-While we demonstrate that a simple instruction-based approach is effective, the design of these instructions can still influence outcomes.
-Additionally, the few-shot variant of our algorithm assumes that a small number of samples can sufficiently capture the overall structure of the clusters.
-While this is generally practical, it may become a limitation when dealing with highly complex or heterogeneous clusters.
+Our k-LLMmeans relies on LLM-generated summaries; thus, biases or errors in the underlying model can propagate to cluster prototypes and downstream analyses.
+Although simple instruction prompts are effective (Appendix B.2), prompt design can still influence outcomes.
+The few-shot variant assumes that a small set of representative samples captures cluster structure—practical in many settings but potentially limiting for heterogeneous clusters.
+Regarding potential pretraining exposure, note that the LLM in our pipeline is used solely to summarize texts already assigned to each cluster; it does not inject external labels or unseen content, so any exposure does not confer an unfair advantage.
+As indicated by suboptimal results on the StackExchange dataset (Table 4), k-NLPmeans may benefit from applying text-cleaning preprocessing prior to use.
+Cost.
+Our k-NLPmeans is LLM-free; runtime is dominated by standard k-means and a lightweight summarization step.
+In contrast, k-LLMmeans incurs one LLM call per cluster per summarization step.
+These calls are highly parallelizable, so wall-clock time is roughly the latency of a single LLM call multiplied by the number of summarization steps.
+In practice, running any static dataset with five summarization rounds using GPT-4o and text-embedding-3-small cost $\mathdollar 1$ and completed in $\approx 1$ minute on a single laptop without parallelization; for the Stack Overflow dataset, the end-to-end run cost $\mathdollar 2.5$ and took $\approx 8$ minutes under the same conditions.
+By comparison, competing LLM-heavy approaches would require $\mathdollar 18\text{--}\mathdollar 25$ and >40>40 minutes on identical hardware and API settings.
+Thus, k-LLMmeans remains interpretable and cost-efficient among LLM-assisted methods, while k-NLPmeans provides a strong, interpretable, zero-LLM alternative when budget or latency is constrained.
+Reproducibility Statement
+We release code to reproduce all results in this paper at https://github.com/jairoadiazr/summaryCentroids (see the README for instructions).
 Appendix A Datasets
@@
 Stack Exchange content is licensed under the Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) license.
-Appendix B Extended simulation results
-B.1 Static experiments
-Tables 59 present extended results from the simulations described in Section 5.
-For each dataset, we report the average Accuracy (ACC), Normalized Mutual Information (NMI), and distance between final and true centroids ($dist$), along with standard deviations computed over 10 random seeds.
-Results are provided for all k-LLMmeans variants using GPT-4o, as well as for traditional clustering algorithms, evaluated under four different embedding types.
-Table 10 show extended results for all benchmark datasets and multiple LLMs.
-B.2 Streaming (mini-batch) experiments
-Table 11 report the standard deviation for results presented on Table 4.
-Appendix C Computational resources for experiments
-Our simulations span approximately one week on a standard laptop without specialized hardware, with the primary computational bottleneck arising from LLM usage.
-We conducted API-based experiments using GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3.
-Most experiments relied on OpenAI’s API (GPT-4o and text-embedding-3-small), incurring a cost of roughly \$700 and consuming around 120 million tokens.
-Experiments involving the other LLMs collectively cost under \$30.
-Appendix D Broader Impacts
-While k-LLMmeans is primarily a foundational method aimed at improving scalability and interpretability in LLM-based text clustering, potential societal risks arise from the misuse of generated cluster summaries.
-For example, if used in sensitive domains such as legal, medical, or financial text analysis, misleading or biased summaries could propagate incorrect conclusions.
-Additionally, because the method relies on pretrained language models, any biases or factual inaccuracies embedded in those models may influence clustering outcomes.
-To mitigate these risks, we recommend cautious deployment in high-stakes settings, transparency about the LLMs used, and regular human-in-the-loop validation when interpretability affects downstream decisions.
-Table 5: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids (dist) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on Bank77 using four different embedding models (standard deviations in parenthesis).
-embedding/Method ACC NMI $dist$
-DistilBERT
-k-LLMmeans
-   single 43.8(1.12) 64.8(0.493) 0.351(0.013)
-   multiple 43.5(0.991) 64.8(0.516) 0.358(0.006)
-   FS-single 43.7(0.75) 64.8(0.48) 0.352(0.007)
-   FS-multiple 43.8(1.032) 64.9(0.362) 0.363(0.011)
-k-means 43.653(0.714) 64.389(0.456) 0.335(0.007)
-k-medoids 32.89(0.0) 55.991(0.0) 0.591(0.0)
-GMM 44.701(1.163) 64.481(0.63) 0.318(0.005)
-Agglomerative 45.974(0.0) 67.064(0.0) 0.33(0.0)
-Spectral 45.925(0.201) 66.057(0.241) 0.365(0.001)
-text-embedding-3-small
-k-LLMmeans
-   single 67.1(1.216) 83.6(0.459) 0.221(0.012)
-   multiple 66.7(0.778) 83.6(0.2) 0.226(0.009)
-   FS-single 67.5(1.072) 83.8(0.358) 0.219(0.011)
-   FS-multiple 67.9(1.546) 84.1(0.412) 0.22(0.014)
-k-means 66.25(1.815) 83.005(0.559) 0.225(0.014)
-k-medoids 41.688(0.0) 69.457(0.0) 0.639(0.0)
-GMM 68.792(1.784) 83.048(0.543) 0.198(0.013)
-Agglomerative 69.87(0.0) 83.741(0.0) 0.2(0.0)
-Spectral 68.562(0.562) 83.746(0.38) 0.227(0.003)
-e5-large
-k-LLMmeans
-   single 61.1(1.041) 78.6(0.412) 0.138(0.006)
-   multiple 61.6(0.783) 79.0(0.355) 0.136(0.005)
-   FS-single 61.5(1.218) 78.8(0.484) 0.136(0.004)
-   FS-multiple 62.7(1.706) 79.5(0.646) 0.134(0.007)
-k-means 58.097(1.732) 76.703(0.795) 0.142(0.007)
-k-medoids 37.175(0.0) 64.487(0.0) 0.331(0.0)
-GMM 56.977(1.384) 75.801(0.502) 0.142(0.006)
-Agglomerative 61.623(0.0) 79.259(0.0) 0.138(0.0)
-Spectral 63.195(0.69) 79.684(0.312) 0.144(0.003)
-S-BERT
-k-LLMmeans
-   single 64.4(1.83) 81.7(0.563) 0.253(0.014)
-   multiple 64.6(1.393) 81.8(0.389) 0.253(0.013)
-   FS-single 65.4(1.693) 82.0(0.515) 0.246(0.013)
-   FS-multiple 65.7(1.841) 82.2(0.648) 0.247(0.013)
-k-means 63.643(2.272) 80.937(0.744) 0.255(0.016)
-k-medoids 41.299(0.0) 68.985(0.0) 0.64(0.0)
-GMM 63.565(2.014) 80.504(0.579) 0.248(0.017)
-Agglomerative 64.643(0.0) 81.601(0.0) 0.255(0.0)
-Spectral 65.737(0.652) 83.247(0.395) 0.268(0.006)
-Table 6: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids ($dist$) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on CLINC using four different embedding models (standard deviations in parenthesis).
-embedding/Method ACC NMI $dist$
-DistilBERT
-k-LLMmeans
-   single 53.3(1.468) 77.4(0.547) 0.352(0.01)
-   multiple 53.5(1.141) 77.7(0.437) 0.356(0.01)
-   FS-single 54.7(1.117) 78.1(0.552) 0.346(0.008)
-   FS-multiple 55.3(1.104) 78.7(0.531) 0.343(0.009)
-k-means 53.889(0.956) 77.188(0.494) 0.34(0.007)
-k-medoids 36.889(0.0) 66.765(0.0) 0.69(0.0)
-GMM 53.42(1.362) 76.534(0.442) 0.334(0.01)
-Agglomerative 55.333(0.0) 78.805(0.0) 0.343(0.0)
-Spectral 52.289(0.395) 76.215(0.268) 0.364(0.004)
-text-embedding-3-small
-k-LLMmeans
-   single 78.1(1.67) 92.5(0.474) 0.194(0.015)
-   multiple 79.1(1.675) 92.8(0.48) 0.186(0.015)
-   FS-single 79.2(1.725) 92.8(0.427) 0.186(0.015)
-   FS-multiple 80.2(1.523) 93.1(0.323) 0.179(0.015)
-k-means 77.253(1.828) 91.965(0.351) 0.2(0.014)
-k-medoids 49.289(0.0) 77.718(0.0) 0.694(0.0)
-GMM 80.484(1.181) 92.492(0.222) 0.168(0.011)
-Agglomerative 81.0(0.0) 92.544(0.0) 0.169(0.0)
-Spectral 76.02(0.401) 90.812(0.139) 0.214(0.003)
-e5-large
-k-LLMmeans
-   single 75.2(1.917) 91.6(0.48) 0.127(0.008)
-   multiple 75.4(1.858) 92.0(0.472) 0.127(0.009)
-   FS-single 75.4(1.746) 91.7(0.46) 0.127(0.008)
-   FS-multiple 77.2(1.604) 92.5(0.399) 0.119(0.007)
-k-means 73.831(2.338) 90.77(0.739) 0.131(0.01)
-k-medoids 41.911(0.0) 73.371(0.0) 0.371(0.0)
-GMM 77.647(1.286) 91.41(0.279) 0.106(0.006)
-Agglomerative 80.333(0.0) 92.544(0.0) 0.101(0.0)
-Spectral 72.853(0.252) 89.828(0.092) 0.128(0.002)
-S-BERT
-k-LLMmeans
-   single 77.3(1.475) 91.6(0.381) 0.215(0.013)
-   multiple 77.9(1.373) 91.9(0.436) 0.21(0.012)
-   FS-single 78.0(1.115) 91.8(0.332) 0.208(0.01)
-   FS-multiple 79.7(0.734) 92.5(0.247) 0.198(0.008)
-k-means 76.853(1.63) 91.036(0.496) 0.215(0.013)
-k-medoids 51.156(0.0) 78.506(0.0) 0.68(0.0)
-GMM 77.333(0.93) 91.025(0.254) 0.203(0.009)
-Agglomerative 82.622(0.0) 92.78(0.0) 0.158(0.0)
-Spectral 77.167(0.36) 90.794(0.105) 0.215(0.002)
-Table 7: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids ($dist$) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on GoEmo using four different embedding models (standard deviations in parenthesis).
-embedding/Method ACC NMI $dist$
-DistilBERT
-k-LLMmeans
-   single 17.5(0.467) 18.3(0.447) 0.358(0.009)
-   multiple 17.9(0.861) 18.3(0.62) 0.354(0.006)
-   FS-single 18.0(0.681) 18.9(0.407) 0.357(0.006)
-   FS-multiple 18.2(0.52) 18.8(0.733) 0.351(0.004)
-k-means 17.755(0.857) 18.267(0.667) 0.364(0.005)
-k-medoids 15.851(0.0) 13.964(0.0) 0.769(0.0)
-GMM 18.07(0.819) 18.729(0.471) 0.36(0.005)
-Agglomerative 16.857(0.0) 15.402(0.0) 0.352(0.0)
-Spectral 17.198(0.21) 17.23(0.189) 0.42(0.002)
-text-embedding-3-small
-k-LLMmeans
-   single 24.0(1.346) 22.3(0.572) 0.278(0.007)
-   multiple 24.1(1.326) 22.1(0.651) 0.275(0.004)
-   FS-single 23.0(0.937) 21.9(0.638) 0.283(0.011)
-   FS-multiple 24.2(1.156) 22.3(0.67) 0.278(0.011)
-k-means 20.734(0.997) 20.466(0.672) 0.287(0.005)
-k-medoids 15.65(0.0) 15.943(0.0) 0.852(0.0)
-GMM 20.831(1.418) 20.123(0.524) 0.296(0.007)
-Agglomerative 15.818(0.0) 14.098(0.0) 0.286(0.0)
-Spectral 17.661(0.328) 15.045(0.319) 0.341(0.004)
-e5-large
-k-LLMmeans
-   single 23.4(0.875) 23.4(0.694) 0.176(0.007)
-   multiple 23.4(0.9) 23.7(0.521) 0.174(0.004)
-   FS-single 23.4(1.089) 23.7(0.356) 0.174(0.003)
-   FS-multiple 24.2(0.761) 24.3(0.702) 0.168(0.003)
-k-means 22.778(0.917) 22.761(0.729) 0.176(0.003)
-k-medoids 14.812(0.0) 15.667(0.0) 0.447(0.0)
-GMM 21.367(0.986) 22.197(0.69) 0.181(0.003)
-Agglomerative 19.571(0.0) 18.544(0.0) 0.177(0.0)
-Spectral 23.388(0.524) 21.362(0.257) 0.192(0.001)
-S-BERT
-k-LLMmeans
-   single 14.4(0.739) 13.6(0.547) 0.351(0.006)
-   multiple 14.4(0.75) 13.7(0.59) 0.348(0.005)
-   FS-single 14.5(0.79) 13.9(0.413) 0.348(0.006)
-   FS-multiple 14.7(0.551) 13.9(0.501) 0.346(0.004)
-k-means 13.737(0.617) 13.269(0.574) 0.355(0.005)
-k-medoids 13.941(0.0) 11.974(0.0) 0.899(0.0)
-GMM 14.31(0.744) 13.628(0.482) 0.353(0.003)
-Agglomerative 12.399(0.0) 10.766(0.0) 0.355(0.0)
-Spectral 13.05(0.211) 11.777(0.123) 0.416(0.003)
-Table 8: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids ($dist$) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on MASSIVE (Domain) using four different embedding models (standard deviations in parenthesis).
-embedding/Method ACC NMI $dist$
-DistilBERT
-k-LLMmeans
-   single 45.8(1.508) 46.0(1.164) 0.294(0.016)
-   multiple 45.2(2.264) 45.6(1.199) 0.306(0.017)
-   FS-single 45.3(1.789) 45.7(1.099) 0.299(0.021)
-   FS-multiple 46.2(1.348) 46.4(1.119) 0.295(0.015)
-k-means 44.368(1.528) 45.065(0.988) 0.309(0.021)
-k-medoids 33.457(0.0) 27.223(0.0) 0.81(0.0)
-GMM 45.451(1.892) 46.966(0.74) 0.293(0.02)
-Agglomerative 35.777(0.0) 39.931(0.0) 0.355(0.0)
-Spectral 40.212(2.693) 44.476(1.514) 0.473(0.009)
-text-embedding-3-small
-k-LLMmeans
-   single 61.1(3.393) 69.6(1.604) 0.232(0.018)
-   multiple 60.6(2.822) 69.5(1.661) 0.238(0.016)
-   FS-single 60.8(3.206) 69.4(1.091) 0.233(0.02)
-   FS-multiple 63.2(2.843) 70.6(1.379) 0.227(0.022)
-k-means 59.391(4.026) 67.869(1.671) 0.246(0.015)
-k-medoids 37.559(0.0) 38.788(0.0) 0.797(0.0)
-GMM 59.391(2.883) 68.0(1.01) 0.253(0.014)
-Agglomerative 62.845(0.0) 67.07(0.0) 0.222(0.0)
-Spectral 61.51(0.022) 66.987(0.023) 0.331(0.0)
-e5-large
-k-LLMmeans
-   single 60.6(2.856) 64.8(1.611) 0.134(0.014)
-   multiple 61.5(4.447) 66.0(2.105) 0.131(0.016)
-   FS-single 61.0(3.441) 64.9(1.42) 0.136(0.019)
-   FS-multiple 62.3(1.358) 65.9(0.776) 0.133(0.007)
-k-means 58.426(3.889) 63.663(1.748) 0.138(0.02)
-k-medoids 41.19(0.0) 36.343(0.0) 0.405(0.0)
-GMM 63.073(3.473) 66.018(1.588) 0.132(0.008)
-Agglomerative 66.073(0.0) 66.279(0.0) 0.132(0.0)
-Spectral 57.001(0.02) 63.327(0.018) 0.202(0.0)
-S-BERT
-k-LLMmeans
-   single 61.0(3.797) 65.6(1.36) 0.248(0.026)
-   multiple 60.1(2.016) 65.2(0.735) 0.25(0.019)
-   FS-single 61.7(2.529) 66.2(1.134) 0.243(0.026)
-   FS-multiple 59.7(2.543) 65.6(0.795) 0.254(0.019)
-k-means 58.248(2.029) 64.608(1.09) 0.271(0.023)
-k-medoids 45.763(0.0) 44.605(0.0) 0.8(0.0)
-GMM 57.935(3.645) 64.404(1.094) 0.27(0.032)
-Agglomerative 63.416(0.0) 64.72(0.0) 0.222(0.0)
-Spectral 58.487(0.85) 63.148(0.49) 0.367(0.005)
-Table 9: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids ($dist$) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on MASSIVE (Intention) using four different embedding models (standard deviations in parenthesis).
-embedding/Method ACC NMI $dist$
-DistilBERT
-k-LLMmeans
-   single 44.1(1.099) 58.6(0.511) 0.362(0.014)
-   multiple 43.8(1.494) 59.0(0.46) 0.363(0.013)
-   FS-single 43.5(1.196) 58.8(0.59) 0.365(0.01)
-   FS-multiple 43.6(1.343) 59.0(0.594) 0.362(0.012)
-k-means 43.124(1.465) 58.102(0.516) 0.366(0.014)
-k-medoids 25.084(0.0) 42.582(0.0) 0.847(0.0)
-GMM 41.379(1.249) 57.375(0.452) 0.367(0.013)
-Agglomerative 39.139(0.0) 55.378(0.0) 0.386(0.0)
-Spectral 46.022(0.482) 58.237(0.234) 0.378(0.004)
-text-embedding-3-small
-k-LLMmeans
-   single 54.0(1.513) 73.0(0.582) 0.314(0.013)
-   multiple 55.8(1.396) 73.5(0.762) 0.307(0.013)
-   FS-single 55.3(1.759) 73.6(0.774) 0.314(0.014)
-   FS-multiple 56.4(1.32) 73.9(0.7) 0.302(0.012)
-k-means 52.878(1.072) 72.373(0.541) 0.32(0.011)
-k-medoids 35.911(0.0) 52.442(0.0) 0.86(0.0)
-GMM 54.56(1.918) 73.105(0.798) 0.306(0.015)
-Agglomerative 56.59(0.0) 70.461(0.0) 0.29(0.0)
-Spectral 58.453(0.837) 71.797(0.23) 0.28(0.002)
-e5-large
-k-LLMmeans
-   single 55.2(1.625) 72.0(0.515) 0.17(0.01)
-   multiple 55.6(2.161) 72.1(0.725) 0.166(0.012)
-   FS-single 54.3(2.395) 71.6(0.833) 0.173(0.012)
-   FS-multiple 55.8(1.995) 72.4(0.912) 0.166(0.01)
-k-means 52.885(1.264) 70.917(0.708) 0.175(0.01)
-k-medoids 32.112(0.0) 50.313(0.0) 0.437(0.0)
-GMM 50.595(2.28) 70.15(1.225) 0.183(0.007)
-Agglomerative 55.447(0.0) 70.315(0.0) 0.17(0.0)
-Spectral 56.594(0.585) 69.034(0.417) 0.152(0.003)
-S-BERT
-k-LLMmeans
-   single 52.2(2.754) 71.2(0.785) 0.334(0.022)
-   multiple 52.9(2.005) 71.5(0.675) 0.325(0.015)
-   FS-single 52.0(2.552) 71.4(0.507) 0.333(0.019)
-   FS-multiple 53.3(1.768) 71.9(0.466) 0.337(0.011)
-k-means 50.908(1.747) 70.708(0.296) 0.344(0.019)
-k-medoids 35.978(0.0) 54.587(0.0) 0.817(0.0)
-GMM 50.494(1.22) 70.559(0.488) 0.338(0.012)
-Agglomerative 52.186(0.0) 69.544(0.0) 0.313(0.0)
-Spectral 56.126(1.395) 70.227(0.463) 0.304(0.005)
-Table 10: Average accuracy (ACC) and normalized mutual information (NMI) for k-LLMmeans FS-multiple using various LLMs and e5-large embeddings on all benchmark datasets (standard deviations in parenthesis).
+Appendix B Supplementary results
+B.1 Main results with standard deviations
+In Tables 5, 6, 7 and 4 we report the same results as in Tables 1, 2, 3 and 8 including standard deviations.
+Table 5: Average ACC and NMI for four k-LLMmeans variants using GPT-4o, compared against traditional baselines, with text-embedding-3-small embeddings, evaluated on benchmark datasets.
+Standard deviations of ACC and NMI in parenthesis
 Dataset/Method BANK77 CLINC GoEmo Massive (D) Massive (I)
 ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI
+k-NLPmeans
+     TextRank-single 66.4 83.6 78.7 92.4 21.5 20.9 60.6 68.9 53.0 72.5
+(0.94) (0.47) (1.56) (0.41) (0.63) (0.29) (3.47) (1.27) (1.8) (0.71)
+     TextRank-multiple 66.8 83.9 79.6 92.8 21.6 21.0 60.5 69.3 53.7 72.8
+(0.94) (0.45) (1.51) (0.45) (1.29) (0.65) (3.66) (1.22) (1.65) (0.58)
+     Centroid-single 67.2 84.0 78.5 92.2 22.1 21.2 60.9 68.3 53.5 73.0
+(0.81) (0.39) (1.68) (0.37) (0.6) (0.41) (3.02) (0.91) (2.07) (1.13)
+     Centroid-multiple 67.5 84.0 79.0 92.4 21.4 21.1 60.8 69.1 54.3 73.5
+(0.57) (0.3) (1.95) (0.46) (0.87) (0.12) (2.72) (1.09) (1.72) (0.84)
+     LSA-single 66.7 83.7 78.8 92.5 21.9 20.5 61.2 68.7 54.1 73.0
+(0.99) (0.47) (2.07) (0.63) (1.11) (0.87) (2.66) (1.03) (1.86) (0.91)
+     LSA-multiple 67.1 84.0 80.2 92.9 22.3 20.2 63.3 70.0 55.3 73.4
+(0.98) (0.42) (0.98) (0.41) (0.85) (0.56) (3.06) (1.02) (1.32) (0.72)
 k-LLMmeans
-   GPT-3.5 56.6 77.6 76.0 92.1 23.9 24.1 63.3 66.2 56.9 72.5
+     single 67.1 83.6 78.1 92.5 24.0 22.3 61.1 69.6 54.0 73.0
 (1.22) (0.46) (1.67) (0.47) (1.35) (0.57) (3.39) (1.6) (1.51) (0.58)
-   GPT-4o 62.7 79.5 77.2 92.5 24.2 24.3 62.3 65.9 55.8 72.4
+     multiple 66.7 83.6 79.1 92.8 24.1 22.1 60.6 69.5 55.8 73.5
 (0.78) (0.2) (1.68) (0.48) (1.33) (0.65) (2.82) (1.66) (1.4) (0.76)
-   Llama-3.3 61.6 79.2 77.2 92.3 23.8 23.1 62.0 66.3 56.8 72.9
+     FS-single 67.5 83.8 79.2 92.8 23.0 21.9 60.8 69.4 55.3 73.6
 (1.07) (0.36) (1.73) (0.43) (0.94) (0.64) (3.21) (1.09) (1.76) (0.77)
-   DeepSeek-V3 63.1 79.9 69.7 90.8 22.8 22.7 62.6 66.3 57.6 72.0
+     FS-multiple 67.9 84.1 80.2 93.1 24.2 22.3 63.2 70.6 56.7 73.9
 (1.55) (0.41) (1.52) (0.32) (1.16) (0.67) (2.84) (1.38) (1.32) (0.7)
-   Claude-3.7 61.0 78.8 76.9 92.5 24.2 23.7 61.8 66.0 57.3 72.9
+k-means 66.2 83.0 77.3 92.0 20.7 20.5 59.4 67.9 52.9 72.4
 (1.82) (0.56) (1.83) (0.35) (1.0) (0.67) (4.03) (1.67) (1.07) (0.54)
-Table 11: Standard deviation of the accuracy (ACC) and Normalized Mutual Information (NMI) for four sequential mini-batch variants, k-means, mini-batch k-means, sequential mini-batch k-means on the yearly StackExchange data.
+k-medoids 41.7 69.5 49.3 77.7 15.7 15.9 37.6 38.8 35.9 52.4
+(0.0) (0.0) (0.0) (0.0) (0.0) (0.0) (0.0) (0.0) (0.0) (0.0)
+GMM 67.7 83.0 78.9 92.7 21.5 20.6 56.2 68.6 53.9 73.2
+(1.78) (0.54) (1.18) (0.22) (0.82) (0.47) (2.88) (1.01) (1.91) (0.80)
+Agglomerative 69.9 83.7 81.0 92.5 15.8 14.1 62.8 67.1 56.6 70.5
+(0.0) (0.0) (0.0) (0.0) (0.0) (0.0) (0.0) (0.0) (0.0) (0.0)
+Spectral 68.2 83.3 76.3 90.9 17.6 15.2 61.5 67.0 56.6 71.4
+(0.56) (0.38) (0.401) (0.14) (0.32) (0.32) (0.02) (0.02) (0.84) (0.23)
+Table 6: Average ACC, NMI, and distdist for k-means and k-LLMmeans FS-multiple, evaluated on three datasets using four different embedding models.
+Standard deviations of ACC, NMI and distdist in parenthesis.
+Dataset CLINC GoEmo Massive (D)
+embedding/Method ACC NMI distdist ACC NMI distdist ACC NMI distdist
+DistilBERT
+     k-means 53.9 77.2 0.34 17.8 18.3 0.364 44.4 45.1 0.309
+(0.96) (0.49) (0.007) (0.86) (0.67) (0.005) (1.53) (0.99) (0.021)
+     k-NLPmeans 52.9 77.3 0.353 18.5 18.2 0.365 44.7 45.3 0.311
+(0.25) (0.22) (0.003) (0.58) (0.38) (0.004) (2.44) (1.22) (0.02)
+     k-LLMmeans 55.3 78.7 0.343 18.2 18.8 0.351 46.2 46.4 0.295
+(1.1) (0.53) (0.009) (0.52) (0.73) (0.004) (1.35) (1.12) (0.015)
+text-embedding-3-small
+     k-means 77.3 92.0 0.2 20.7 20.5 0.287 59.4 67.9 0.246
+(1.83) (0.35) (0.014) (1.0) (0.67) (0.005) (4.03) (1.67) (0.015)
+     k-NLPmeans 80.2 92.9 0.173 22.3 20.2 0.29 63.3 70.0 0.227
+(0.98) (0.41) (0.006) (0.85) (0.56) (0.007) (3.06) (1.02) (0.018)
+     k-LLMmeans 80.2 93.1 0.179 24.2 22.3 0.278 63.2 70.6 0.227
+(1.52) (0.32) (0.015) (1.16) (0.67) (0.011) (2.84) (1.38) (0.022)
+e5-large
+     k-means 73.8 90.8 0.131 22.8 22.8 0.176 58.4 63.7 0.138
+(2.34) (0.74) (0.01) (0.92) (0.73) (0.003) (3.89) (1.75) (0.02)
+     k-NLPmeans 76.5 92.0 0.119 22.6 22.6 0.179 60.4 65.7 0.137
+(1.76) (0.58) (0.009) (0.7) (0.32) (0.003) (3.2) (1.79) (0.008)
+     k-LLMmeans 77.2 92.5 0.119 24.2 24.3 0.168 62.3 65.9 0.133
+(1.6) (0.4) (0.007) (0.76) (0.7) (0.003) (1.36) (0.78) (0.007)
+S-BERT
+     k-means 76.9 91.0 0.215 13.7 13.3 0.355 58.2 64.6 0.271
+(1.63) (0.5) (0.013) (0.62) (0.57) (0.005) (2.03) (1.09) (0.023)
+     k-NLPmeans 79.0 91.9 0.201 14.1 12.8 0.363 58.5 64.6 0.272
+(1.65) (0.31) (0.017) (0.3) (0.29) (0.003) (2.24) (1.29) (0.022)
+     k-LLMmeans 79.7 92.5 0.198 14.7 13.9 0.346 59.7 65.6 0.254
+(0.73) (0.25) (0.008) (0.55) (0.5) (0.004) (2.54) (0.79) (0.019)
+Table 7: Number of LLM calls (prompts), average ACC, and average NMI for k-LLMmeans FS-multiple using various LLMs and e5-large embeddings, compared against other LLM-based text clustering methods on three benchmark datasets.
+Standard deviations of ACC and NMI in parenthesis (not available for baseline LLM-based methods).
+Dataset/Method CLINC GoEmo Massive (D)
+prompts ACC NMI prompts ACC NMI prompts ACC NMI
+k-NLPmeans 0 76.5 92.0 0 22.6 22.6 0 60.4 65.7
+(0.02) (0.01) (0.01) (0.0) (0.03) (0.02)
+k-LLMmeans
+     GPT-3.5 750 76.0 92.1 135 23.9 24.1 90 63.3 66.2
+(1.16) (0.28) (0.65) (0.39) (1.47) (0.62)
+     GPT-4o 750 77.2 92.5 135 24.2 24.3 90 62.3 65.9
+(1.6) (0.4) (0.76) (0.7) (1.36) (0.78)
+     Llama-3.3 750 77.2 92.3 135 23.8 23.1 90 62.0 66.3
+(1.6) (0.44) (0.66) (0.62) (2.97) (1.56)
+     DeepSeek-V3 750 69.7 90.8 135 22.8 22.7 90 62.6 66.3
+(1.02) (0.29) (1.68) (0.88) (2.24) (0.71)
+     Claude-3.7 750 76.9 92.5 135 24.2 23.7 90 61.8 66.0
+(1.83) (0.52) (1.0) (0.65) (4.83) (1.99)
+ClusterLLM 1618 83.8 94.0 1618 26.8 23.9 1618 60.9 68.8
+IDAS 4650 81.4 92.4 3011 30.6 25.6 2992 53.5 63.9
+LLMEdgeRefine 1350 86.8 94.9 895 34.8 29.7 892 63.05 68.67
+Table 8: Average ACC, and average NMI for four sequential mini-batch variants, k-means, mini-batch k-means, sequential mini-batch k-means on the yearly StackExchange data.
+Standard deviations of ACC and NMI in parenthesis.
 Year/Method 2020 2021 2022 2023
@@
 ACC NMI ACC NMI ACC NMI ACC NMI
+mini-batch k-NLPmeans 68.0 79.5 67.9 78.5 69.0 78.9 71.6 78.8
+(3.09) (0.61) (2.5) (0.27) (4.15) (0.71) (2.74) (0.43)
 mini-batch k-LLMmeans
-   single (3.19) (0.57) (2.48) (0.42) (2.72) (0.86) (1.59) (0.46)
-   multiple (3.02) (0.65) (2.18) (0.46) (2.48) (0.87) (2.24) (0.81)
-   FS + single (2.86) (0.39) (2.15) (0.35) (2.68) (0.9) (1.68) (0.59)
-   FS + multiple (2.02) (0.48) (2.39) (0.7) (2.28) (0.91) (1.38) (0.45)
-k-means (3.79) (0.75) (2.8) (0.25) (4.03) (0.69) (3.79) (0.56)
-mini-batch k-means (2.09) (0.9) (2.12) (0.84) (4.02) (0.98) (2.76) (0.97)
-seq. mini-batch k-means (2.39) (0.57) (4.61) (1.54) (1.23) (0.78) (2.55) (0.9)
+     multiple 72.5 80.9 73.6 80.5 74.4 80.5 73.4 80.3
+(3.02) (0.65) (2.18) (0.46) (2.48) (0.87) (2.24) (0.81)
+     FS + multiple 75.4 81.6 73.5 80.2 72.8 80.1 72.7 80.1
+(2.02) (0.48) (2.39) (0.7) (2.28) (0.91) (1.38) (0.45)
+k-means 73.4 80.6 67.7 79.0 68.6 79.0 72.0 79.6
+(3.79) (0.75) (2.8) (0.25) (4.03) (0.69) (3.79) (0.56)
+mini-batch k-means 67.0 78.2 67.7 77.4 67.5 77.6 67.2 77.0
+(2.09) (0.9) (2.12) (0.84) (4.02) (0.98) (2.76) (0.97)
+seq. mini-batch k-means 67.0 76.6 66.7 75.2 65.6 75.6 65.8 74.8
+(2.39) (0.57) (4.61) (1.54) (1.23) (0.78) (2.55) (0.9)
+B.2 Additional experiments
+We evaluated the sensitivity of k-LLMmeans to the instruction prompt $I$ on BANK77 by testing five prompt variants using k-LLMmeans (FS-single) with text-embedding-3-small and GPT-4o.
+We also examined the effect of the parameter $q$ on the same dataset and embeddings.
+As shown in Table 9, performance remains remarkably stable across prompt choices and $q$ settings, indicating that the methods are insensitive to these hyperparameters.
+Table 9: Average ACC, and average NMI for multiple instruction prompts II for k-LLMmeans; and multiple values of $q$ for k-NLPmeans on BANK77. Standard deviations in parenthesis.
+ACC NMI
+k-LLMmeans (Prompt $I$)
+The following is a cluster of online banking queries.
+Write a text that summarizes the following cluster: 67.5 83.8
+(1.07) (0.36)
+The following is a cluster of texts.
+Write a text that summarizes the following cluster: 66.3 83.6
+(0.2) (0.3)
+Write a query that summarizes the following online banking queries: 67.4 83.8
+(0.7) (0.3)
+Write a text that summarizes the following texts: 66.9 83.7
+(0.6) (0.2)
+Summarize the following list of online banking queries: 67.0 83.7
+(0.4) (0.1)
+Summarize the following list of texts: 66.4 83.8
+(0.6) (0.1)
+k-NLPmeans ($q$)
+3 66.7 84.0
+(1.8) (0.8)
+5 66.4 83.8
+(0.5) (0.3)
+10 66.6 83.8
+(0.7) (0.4)
+15 66.4 83.7
+(0.6) (0.3)
+Appendix C Algorithms
+This section includes Algorithm 1 (k-NLPmeans / k-LLMmeans) and Algorithm 2 (mini batch version).
+input : $D=\{d_{1},\dots,d_{n}\},k,I,m,l,T$
+for $i\leftarrow 1$ to $n$ do
+ $\mathbf{x}_{i}=\text{Embedding}(d_{i})$;
+end for
+for $t\leftarrow 1$ to $T$ do
+ if $t=1$ then
+    // Initialize using k-means++
+    $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\texttt{k-means++}(\{d_{1},\dots,d_{n}\},k)$;
+   end if
+ else if $t\bmod l=0$ then
+    // Summarization step every $l$ iterations
+    for $j\leftarrow 1$ to $k$ do
+       $\boldsymbol{\mu}_{j}\leftarrow\text{Embedding}\left(\text{j}^{\text{th}}\text{ cluster summary}\right)$;
+      end for
+   end if
+ else
+    // k-means step
+    for $j\leftarrow 1$ to $k$ do
+       $\boldsymbol{\mu}_{j}\leftarrow\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}$;
+      end for
+   end if
+ for $j\leftarrow 1$ to $k$ do
+    $C_{j}=\{\}$;
+   end for
+ for $i\leftarrow 1$ to $n$ do
+    $j^{*}\leftarrow\arg\min_{j\in\{1,\dots,k\}}d(x_{i},\boldsymbol{\mu}_{j})$;
+    // Assign $x_{i}$ to cluster $C_{j^{*}}$
+    $C_{j^{*}}\leftarrow C_{j^{*}}\cup\{x_{i}\}$;
+   end for
+end for
+return $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\},\{s_{1},\dots,s_{k}\}$
+Algorithm 1 k-NLPmeans / k-LLMmeans
+input : $\{D_{1},\cdots,D_{b}\},k,I,m,l,T\$// $b$ batches of documents
+for $j\leftarrow 1$ to $k$ do
+ $C_{j}=\{\}$;
+end for
+$\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\{{\bf 0},\dots,{\bf 0}\}$;
+for $i\leftarrow 1$ to $b$ do
+// Compute k-NLPmeans / k-LLMmeans with documents in batch11 1Here k-NLPMmeans/k-LLMmeans is initialized with the final centroids of the previous batch
+ $\{\boldsymbol{\mu}^{*}_{1},\dots,\boldsymbol{\mu}^{*}_{k}\},\{C^{*}_{1},\dots,C^{*}_{k}\},S_{b}\leftarrow\texttt{k-NLPmeans}(D_{i},k,I,m,l,T)\text{ or }\texttt{k-LLMmeans}(D_{i},k,I,m,l,T)$;
+ // Update centroids proportional to cluster and batch sizes
+ for $j\leftarrow 1$ to $k$ do
+    $\eta\leftarrow\frac{|C^{*}_{j}|}{|C_{j}|+|C^{*}_{j}|}$;
+    $\boldsymbol{\mu}_{j}\leftarrow\boldsymbol{\mu}_{j}(1-\eta)+\eta\boldsymbol{\mu}^{*}_{j}$;
+   end for
+end for
+return $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\},\{S_{1},\dots,S_{b}\}$
+Algorithm 2 Mini-batch k-NLPmeans / k-LLMmeans

### v4

@@
 Figure 1 illustrates our proposal and shows how even a single summarization step can redirect k-means toward a qualitatively improved solution.
+Our goal is not to strictly dominate every clustering algorithm, but to provide a simple novel approach that (i) improves clustering quality over standard k-means–style methods, (ii) yields human-readable centroids, and (iii) scales to large and streaming datasets with limited LLM usage.
 Figure 1: Illustration of k-NLPmeans/k-LLMmeans with a single summarization step.
 First panel shows the text embeddings with stars marking the initial centroids; second shows the partition reached after k-means iterations (a local minimum); third performs the summarization step, each previous cluster is summarized into a textual prototype and re-embedded; final panel runs one more k-means iteration using these summaries as centroids, yielding a qualitatively improved partition.
+This example is illustrative; in practice, clusters can be multi-topic, and summaries may describe multiple prominent subtopics for a single cluster.
 Summarization step.
-We consider two families of summarizers: (1) Classical NLP—centroid-based summarization [48], graph-based methods such as TextRank [37], and LSA-style techniques [14]—which are fast and deterministic (yielding k-NLPmeans); and (2) LLM-based summaries that can capture richer context and filter secondary or noisy content [71, 50, 28] (yielding k-LLMmeans).
+We consider two families of summarizers: (1) Classical NLP such as centroid-based summarization [48], graph-based methods such as TextRank [37], and LSA-style techniques [14], which are fast and deterministic (yielding k-NLPmeans); and (2) LLM-based summaries that can capture richer context and filter secondary or noisy content [71, 50, 28] (yielding k-LLMmeans).
 In both variants, summaries are recomputed only at scheduled iterations, so the algorithms share the same summary-as-centroid mechanism and differ solely in the summarizer.
@@
 Contributions.
-In summary, we (i) propose k-NLPmeans and k-LLMmeans, a summary-as-centroid variant of k-means that periodically replaces numeric centroids with textual prototypes re-embedded by the same encoder—an LLM-optional design that, by construction, preserves the standard k-means objective between summary steps; interpretability follows from human-readable centroids and transparent intermediate outputs; (ii) extend the approach to mini-batch k-means for efficient, online clustering of streams; (iii) present a comprehensive empirical study across datasets, embeddings, and summarizers, showing consistent gains over classical baselines and competitiveness with recent LLM-based clustering under a dataset-size–independent LLM budget; and (iv) provide a case study on sequential text streams and release a StackExchange-derived benchmark for evaluating streaming text clustering.
+In summary, we (i) propose k-NLPmeans and k-LLMmeans, a summary-as-centroid variant of k-means that periodically replaces numeric centroids with textual prototypes re-embedded by the same encoder, an LLM-optional design that, by construction, preserves the standard k-means objective between summary steps; interpretability follows from human-readable centroids and transparent intermediate outputs; (ii) extend the approach to mini-batch k-means for efficient, online clustering of streams; (iii) present a comprehensive empirical study across datasets, embeddings, and summarizers, showing consistent gains over classical baselines and competitiveness with recent LLM-based clustering under a dataset-size–independent LLM budget; and (iv) provide a case study on sequential text streams and release a StackExchange-derived benchmark for evaluating streaming text clustering.
 2 Preliminaries: k-means for text clustering
@@
 $\boldsymbol{\mu}_{j}=\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}.$ (2)
-Lloyd’s algorithm [35]—the standard heuristic for k-means—alternates between assigning each document embedding $\mathbf{x}_{i}$ to its nearest centroid and recomputing each centroid as the mean of its assigned points.
+Lloyd’s algorithm [35], the standard heuristic for k-means, alternates between assigning each document embedding $\mathbf{x}_{i}$ to its nearest centroid and recomputing each centroid as the mean of its assigned points.
 Repeating these two steps for $T$ iterations monotonically reduces the within‑cluster sum of squared distances, steering the procedure toward a (locally) optimal set of centroids.
@@
 We enhance k-means for text clustering by periodically replacing the numerical centroid update with summarization steps that yield textual-summary–based centroids.
-The procedure (Algorithm 1 in Appendix C) is identical to k-means algorithm except that every $l$ iterations the mean update in (2) is replaced by the embedding of a cluster textual summary.
+The procedure (Algorithm 1 in Appendix D) is identical to k-means algorithm except that every $l$ iterations the mean update in (2) is replaced by the embedding of a cluster textual summary.
 At all other iterations, the standard update is used (see Figure 1 for an illustration with one summarization step).
-The summarizer can be instantiated with classical, deterministic methods (k-NLPmeans) or with LLM-based summaries (k-LLMmeans), and is not restricted to the specific techniques we discuss—any textual summarization operator can be dropped in.
+The summarizer can be instantiated with classical, deterministic methods (k-NLPmeans) or with LLM-based summaries (k-LLMmeans), and is not restricted to the specific techniques we discuss, any textual summarization operator can be dropped in.
 3.1 k-NLPmeans.
@@
 Instead of relying solely on Euclidean means, a textual summary captures contextual cues present in the underlying documents; re-embedding this summary yields centroids that better reflect cluster semantics.
-In practice, summary-based centroids produce more interpretable and often more semantically coherent—partitions, even when k-means++ seeding is suboptimal.
+In practice, summary-based centroids produce more interpretable and often more semantically coherent partitions, even when k-means++ seeding is suboptimal.
 Apart from the summary steps, the procedure follows standard k-means: assignments and numeric updates are unchanged and the usual objective is preserved between summary iterations.
@@
 Mini-batch k-means [56] is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset.
-This approach substantially reduces memory usage and computational cost, making it well suited for continuously generated text streams—such as those from social media, news, or customer feedback—where data must be clustered incrementally without full dataset access.
+This approach substantially reduces memory usage and computational cost, making it well suited for continuously generated text streams, such as those from social media, news, or customer feedback; where data must be clustered incrementally without full dataset access.
 Mini-batch k-means exhibits convergence properties comparable to standard k-means while offering superior scalability.
@@
 We use ground-truth number of clusters and ten different seeds.
+We also include BERTopic [23] as a strong topic-modeling baseline.
+In addition, we report results obtained by applying our proposed methods to the document embeddings produced by BERTopic.
 For advanced LLM-based methods, we compare with results obtained by Feng et al.
@@
 All k-means based methods are initialized with k-means++.
+We evaluate performance using clustering accuracy (ACC), which measures the best one-to-one alignment between predicted clusters and gold labels, and Normalized Mutual Information (NMI), which quantifies the mutual information between them normalized to $\left[0,1\right]$ (See Appendix C for details on the metrics).
 We now provide a summary of the results.
 5.1 Results with static data
-Table 1: Average ACC and NMI for four k-LLMmeans variants using GPT-4o, compared against traditional baselines, with text-embedding-3-small embeddings, evaluated on benchmark datasets.
+Table 1: Average ACC and NMI for k-NLPmeans and k-LLMmeans variants using GPT-4o, compared against traditional baselines, BERTopic, and our k-NLPmeans LSA-multiple and k-LLMmeans FS-multiple variants applied to BERTopic embeddings, using text-embedding-3-small embeddings on benchmark datasets.
 Dataset/Method BANK77 CLINC GoEmo Massive (D) Massive (I)
@@
      FS-multiple 67.9 84.1 80.2 93.1 24.2 22.3 63.2 70.6 56.7 73.9
+BERTopic 71.2 83.0 79.0 91.6 18.2 16.6 36.1 61.6 54.8 70.6
+BERTopic+kNLPmeans 76.0 87.3 82.1 93.0 17.4 17.2 65.0 70.2 56.4 74.3
+BERTopic+kLLMmeans 75.7 87.4 82.1 93.0 17.7 17.2 63.9 69.8 56.4 74.2
 k-means 66.2 83.0 77.3 92.0 20.7 20.5 59.4 67.9 52.9 72.4
@@
 Spectral 68.2 83.3 76.3 90.9 17.6 15.2 61.5 67.0 56.6 71.4
-Table 2: Average ACC, NMI, and distdist for k-means and k-LLMmeans FS-multiple, evaluated on three datasets using four different embedding models.
+Table 2: Average ACC, NMI, and distdist for k-means, k-NLPmeans LSA-multiple and k-LLMmeans FS-multiple, evaluated on three datasets using four different embedding models.
 Dataset CLINC GoEmo Massive (D)
@@
      k-LLMmeans 79.7 92.5 0.198 14.7 13.9 0.346 59.7 65.6 0.254
-Table 3: Number of LLM calls (prompts), average ACC, and average NMI for k-LLMmeans FS-multiple using various LLMs and e5-large embeddings, compared against other LLM-based text clustering methods on three benchmark datasets.
+Table 3: Number of LLM calls (prompts), average ACC, and average NMI for k-NLPmeans (LSA-multiple) and k-LLMmeans (FS-multiple) using various LLMs with e5-large embeddings, compared against BERTopic, our variants applied to BERTopic embeddings, and other state-of-the-art LLM-based clustering methods on three benchmark datasets.
 Dataset/Method CLINC GoEmo Massive (D)
@@
      Claude-3.7 750 76.9 92.5 135 24.2 23.7 90 61.8 66.0
+BERTopic 0 77.8 90.8 0 20.3 20.8 0 38.5 61.7
+BERTopic+kNLPmeans 0 81.5 93.0 0 23.2 22.6 0 60.8 66.1
+BERTopic+kLLMmeans 750 81.4 93.0 135 24.0 23.0 90 58.6 65.6
 ClusterLLM 1618 83.8 94.0 1618 26.8 23.9 1618 60.9 68.8
@@
 Overall, both k-NLPmeans and k-LLMmeans improve upon traditional algorithms, with k-LLMmeans delivering the best results, and remaining efficient, as few-shot summarization appears sufficient without needing extensive prompt length.
+We also compare with BERTopic and apply our methods to its document embeddings.
+While BERTopic is strong on its own, adding our variants consistently improves its performance.
+Across all datasets, the best results come from either our stand-alone methods or their BERTopic-based versions.
 Comparing our approaches with k-means using different embeddings.
@@
 Table 3 reports the number of LLM calls (prompts) and mean ACC/NMI for k-NLPmeans (LSA-multiple) and k-LLMmeans (FS-multiple) on three benchmarks with e5-large embeddings.
-For k-LLMmeans we vary the backbone generator—including GPT-3.5—and observe stable performance across LLMs, indicating robustness to the specific model.
+For k-LLMmeans we vary the LLM generator, including GPT-3.5, and observe stable performance across LLMs, indicating robustness to the specific model.
 By design, k-NLPmeans uses zero LLM calls.
@@
 k-NLPmeans trails k-LLMmeans slightly but remains competitive without any LLM usage.
+We also evaluate BERTopic and apply our variants to its embeddings.
+In this setting, our methods achieve results extremely close to the best LLM-based models while requiring none or only a small number of LLM calls.
 Overall, our framework offers a favorable quality–cost trade-off: k-NLPmeans for zero-inference-cost deployments, and k-LLMmeans for higher accuracy under a small, fixed summarization budget.
+Additional results and experiments.
 Results with standard deviations are reported in Tables 5, 6, and 7 of Appendix B.1.
-We also assess sensitivity to the instruction prompt II in k-LLMmeans and to parameter $q$ in k-NLPmeans.
-As shown in Table 9 (Appendix B.2), performance is generally stable across prompt choices and $q$ values.
+We also assess sensitivity to the selection of the number of clusters $k$ in Appendix B.2.1.
+Table 9 shows robustness when the number of clusters mildly differs from the ground truth clusters, consistently better than vanilla k-means.
+We also evaluate sensitivity to the instruction prompt $I$ in k-LLMmeans and to parameter $q$ in k-NLPmeans in Appendix B.2.2.
+As shown in Table 10, performance remains generally stable across prompt choices and $q$ values.
+Table 11 in Appendix B.2.3 also compares several sampling strategies for the few-shot (FS) version of k-LLMmeans, showing that our k-means++ strategy has consistent performance.
+Summary of static experimental results
+Across our experiments, our methods (i) improve over classical and topic-modeling baselines (Table 1); (ii) outperforms k-means across embeddings (Table 2); and (iii) approaches strong LLM-based clustering methods at much lower LLM cost (Table 3).
 6 Sequential (mini-batch) experiments
@@
 Table 4 reports average ACC and NMI on the yearly StackExchange corpus, comparing mini‑batch k-NLPmeans (LSA-multiple) and mini‑batch k-LLMmeans variants against standard k-means, mini‑batch k-means, and sequential mini‑batch k-means.
-Mini-batch k-LLMmeans variants achieve the highest ACC and NMI—consistently outperforming baselines, and even surpassing full-dataset k-means.
+Mini-batch k-LLMmeans variants achieve the highest ACC and NMI, consistently outperforming baselines, and even surpassing full-dataset k-means.
 Despite operating in a streaming regime, our approach clusters the entire corpus of 205,943 posts using no more than 3,850 LLM calls, demonstrating a favorable accuracy-per-LLM-token trade-off.
 The few-shot summarization design keeps prompts short, sidestepping context-window limits while preserving cluster quality.
-Mini-batch k-NLPmeans also improves over all mini-batch baselines, though it trails the k-LLMmeans variants—likely because its extractive summaries are more sensitive to the noisiness and heterogeneity of community posts than LLM-generated centroids.
+Mini-batch k-NLPmeans also improves over all mini-batch baselines, though it trails the k-LLMmeans variants, likely because its extractive summaries are more sensitive to the noisiness and heterogeneity of community posts than LLM-generated centroids.
 Overall, these results highlight that our interpretable, mini-batch formulation scales to long sequential streams with strong accuracy and limited LLM usage.
@@
 Image Model Optimization begins with robustness anxieties (adversaries, occlusion, scale) around mainstream detectors/segmenters, pivots mid-year to “how do I implement this well?” (transfer learning, augmentation, preprocessing), and ends with production hardening (class imbalance, specialized losses, edge efficiency) as strong pretrained models and libraries made setup routine and pushed bottlenecks to data and deployment [8, 32].
-AI Evolution and Challenges moves from historical “what is AI becoming?” to curiosity about large generative/multimodal systems—probably sparked by public releases like CLIP/DALL·E and their failure modes—and finishes with system-level synthesis (neuro-symbolic integration, RL control, compute efficiency, safety) as enterprise adoption and policy attention grow [49, 51, 40, 42, 19].
-Advanced NLP Techniques tracks the transformer boom: early questions focus on which models and preprocessing to use; mid-year attention shifts to measuring semantics beyond BLEU (e.g., BERTScore, BLEURT) as long-context and domain-shift issues surface; late-year concerns are product-driven—sequence limits, domain vocab, multilingual/cross-modal alignment—reflecting rapid uptake of pretrained encoders/decoders in industry [10, 72, 57, 68].
+AI Evolution and Challenges moves from historical “what is AI becoming?” to curiosity about large generative/multimodal systems, probably sparked by public releases like CLIP/DALL·E and their failure modes, and finishes with system-level synthesis (neuro-symbolic integration, RL control, compute efficiency, safety) as enterprise adoption and policy attention grow [49, 51, 40, 42, 19].
+Advanced NLP Techniques tracks the transformer boom: early questions focus on which models and preprocessing to use; mid-year attention shifts to measuring semantics beyond BLEU (e.g., BERTScore, BLEURT) as long-context and domain-shift issues surface; late-year concerns are product-driven, sequence limits, domain vocab, multilingual/cross-modal alignment; reflecting rapid uptake of pretrained encoders/decoders in industry [10, 72, 57, 68].
 Practically, these trajectories can power better post categorization, searchability, answer routing, and trend detection.
@@
 Density-based approaches like DBSCAN [18] and graph-based methods detect clusters of arbitrary shapes, while spectral clustering [41] leverages eigen-decomposition to uncover complex structures.
-Model-based techniques—including Gaussian mixture models [15] and recent neural network frameworks [74, 25, 69, 70, 67]—provide probabilistic clustering formulations.
+Model-based techniques; including Gaussian mixture models [15] and recent neural network frameworks [74, 25, 69, 70, 67]; provide probabilistic clustering formulations.
 Additionally, topic modeling methods, from probabilistic latent semantic analysis [24] to latent Dirichlet allocation [7], capture word co-occurrence patterns and latent topics.
@@
 Although simple instruction prompts are effective (Appendix B.2), prompt design can still influence outcomes.
-The few-shot variant assumes that a small set of representative samples captures cluster structure—practical in many settings but potentially limiting for heterogeneous clusters.
+The few-shot variant assumes that a small set of representative samples captures cluster structure, practical in many settings but potentially limiting for heterogeneous clusters.
 Regarding potential pretraining exposure, note that the LLM in our pipeline is used solely to summarize texts already assigned to each cluster; it does not inject external labels or unseen content, so any exposure does not confer an unfair advantage.
 As indicated by suboptimal results on the StackExchange dataset (Table 4), k-NLPmeans may benefit from applying text-cleaning preprocessing prior to use.
+Finally, we illustrate qualitative failure cases in Appendix E.
 Cost.
@@
 Reproducibility Statement
-We release code to reproduce all results in this paper at https://github.com/jairoadiazr/summaryCentroids (see the README for instructions).
+We release anonymized code to reproduce all results in this paper at https://anonymous.4open.science/r/summaryCentroids-877C/ (see the README for instructions).
 Appendix A Datasets
@@
 In Tables 5, 6, 7 and 4 we report the same results as in Tables 1, 2, 3 and 8 including standard deviations.
-Table 5: Average ACC and NMI for four k-LLMmeans variants using GPT-4o, compared against traditional baselines, with text-embedding-3-small embeddings, evaluated on benchmark datasets.
+Table 5: Average ACC and NMI for k-NLPmeans and k-LLMmeans variants using GPT-4o, compared against traditional baselines, BERTopic, and our k-NLPmeans LSA-multiple and k-LLMmeans FS-multiple variants applied to BERTopic embeddings, using text-embedding-3-small embeddings on benchmark datasets.
 Standard deviations of ACC and NMI in parenthesis
@@
 (1.55) (0.41) (1.52) (0.32) (1.16) (0.67) (2.84) (1.38) (1.32) (0.7)
+BERTopic 71.2 83.0 79.0 91.6 18.2 16.6 36.1 61.6 54.8 70.6
+(0.0) (0.0) (0.0) (0.0) (0.0) (0.0) (0.0) (0.0) (0.0) (0.0)
+BERTopic+kNLPmeans 76.0 87.3 82.1 93.0 17.4 17.2 65.0 70.2 56.4 74.3
+(0.76) (0.11) (0.84) (0.09) (0.7) (0.3) (2.39) (0.66) (1.86) (0.27)
+BERTopic+kLLMmeans 75.7 87.4 82.1 93.0 17.7 17.2 63.9 69.8 56.4 74.2
+(1.26) (0.17) (0.83) (0.2) (0.43) (0.3) (1.35) (0.63) (1.19) (0.12)
 k-means 66.2 83.0 77.3 92.0 20.7 20.5 59.4 67.9 52.9 72.4
@@
 (0.56) (0.38) (0.401) (0.14) (0.32) (0.32) (0.02) (0.02) (0.84) (0.23)
-Table 6: Average ACC, NMI, and distdist for k-means and k-LLMmeans FS-multiple, evaluated on three datasets using four different embedding models.
+Table 6: Average ACC, NMI, and distdist for k-means, k-NLPmeans LSA-multiple and k-LLMmeans FS-multiple, evaluated on three datasets using four different embedding models.
 Standard deviations of ACC, NMI and distdist in parenthesis.
@@
 (0.73) (0.25) (0.008) (0.55) (0.5) (0.004) (2.54) (0.79) (0.019)
-Table 7: Number of LLM calls (prompts), average ACC, and average NMI for k-LLMmeans FS-multiple using various LLMs and e5-large embeddings, compared against other LLM-based text clustering methods on three benchmark datasets.
+Table 7: Number of LLM calls (prompts), average ACC, and average NMI for k-NLPmeans (LSA-multiple) and k-LLMmeans (FS-multiple) using various LLMs with e5-large embeddings, compared against BERTopic, our variants applied to BERTopic embeddings, and other state-of-the-art LLM-based clustering methods on three benchmark datasets.
 Standard deviations of ACC and NMI in parenthesis (not available for baseline LLM-based methods).
@@
      Claude-3.7 750 76.9 92.5 135 24.2 23.7 90 61.8 66.0
+(1.83) (0.52) (1.0) (0.65) (4.83) (1.99)
+BERTopic 0 77.8 90.8 0 20.3 20.8 0 38.5 61.7
+(0.0) (0.0) (0.0) (0.0) (0.0) (0.0)
+BERTopic+kNLPmeans 0 81.5 93.0 0 23.2 22.6 0 60.8 66.1
+(0.78) (0.18) (0.62) (0.13) (1.42) (0.58)
+BERTopic+kLLMmeans 750 81.4 93.0 135 24.0 23.0 90 58.6 65.6
 (0.55) (0.11) (0.36) (0.21) (2.96) (0.98)
@@
 B.2 Additional experiments
+B.2.1 Sensitivity to number of clusters $k$
+We also examined the sensitivity to the parameter $k$ (the number of clusters) by assessing the effect when the number of clusters $k$ is set below the ground truth ($k{-}20\%$, $k{-}10\%$), at the ground truth ($k$), and above it ($k{+}10\%$, $k{+}20\%$).
+The corresponding results are reported in Table 9.
+As shown there, both k-NLPmeans and k-LLMmeans consistently outperform kmeans regardless of the estimation error in $k$, and the overall impact on ACC and NMI remains minor.
+B.2.2 Sensitivity to prompt $I$ and parameter $q$
 We evaluated the sensitivity of k-LLMmeans to the instruction prompt $I$ on BANK77 by testing five prompt variants using k-LLMmeans (FS-single) with text-embedding-3-small and GPT-4o.
@@
 As shown in Table 10, performance remains remarkably stable across prompt choices and $q$ settings, indicating that the methods are insensitive to these hyperparameters.
-Table 9: Average ACC, and average NMI for multiple instruction prompts II for k-LLMmeans; and multiple values of $q$ for k-NLPmeans on BANK77. Standard deviations in parenthesis.
+B.2.3 Sensitivity to number sampling strategy for the few-shot (FS) variant of k-LLMmeans
+We also examine the effect of alternative sampling procedures for the few shot variant.
+We propose in the manuscript k-means++ but here we also explore uniform random selection (random), choosing closest to the cluster centroid (centroid) and choosing the farthest ones (edge).
+The corresponding results are reported in Table 11.
+As shown there, our proposed k-means++ seems to have a consistent performance.
+Table 9: Average ACC and NMI when the number of clusters $k$ is set below the ground truth ($k{-}20\%$, $k{-}10\%$), at the ground truth ($k$), and above it ($k{+}10\%$, $k{+}20\%$), for k-means, k-NLPmeans (LSA-single), and k-LLMmeans (FS-single, GPT-4o) using text-embedding-3-small, evaluated on benchmark datasets.
+Standard deviations of ACC and NMI in parenthesis.
+Dataset/Method BANK77 CLINC GoEmo Massive (D) Massive (I)
+ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI
+k-means
+     kk-20% 61.1 81.6 70.4 91.1 21.4 19.7 61.4 66.8 53.0 71.7
+(0.84) (0.21) (1.17) (0.43) (1.25) (0.46) (2.62) (2.04) (0.95) (0.57)
+     kk-10% 63.5 82.3 73.4 91.6 20.6 19.9 60.7 67.2 53.2 72.3
+(2.25) (0.77) (1.33) (0.54) (1.03) (0.47) (3.38) (1.7) (1.77) (0.63)
+     $k$ 65.1 83.0 77.5 92.0 20.7 20.5 60.6 67.9 53.2 72.5
+(1.36) (0.62) (1.55) (0.41) (0.21) (0.73) (2.93) (1.22) (0.94) (0.45)
+     kk+10% 64.8 83.1 77.8 91.8 20.0 20.3 59.1 67.7 53.1 72.7
+(0.96) (0.49) (1.13) (0.15) (1.02) (0.79) (3.95) (1.59) (1.62) (0.78)
+     kk+20% 64.6 83.1 77.2 91.8 19.4 20.7 59.8 69.8 52.1 72.4
+(1.35) (0.54) (1.23) (0.25) (0.68) (0.69) (0.91) (0.38) (0.83) (0.64)
+k-NLPmeans
+     kk-20 62.1 82.2 71.3 91.4 21.9 19.4 62.1 67.4 53.7 72.4
+(0.42) (0.37) (1.23) (0.43) (1.08) (0.63) (2.03) (1.54) (2.02) (0.58)
+     kk-10% 65.5 83.1 74.3 91.9 21.2 19.5 61.3 67.7 53.4 72.8
+(1.61) (0.7) (1.25) (0.47) (1.22) (0.97) (2.94) (1.71) (0.92) (0.63)
+     $k$ 66.7 83.7 78.4 92.4 21.7 20.5 61.1 68.6 54.1 73.0
+(0.99) (0.47) (2.06) (0.66) (1.19) (0.97) (2.96) (1.11) (2.08) (1.01)
+     kk+10% 66.9 83.9 78.6 92.3 20.1 20.4 60.8 69.2 53.9 73.3
+(1.03) (0.4) (1.22) (0.23) (0.38) (0.28) (3.48) (1.02) (1.76) (0.86)
+     kk+20 66.7 83.9 78.0 92.3 20.3 20.6 60.2 70.2 52.3 73.1
+(1.44) (0.31) (1.3) (0.15) (0.98) (0.49) (1.43) (0.76) (0.76) (0.69)
+k-LLMmeans
+     kk-20% 62.4 82.5 71.4 91.5 24.2 20.8 60.9 67.2 55.3 72.9
+(0.89) (0.38) (1.24) (0.57) (1.92) (1.14) (2.56) (1.51) (1.18) (0.81)
+     kk-10% 65.3 83.2 74.8 92.0 23.2 21.0 60.4 68.2 54.3 73.0
+(1.19) (0.42) (1.61) (0.47) (0.11) (0.49) (4.04) (1.94) (1.99) (0.34)
+     $k$ 66.8 83.8 79.1 92.7 22.9 21.6 60.8 69.2 55.3 73.7
+(0.94) (0.27) (0.75) (0.29) (0.32) (0.59) (2.79) (1.14) (2.08) (0.97)
+     kk+10% 67.4 84.2 80.7 92.9 23.1 22.2 60.1 69.8 55.1 73.9
+(0.85) (0.31) (0.09) (0.42) (0.54) (0.52) (3.78) (1.2) (2.14) (0.96)
+     kk+20% 66.6 84.1 79.2 92.9 22.2 22.6 58.7 70.0 53.5 73.6
+(1.23) (0.36) (1.06) (0.27) (0.86) (0.41) (0.9) (0.5) (2.07) (1.39)
+Table 10: Average ACC, and average NMI for multiple instruction prompts II for k-LLMmeans FS-multiple; and multiple values of $q$ for k-NLPmeans LSA-multiple on BANK77.
+Standard deviations in parenthesis.
 ACC NMI
@@
 (0.6) (0.3)
-Appendix C Algorithms
+Table 11: Average ACC and NMI for multiple sampling strategies for k-LLMmeans FS-single using GPT-4o with text-embedding-3-small embeddings, evaluated on benchmark datasets.
+Standard deviations of ACC and NMI in parenthesis.
+Dataset/Method BANK77 CLINC GoEmo Massive (D) Massive (I)
+ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI
+k-LLMmeans
+     kmeans++ 67.3 84.0 79.5 92.8 23.1 21.8 60.3 69.1 55.0 73.5
+(1.35) (0.41) (1.18) (0.34) (0.63) (0.67) (2.66) (1.03) (1.92) (0.95)
+     random 66.2 83.6 80.6 93.0 23.4 21.9 62.2 68.0 53.5 72.8
+(0.46) (0.31) (1.27) (0.3) (0.86) (0.51) (2.95) (1.16) (1.18) (0.64)
+     centroid 67.2 84.0 79.3 92.5 22.9 22.5 61.6 69.5 53.9 72.8
+(0.54) (0.31) (1.5) (0.34) (1.26) (0.53) (3.33) (1.7) (1.11) (0.28)
+     edge 66.4 83.3 78.1 92.5 22.6 21.7 63.1 70.2 53.6 72.7
+(0.58) (0.27) (1.62) (0.44) (1.33) (0.47) (2.08) (0.64) (0.99) (0.15)
+Appendix C Evaluation metrics
+For evaluation, we report clustering accuracy (ACC) and normalized mutual information (NMI).
+NMI is computed with normalized_mutual_info_score from sklearn.metrics, which implements the standard mutual-information–based normalization.
+Let $Y$ denote the ground-truth labels and $\hat{Y}$ the predicted cluster labels.
+We first form the contingency matrix $W\in\mathbb{N}^{D\times D}$ with entries $W_{ij}=|\{n:\hat{y}_{n}=i,\;y_{n}=j\}|$, where $D$ is the maximum label index plus one and $N$ is the number of data points.
+The empirical mutual information between $Y$ and $\hat{Y}$ is
+$\mathrm{MI}(Y,\hat{Y})\;=\;\sum_{i,j}\frac{W_{ij}}{N}\log\frac{N\,W_{ij}}{\left(\sum_{j^{\prime}}W_{ij^{\prime}}\right)\left(\sum_{i^{\prime}}W_{i^{\prime}j}\right)},$
+and the entropies are
+$H(Y)=-\sum_{j}\frac{\sum_{i}W_{ij}}{N}\log\frac{\sum_{i}W_{ij}}{N},\qquad H(\hat{Y})=-\sum_{i}\frac{\sum_{j}W_{ij}}{N}\log\frac{\sum_{j}W_{ij}}{N}.$
+Following the default in sklearn, NMI is then defined as
+$\mathrm{NMI}(Y,\hat{Y})\;=\;\frac{\mathrm{MI}(Y,\hat{Y})}{\tfrac{1}{2}\bigl(H(Y)+H(\hat{Y})\bigr)}$.
+ACC is defined as the fraction of correctly assigned points after the best one-to-one relabeling of clusters.
+Concretely, given the same contingency matrix $W$, we solve a maximum-weight bipartite matching using the Hungarian algorithm on $W$ and compute
+$\mathrm{ACC}\;=\;\frac{1}{N}\sum_{(i,j)\in\mathcal{M}}W_{ij},$
+where $\mathcal{M}$ is the set of matched label pairs returned by the Hungarian algorithm.
+Appendix D Algorithms
 This section includes Algorithm 1 (k-NLPmeans / k-LLMmeans) and Algorithm 2 (mini batch version).
@@
 Algorithm 2 Mini-batch k-NLPmeans / k-LLMmeans
+Appendix E Qualitative Analysis and Failure Modes
+In the main paper we focused on successful examples to illustrate how summary-as-centroid can make clusters more interpretable.
+For completeness, we provide qualitative failure cases that highlight typical limitations of our approach.
+The examples below are representative, not exhaustive, and are lightly anonymized / paraphrased for clarity.
+E.1 Prompt-Echo and Meta Summaries
+The first failure mode we mention (for k-LLMmeans) is “prompt echoing”, where the LLM partially repeats the instruction instead of summarizing the actual content.
+Example 1 (Prompt-echo / meta summary).
+Cluster snippets (paraphrased) Generated summary
+“My subscription was renewed without my consent.”
+“Please cancel my premium plan and refund the last payment.”
+“I was charged after I thought I had cancelled.”
+“A group of user support messages asking about the topics mentioned above, focusing on account and billing issues.”
+Meta phrases like “messages” and “topics mentioned above” make the summary less useful as a concise description.
+Slightly tightening the prompt (e.g., “Do not mention that you are summarizing, and avoid meta phrases like ‘questions’ or ‘messages”’) reduces this behavior.
+E.2 Spurious Detail and Hallucinated Constraints
+We also observe occasional cases where the summary introduces details that are not consistently supported by the cluster.
+Example 2 (Hallucinated detail).
+Cluster snippets (paraphrased) Generated summary
+“The app keeps crashing when I open the camera.”
+“It freezes on the loading screen after the last update.”
+“The app closes automatically when I try to log in.”
+“Complaints about the mobile app crashing on Android after the latest security update.”
+The summary correctly captures crashes but adds a misleading detail (“on Android”) that is not consistently supported.
+Mitigations include (i) asking the LLM to avoid unsupported specifics (“avoid making up details such as platforms or versions”), and (ii) using extractive summarization in high-stakes settings.
+E.3 Over-Compressed Multi-Topic Summaries
+Another failure mode appears when a cluster genuinely mixes several distinct topics of similar frequency.
+In this case, the summarizer sometimes tries to “cover everything at once”, producing an over-compressed summary that mentions multiple things but is too broad to be useful.
+Example 3 (Over-compressed multi-topic summary).
+Cluster snippets (paraphrased) Generated summary
+“How do I change my password?”
+“Where can I update my email address?”
+“How do I delete my account permanently?”
+“Can I change my username without losing data?”
+“Questions about managing and modifying user accounts, including changing settings and making updates.”
+Here, the summary technically mentions “managing and modifying user accounts” but compresses several distinct operations (password reset, email change, account deletion, username change) into a single vague description.
+This makes the prototype less informative for downstream users who might want to distinguish, for example, deletion vs.
+simple edits.
+In practice, we found that allowing slightly more detailed summaries (e.g., “changing passwords, emails, usernames, and deleting accounts”) or generating short bullet-style summaries can alleviate this issue, but it also highlights the inherent tension between brevity and specificity when clusters are truly multi-topic.
+E.4 Overly Generic Summaries
+This failure mode is opposite to the previous one.
+It happens when summaries are too generic and do not convey the specific intent or topic of the cluster.
+This often occurs when the cluster is relatively heterogeneous or when the summarizer is forced to be very short.
+Example 4 (Overly generic summary).
+Cluster snippets (paraphrased) Generated summary
+“I lost access to my card, can you freeze it?”
+“My card was stolen and I need a replacement.”
+“Please block my card, someone used it without permission.”
+“Questions about using the banking service.”
+The cluster is clearly about card freezing / blocking, but the summary collapses this into a vague description.
+The summary is still related to the domain but less helpful for human interpretation than it could be.
+Allowing slightly longer summaries and explicitly prompting for “key actions or problems” tends to reduce this failure mode.
+E.5 Minority-Topic Overshadowing
+A fifth failure mode appears when a cluster contains a dominant subtopic alongside a smaller but important minority subtopic.
+Because our approach uses a single prototype per cluster, the summary can under-represent the minority.
+Example 5 (Minority-topic overshadowing).
+Cluster snippets (paraphrased) Generated summary
+“How do I book a train ticket for tomorrow?”
+“Can I book a round trip in one payment?”
+“I need to cancel my ticket and get a refund.”
+“Questions about booking train tickets.”
+Here, the summary focuses on the majority “booking” pattern and omits cancellation/refund, which might be crucial for downstream labeling.
+This limitation is inherent to any single-prototype method (numeric or textual).
+Diversity-aware sampling within the cluster (for k-LLMmeans) and permitting a brief list of subtopics (e.g., “booking and cancelling train tickets”) can partially mitigate this issue.
+E.6 Mildly Misleading Emphasis
+A sixth failure mode arises when the summary is broadly correct but emphasizes a less central aspect of the cluster, which can be slightly misleading for a human reader.
+Example 6 (Mildly misleading emphasis).
+Cluster snippets (paraphrased) Generated summary
+“How can I change the delivery address for my order?”
+“I moved, can you update the shipping address?”
+“Can I edit the address before the package is sent?”
+“Questions about tracking online orders.”
+The summary remains in the e-commerce domain but shifts the focus from updating addresses to tracking.
+This type of error is usually easy to spot and correct by a human annotator, but it shows that summaries are not guaranteed to perfectly match the most salient aspect of a cluster.
+E.7 Cross-Lingual and Code-Mixed Drift
+Finally, on multilingual or code-mixed data (e.g., MASSIVE or StackExchange posts mixing English and another language), we occasionally see summaries drifting toward the dominant language or dropping non-English nuances.
+Example 7 (Cross-lingual drift).
+Cluster snippets (paraphrased) Generated summary
+“¿Puedo cambiar el idioma de la app a español?”
+“How do I switch the interface to French?”
+“Comment changer la langue par défaut de l’application ?”
+“Questions about using the app.”
+Here, the summary fails to capture the central concept of changing the app language and collapses to a very generic description.
+Using multilingual embeddings and explicitly mentioning in the prompt that the summary should “describe the main problem or request, even if the questions are in different languages” improves such cases but does not completely eliminate them.
+E.8 Discussion on failure modes.
+These qualitative cases illustrate that our textual prototypes are not infallible: they can be too generic, under-represent minority subtopics, echo the prompt, hallucinate details, or struggle with multilingual nuance.
+Nonetheless, our quantitative results show that summary-based centroids consistently outperform numeric centroids across datasets and embeddings, and provide an interpretable handle on cluster semantics.
+In the worst case, a poor summary behaves like a suboptimal centroid update and the procedure effectively falls back toward a vanilla k-means solution, so performance does not catastrophically degrade.
+We view developing more robust prompts, stronger multilingual summarization, and multi-prototype extensions (e.g., per-cluster sub-summaries) as promising directions for future work.

### v5

@@
 We introduce k-NLPmeans and k-LLMmeans, text-clustering variants of k-means that periodically replace numeric centroids with textual summaries.
-The key idea—summary-as-centroid—retains k-means assignments in embedding space while producing human-readable, auditable cluster prototypes.
+The key idea, summary-as-centroid, retains k-means assignments in embedding space while producing human-readable, auditable cluster prototypes.
 The method is LLM-optional: k-NLPmeans uses lightweight, deterministic summarizers, enabling offline, low-cost, and stable operation; k-LLMmeans is a drop-in upgrade that uses an LLM for summaries under a fixed per-iteration budget whose cost does not grow with dataset size.
 We also present a mini-batch extension for real-time clustering of streaming text.
-Across diverse datasets, embedding models, and summarization strategies, our approach consistently outperforms classical baselines and approaches the accuracy of recent LLM-based clustering—without extensive LLM calls.
+Across diverse datasets, embedding models, and summarization strategies, our approach consistently outperforms classical baselines and approaches the accuracy of recent LLM-based clustering without extensive LLM calls.
 Finally, we provide a case study on sequential text streams and release a StackExchange-derived benchmark for evaluating streaming text clustering.
 1 Introduction
-Text clustering is a core problem in natural language processing (NLP), with applications in document organization, topic exploration, and information retrieval [55, 61].
-A standard pipeline embeds documents into vectors [17, 54, 38, 46, 10, 29] and then groups them with a clustering algorithm [47].
-Among these algorithms, k-means [36] remains ubiquitous, iteratively updating each centroid as the mean of its assigned points.
-While effective, this purely numerical averaging can blur contextual nuance present in the original texts [52].
-Prior work has explored alternative centroid definitions and related objectives [27, 9, 31], yet these approaches remain anchored in vector space, which limits interpretability and can induce semantic drift between centroids and their underlying documents.
+Text clustering is a core problem in natural language processing (NLP), with applications in document organization, topic exploration, and information retrieval (Schütze et al., 2008; Steinbach, 2000).
+A standard pipeline embeds documents into vectors (Devlin, 2018; Sanh, 2019; Mikolov, 2013; Pennington et al., 2014; Brown et al., 2020; Jin et al., 2023) and then groups them with a clustering algorithm (Petukhova et al., 2025).
+Among these algorithms, k-means (MacQueen, 1967) remains ubiquitous, iteratively updating each centroid as the mean of its assigned points.
+While effective, this purely numerical averaging can blur contextual nuance present in the original texts (Reimers and Gurevych, 2019).
+Prior work has explored alternative centroid definitions and related objectives (Jain and Dubes, 1988; Bradley et al., 1996; Kaufman and Rousseeuw, 2008), yet these approaches remain anchored in vector space, which limits interpretability and can induce semantic drift between centroids and their underlying documents.
 This motivates centroids that are explicitly textual, aligning prototypes with human-interpretable summaries.
@@
 Summarization step.
-We consider two families of summarizers: (1) Classical NLP such as centroid-based summarization [48], graph-based methods such as TextRank [37], and LSA-style techniques [14], which are fast and deterministic (yielding k-NLPmeans); and (2) LLM-based summaries that can capture richer context and filter secondary or noisy content [71, 50, 28] (yielding k-LLMmeans).
+We consider two families of summarizers: (1) Classical NLP such as centroid-based summarization, graph-based methods such as TextRank , and LSA-style techniques (Radev et al., 2004; Mihalcea and Tarau, 2004; Deerwester et al., 1990), which are fast and deterministic (yielding k-NLPmeans); and (2) LLM-based summaries that can capture richer context and filter secondary or noisy content (Zhang et al., 2020a; Raffel et al., 2020; Jia and Diaz-Rodriguez, 2025) (yielding k-LLMmeans).
 In both variants, summaries are recomputed only at scheduled iterations, so the algorithms share the same summary-as-centroid mechanism and differ solely in the summarizer.
@@
 The summarizer is modular: it can be LLM-free (k-NLPmeans) or LLM-assisted under a fixed per-iteration budget that does not grow with dataset size (k-LLMmeans).
-For large-scale and streaming settings, we adapt mini-batch k-means [56] by inserting summarization steps into the mini-batch update rule, enabling efficient online clustering while retaining interpretability.
+For large-scale and streaming settings, we adapt mini-batch k-means (Sculley, 2010) by inserting summarization steps into the mini-batch update rule, enabling efficient online clustering while retaining interpretability.
 Relation to existing LLM-based clustering.
-Recent work shows that LLMs can deliver strong unsupervised clustering [73, 20, 13, 64, 58, 62, 39].
+Recent work shows that LLMs can deliver strong unsupervised clustering (Zhang et al., 2023; Feng et al., 2024; De Raedt et al., 2023; Viswanathan et al., 2024; Shi and Sakai, 2023; Tarekegn et al., 2024; Nakshatri et al., 2023).
 However, many pipelines face two practical issues: (i) scalability, as they may rely on semi-supervision, iterative labeling, or a number of LLM calls that grows with dataset size; and (ii) opaque optimization, since they often combine prompts, greedy merges, and similarity thresholds without an explicit objective, making convergence behavior hard to analyze.
@@
 The goal of k-means clustering is to partition these $n$ document embeddings into $k$ clusters, minimizing the intra-cluster variance. Formally, we define the clustering objective as:
-$\min_{C_{1},C_{2},\dots,C_{k}}\sum_{j=1}^{k}\sum_{i\in[C_{j}]}\|\mathbf{x}_{i}-\boldsymbol{\mu}_{j}\|^{2},$ (1)
-where $C_{j}$ denotes the set of embeddings assigned to cluster $j$, $C_{j}]=\{i|\mathbf{x}_{i}\in C_{j}\}$ denotes the set of embedding indices assigned to cluster $j$ and $\boldsymbol{\mu}_{j}$ is the cluster centroid, computed as the mean of the assigned embeddings:
-$\boldsymbol{\mu}_{j}=\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}.$ (2)
-Lloyd’s algorithm [35], the standard heuristic for k-means, alternates between assigning each document embedding $\mathbf{x}_{i}$ to its nearest centroid and recomputing each centroid as the mean of its assigned points.
+$\min_{C_{1},C_{2},\dots,C_{k}}\sum_{j=1}^{k}\sum_{i\in[C_{j}]}\|\mathbf{x}_{i}-\bm{\mu}_{j}\|^{2},$ (1)
+where $C_{j}$ denotes the set of embeddings assigned to cluster $j$, $C_{j}]=\{i|\mathbf{x}_{i}\in C_{j}\}$ denotes the set of embedding indices assigned to cluster $j$ and $\bm{\mu}_{j}$ is the cluster centroid, computed as the mean of the assigned embeddings:
+$\bm{\mu}_{j}=\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}.$ (2)
+Lloyd’s algorithm (Lloyd, 1982), the standard heuristic for k-means, alternates between assigning each document embedding $\mathbf{x}_{i}$ to its nearest centroid and recomputing each centroid as the mean of its assigned points.
 Repeating these two steps for $T$ iterations monotonically reduces the within‑cluster sum of squared distances, steering the procedure toward a (locally) optimal set of centroids.
-However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima [36, 35].
-Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results [5].
+However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima (MacQueen, 1967; Lloyd, 1982).
+Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results (Arthur and Vassilvitskii, 2006).
 3 k-means with summarization steps
 We enhance k-means for text clustering by periodically replacing the numerical centroid update with summarization steps that yield textual-summary–based centroids.
-The procedure (Algorithm 1 in Appendix D) is identical to k-means algorithm except that every $l$ iterations the mean update in (2) is replaced by the embedding of a cluster textual summary.
+The procedure (Algorithm 1 in Appendix D) is identical to k-means algorithm except that every $l$ iterations the mean update in equation 2 is replaced by the embedding of a cluster textual summary.
 At all other iterations, the standard update is used (see Figure 1 for an illustration with one summarization step).
@@
 3.1 k-NLPmeans.
-Our first variant uses classical extractive summarization to compute a textual prototype in place of the standard centroid update. Formally, we replace (2) with
-$\boldsymbol{\mu}_{j}=\text{Embedding}\!\left(f^{(q)}_{\text{NLP}}(S_{j})\right),$ (3)
+Our first variant uses classical extractive summarization to compute a textual prototype in place of the standard centroid update. Formally, we replace equation 2 with
+$\bm{\mu}_{j}=\text{Embedding}\!\left(f^{(q)}_{\text{NLP}}(S_{j})\right),$ (3)
 where $S_{j}$ is the collection (multiset) of sentences obtained by tokenizing all documents assigned to cluster $j$.
@@
 The operator $f^{(q)}_{\text{NLP}}$ returns a short summary of $q$ sentences; typical instantiations include:
-  • Centroid-based summarization [48]: compute the centroid of the sentence embeddings for $S_{j}$; rank sentences by cosine similarity to this centroid; concatenate the top $q$ sentences (by rank) to form $f^{(q)}_{\text{NLP}}(S_{j})$.
-  • Graph-based (TextRank) [37]: build a graph whose nodes are sentences in $S_{j}$ with edge weights given by pairwise cosine similarity of their embeddings; run a PageRank-style algorithm to score sentences; select and concatenate the top $q$.
-  • LSA-style SVD in embedding space [14]: stack the sentence embeddings for $S_{j}$, apply singular value decomposition, score sentences by contribution to the leading components, and concatenate the top $q$.
-The resulting summary text $f^{(q)}_{\text{NLP}}(S_{j})$ is then embedded in the same vector space as the documents to yield the new centroid $\boldsymbol{\mu}_{j}$.
+  • Centroid-based summarization (Radev et al., 2004): compute the centroid of the sentence embeddings for $S_{j}$; rank sentences by cosine similarity to this centroid; concatenate the top $q$ sentences (by rank) to form $f^{(q)}_{\text{NLP}}(S_{j})$.
+  • Graph-based (TextRank) (Mihalcea and Tarau, 2004): build a graph whose nodes are sentences in $S_{j}$ with edge weights given by pairwise cosine similarity of their embeddings; run a PageRank-style algorithm to score sentences; select and concatenate the top $q$.
+  • LSA-style SVD in embedding space (Deerwester et al., 1990): stack the sentence embeddings for $S_{j}$, apply singular value decomposition, score sentences by contribution to the leading components, and concatenate the top $q$.
+The resulting summary text $f^{(q)}_{\text{NLP}}(S_{j})$ is then embedded in the same vector space as the documents to yield the new centroid $\bm{\mu}_{j}$.
 3.2 k-LLMmeans.
-If instead we generate the summary with an LLM, we have k-LLMmeans. Formally, we replace (2) by
-$\boldsymbol{\mu}_{j}=\text{Embedding}(f_{\text{LLM}}(p_{j}))\\$ (4)
+If instead we generate the summary with an LLM, we have k-LLMmeans. Formally, we replace equation 2 by
+$\bm{\mu}_{j}=\text{Embedding}(f_{\text{LLM}}(p_{j}))\\$ (4)
 where $p_{j}=\text{Prompt}\left(I,\{d_{z_{i}}|z_{i}\sim[C_{j}]\}_{i=1}^{m_{j}}\right)$, and $m_{j}=\min(m,|C_{j}|)$.
-Here, $z_{i}\sim[C_{j}]$ denotes a sampled index of the embeddings assigned to cluster $C_{j}$ (without repetitions) and $m$ is a parameter that represents the maximum number of sampled indices used to compute the cluster centroid $\boldsymbol{\mu}_{j}$.
+Here, $z_{i}\sim[C_{j}]$ denotes a sampled index of the embeddings assigned to cluster $C_{j}$ (without repetitions) and $m$ is a parameter that represents the maximum number of sampled indices used to compute the cluster centroid $\bm{\mu}_{j}$.
 In simple terms, we update a cluster’s centroid by using the embedding of the response generated by an LLM when queried with a prompt containing a summarization instruction $I$ and a representative sample of documents from the cluster.
@@
 (2) Scalability.
-Unlike most state-of-the-art methods, whose LLM usage complexity grows with sample size [20, 13] or require fine-tuning [73].
+Unlike most state-of-the-art methods, whose LLM usage complexity grows with sample size (Feng et al., 2024; De Raedt et al., 2023) or require fine-tuning (Zhang et al., 2023).
 k-NLPmeans requires no LLM calls and k-LLMmeans only performs $k$ LLM calls per summarization step, with even few summarization steps yielding substantial performance gains.
@@
 4 Mini-batch k-NLPmeans and k-LLMmeans
-Mini-batch k-means [56] is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset.
+Mini-batch k-means (Sculley, 2010) is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset.
 This approach substantially reduces memory usage and computational cost, making it well suited for continuously generated text streams, such as those from social media, news, or customer feedback; where data must be clustered incrementally without full dataset access.
 Mini-batch k-means exhibits convergence properties comparable to standard k-means while offering superior scalability.
-Although numerous streaming clustering methods that do not rely on LLMs have been studied [59, 3, 53, 2, 1, 45], only a few incorporate LLMs [62, 39].
+Although numerous streaming clustering methods that do not rely on LLMs have been studied (Silva et al., 2013; Aggarwal, 2018; Ribeiro et al., 2017; Aggarwal et al., 2003; Ackermann et al., 2012; Ordonez, 2003), only a few incorporate LLMs (Tarekegn et al., 2024; Nakshatri et al., 2023).
 Moreover, existing offline LLM-based clustering approaches face scalability issues, highlighting the need for scalable summary-based clustering in an online setting.
@@
 5 Static experiments
-We use four benchmark datasets that span diverse domains and classification granularities: Bank77 [12], CLINC [33], GoEmo [16] and MASSIVE (domain and intent) [21].
+We use four benchmark datasets that span diverse domains and classification granularities: Bank77 (Casanueva et al., 2020), CLINC (Larson et al., 2019), GoEmo (Demszky et al., 2020) and MASSIVE (domain and intent) (FitzGerald et al., 2023).
 See Appendix A.1 for detailed descriptions of the datasets.
@@
 The single variant uses a single summarization step ($l=60$), while the multiple variant performs five summarization steps ($l=20$).
-To demonstrate the robustness of our approach, we compute embeddings with models: DistilBERT[54], e5-large[65], S-BERT[52], and text-embedding-3-small[44].
+To demonstrate the robustness of our approach, we compute embeddings with models: DistilBERT(Sanh, 2019), e5-large(Wang et al., 2022), S-BERT(Reimers and Gurevych, 2019), and text-embedding-3-small(OpenAI, 2024).
 For k-NLPmeans we evaluate using TextRank, Centroid and LSA summarization methods mentioned in Section 3.1, with $q=5$.
-For the LLM component of k-LLMmeans, we use GPT-3.5-turbo[43], GPT-4o[26], Llama-3.3[22], Claude-3.7[4] and DeepSeek-V3[34].
+For the LLM component of k-LLMmeans, we use GPT-3.5-turbo(OpenAI, 2023), GPT-4o(Hurst et al., 2024), Llama-3.3(Grattafiori et al., 2024), Claude-3.7(Anthropic, 2025) and DeepSeek-V3(Liu et al., 2024).
 For the instruction task $I$, we employ a simple summarization prompt depending on the task.
@@
 For each setting we run five different seeds.
-As traditional clustering baselines, we consider k-means, k-medoids [31], spectral clustering [41], agglomerative clustering [30], and Gaussian Mixture Models (GMM) [15].
+As traditional clustering baselines, we consider k-means, k-medoids (Kaufman and Rousseeuw, 2008), spectral clustering (Ng et al., 2001), agglomerative clustering (Johnson, 1967), and Gaussian Mixture Models (GMM) (Dempster et al., 1977).
 We use ground-truth number of clusters and ten different seeds.
-We also include BERTopic [23] as a strong topic-modeling baseline.
+We also include BERTopic (Grootendorst, 2022) as a strong topic-modeling baseline.
 In addition, we report results obtained by applying our proposed methods to the document embeddings produced by BERTopic.
 For advanced LLM-based methods, we compare with results obtained by Feng et al.
-[20] on ClusterLLM [73], IDAS [13], and LLMEdgeRefine [20].
+(2024) on ClusterLLM (Zhang et al., 2023), IDAS (De Raedt et al., 2023), and LLMEdgeRefine (Feng et al., 2024).
 All k-means based methods are initialized with k-means++.
@@
 For context, we also include results from Feng et al.
-[20]: ClusterLLM (fine-tuned) and IDAS use GPT-3.5 with the same e5-large embeddings, while LLMEdgeRefine uses the stronger Instructor embeddings [29].
+(2024): ClusterLLM (fine-tuned) and IDAS use GPT-3.5 with the same e5-large embeddings, while LLMEdgeRefine uses the stronger Instructor embeddings (Jin et al., 2023).
 Our GPT-3.5 configuration of k-LLMmeans achieves comparable, sometimes slightly lower ACC/NMI, but with much fewer LLM calls (independent of dataset size) and no fine-tuning.
@@
 6 Sequential (mini-batch) experiments
-We use yearly posts from 2020 to 2023 from 35 StackExchange sites [60].
+We use yearly posts from 2020 to 2023 from 35 StackExchange sites (StackExchange, 2024).
 Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods.
@@
 7 Case study
-To demonstrate the interpretability of our method in capturing the evolution of clusters within sequential data, we present a case study using posts from the AI site in the 2021 Stack Exchange dataset [60].
+To demonstrate the interpretability of our method in capturing the evolution of clusters within sequential data, we present a case study using posts from the AI site in the 2021 Stack Exchange dataset (StackExchange, 2024).
 We apply our mini-batch k-LLMmeans algorithm with three equal-length batches and a total of ten clusters.
@@
 Across 2021, these clusters drift from fundamentals toward integration-at-scale, mirroring how tools, models, and deployment pressures matured.
-Image Model Optimization begins with robustness anxieties (adversaries, occlusion, scale) around mainstream detectors/segmenters, pivots mid-year to “how do I implement this well?” (transfer learning, augmentation, preprocessing), and ends with production hardening (class imbalance, specialized losses, edge efficiency) as strong pretrained models and libraries made setup routine and pushed bottlenecks to data and deployment [8, 32].
-AI Evolution and Challenges moves from historical “what is AI becoming?” to curiosity about large generative/multimodal systems, probably sparked by public releases like CLIP/DALL·E and their failure modes, and finishes with system-level synthesis (neuro-symbolic integration, RL control, compute efficiency, safety) as enterprise adoption and policy attention grow [49, 51, 40, 42, 19].
-Advanced NLP Techniques tracks the transformer boom: early questions focus on which models and preprocessing to use; mid-year attention shifts to measuring semantics beyond BLEU (e.g., BERTScore, BLEURT) as long-context and domain-shift issues surface; late-year concerns are product-driven, sequence limits, domain vocab, multilingual/cross-modal alignment; reflecting rapid uptake of pretrained encoders/decoders in industry [10, 72, 57, 68].
+Image Model Optimization begins with robustness anxieties (adversaries, occlusion, scale) around mainstream detectors/segmenters, pivots mid-year to “how do I implement this well?” (transfer learning, augmentation, preprocessing), and ends with production hardening (class imbalance, specialized losses, edge efficiency) as strong pretrained models and libraries made setup routine and pushed bottlenecks to data and deployment (Bochkovskiy et al., 2020; Kolesnikov et al., 2021).
+AI Evolution and Challenges moves from historical “what is AI becoming?” to curiosity about large generative/multimodal systems, probably sparked by public releases like CLIP/DALL·E and their failure modes, and finishes with system-level synthesis (neuro-symbolic integration, RL control, compute efficiency, safety) as enterprise adoption and policy attention grow (Radford et al., 2021; Ramesh et al., 2021; Nayak, 2021; OpenAI, 2021; European Commission, 2021).
+Advanced NLP Techniques tracks the transformer boom: early questions focus on which models and preprocessing to use; mid-year attention shifts to measuring semantics beyond BLEU (e.g., BERTScore, BLEURT) as long-context and domain-shift issues surface; late-year concerns are product-driven, sequence limits, domain vocab, multilingual/cross-modal alignment; reflecting rapid uptake of pretrained encoders/decoders in industry (Brown et al., 2020; Zhang et al., 2020b; Sellam et al., 2020; Xue et al., 2021).
 Practically, these trajectories can power better post categorization, searchability, answer routing, and trend detection.
 More broadly, they showcase how our mini-batch k-LLMmeans exposes readable centroids enabling end-to-end interpretability for sequential text streams, letting practitioners track topic evolution, attribute causes, and make transparent, timely decisions in dynamic corpora.
-Figure 2: Sequential evolution of the LLM-generated centroids for three primary clusters during the three batches of the sequential mini-batch k-LLMmeans process applied to 2021 posts from the AI Stack Exchange site [60].
+Figure 2: Sequential evolution of the LLM-generated centroids for three primary clusters during the three batches of the sequential mini-batch k-LLMmeans process applied to 2021 posts from the AI Stack Exchange site (StackExchange, 2024).
 Main aspects are manually highlighted.
@@
 Traditional clustering.
-Hierarchical methods [30, 6] build tree-structured representations of nested document relationships.
-Density-based approaches like DBSCAN [18] and graph-based methods detect clusters of arbitrary shapes, while spectral clustering [41] leverages eigen-decomposition to uncover complex structures.
-Model-based techniques; including Gaussian mixture models [15] and recent neural network frameworks [74, 25, 69, 70, 67]; provide probabilistic clustering formulations.
-Additionally, topic modeling methods, from probabilistic latent semantic analysis [24] to latent Dirichlet allocation [7], capture word co-occurrence patterns and latent topics.
-BERTopic [23] offers an embedding–reduction–density pipeline for interpretable topics, while DeepCluster [11] uses clustering to iteratively train encoders in a self-supervised manner (orthogonal to our fixed-embedding setting).
+Hierarchical methods (Johnson, 1967; Blashfield and Aldenderfer, 1978) build tree-structured representations of nested document relationships.
+Density-based approaches like DBSCAN (Ester et al., 1996) and graph-based methods detect clusters of arbitrary shapes, while spectral clustering (Ng et al., 2001) leverages eigen-decomposition to uncover complex structures.
+Model-based techniques; including Gaussian mixture models (Dempster et al., 1977) and recent neural network frameworks (Zhou et al., 2019; Huang et al., 2014; Yang et al., 2016; Zhang et al., 2021; Xie et al., 2016); provide probabilistic clustering formulations.
+Additionally, topic modeling methods, from probabilistic latent semantic analysis (Hofmann, 2001) to latent Dirichlet allocation (Blei et al., 2003), capture word co-occurrence patterns and latent topics.
+BERTopic (Grootendorst, 2022) offers an embedding–reduction–density pipeline for interpretable topics, while DeepCluster (Caron et al., 2018) uses clustering to iteratively train encoders in a self-supervised manner (orthogonal to our fixed-embedding setting).
+Jia and Diaz-Rodriguez (2026) also incorporate modern embeddings into text segmentation within classical change point detection methods.
 LLM-based clustering.
 Viswanathan et al.
-[64] employ LLMs to augment document representations, generate pseudo pairwise constraints, and post-correct low-confidence assignments for query-efficient, few-shot semi-supervised clustering.
+(2024) employ LLMs to augment document representations, generate pseudo pairwise constraints, and post-correct low-confidence assignments for query-efficient, few-shot semi-supervised clustering.
 Zhang et al.
-[73] propose ClusterLLM, which uses instruction-tuned LLMs via interactive triplet and pairwise feedback to cost-effectively refine clustering granularity.
-Complementary approaches [63, 47] show that context-derived representations capture subtle semantic nuances beyond traditional embeddings.
+(2023) propose ClusterLLM, which uses instruction-tuned LLMs via interactive triplet and pairwise feedback to cost-effectively refine clustering granularity.
+Complementary approaches (Tipirneni et al., 2024; Petukhova et al., 2025) show that context-derived representations capture subtle semantic nuances beyond traditional embeddings.
 Wang et al.
-[66] introduce a goal-driven, explainable method that employs natural language descriptions to clarify cluster boundaries, while [13] present IDAS for intent discovery using abstractive summarization.
+(2023) introduce a goal-driven, explainable method that employs natural language descriptions to clarify cluster boundaries, while (De Raedt et al., 2023) present IDAS for intent discovery using abstractive summarization.
 Feng et al.
-[20] propose LLMEdgeRefine an iterative mechanism that forms super-points to mitigate outliers and reassign ambiguous edge points, resulting in clusters with higher coherence and robustness.
+(2024) propose LLMEdgeRefine an iterative mechanism that forms super-points to mitigate outliers and reassign ambiguous edge points, resulting in clusters with higher coherence and robustness.
 9 Discussion
@@
 Reproducibility Statement
-We release anonymized code to reproduce all results in this paper at https://anonymous.4open.science/r/summaryCentroids-877C/ (see the README for instructions).
+We release data and code to reproduce all results in this paper at https://github.com/jairoadiazr/summaryCentroids (see the README for instructions).
+Acknowledgments
+This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) under grant DGECR-2022-04531.
+The author thanks Mumin Jia, the anonymous reviewers, and the session chair for their valuable feedback and insightful suggestions, which greatly improved the quality of this work.
 Appendix A Datasets
@@
 We evaluate our clustering approach on four benchmark datasets.
-Bank77 [12]: Consists of 3,080 customer queries related to banking services, categorized into 77 distinct intents.
-CLINC [33]: A diverse set of 4,500 queries spanning 150 intent classes across multiple domains, designed for open-domain intent classification.
-GoEmo [16]: Contains 2,984 social media posts annotated with 27 fine-grained emotion categories.
+Bank77 (Casanueva et al., 2020): Consists of 3,080 customer queries related to banking services, categorized into 77 distinct intents.
+CLINC (Larson et al., 2019): A diverse set of 4,500 queries spanning 150 intent classes across multiple domains, designed for open-domain intent classification.
+GoEmo (Demszky et al., 2020): Contains 2,984 social media posts annotated with 27 fine-grained emotion categories.
 We removed the neutral expressions to address data imbalance and retained only entries with a single, unique emotion.
-MASSIVE [21]: Comprises 2,974 English-language virtual assistant utterances grouped into 18 domains and 59 intent categories.
+MASSIVE (FitzGerald et al., 2023): Comprises 2,974 English-language virtual assistant utterances grouped into 18 domains and 59 intent categories.
 These datasets provide a robust evaluation setting for text clustering across different domains and classification granularities.
 A.2 New compiled dataset for testing text-Streaming Clustering Algorithms
-We extract and unify a challenging data stream comprising unique archive posts collected from 84 Stack Exchange sites [60].
+We extract and unify a challenging data stream comprising unique archive posts collected from 84 Stack Exchange sites (StackExchange, 2024).
 Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods.
@@
     // Initialize using k-means++
-    $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\texttt{k-means++}(\{d_{1},\dots,d_{n}\},k)$;
+    $\{\bm{\mu}_{1},\dots,\bm{\mu}_{k}\}\leftarrow\texttt{k-means++}(\{d_{1},\dots,d_{n}\},k)$;
    end if
@@
     for $j\leftarrow 1$ to $k$ do
-       $\boldsymbol{\mu}_{j}\leftarrow\text{Embedding}\left(\text{j}^{\text{th}}\text{ cluster summary}\right)$;
+       $\bm{\mu}_{j}\leftarrow\text{Embedding}\left(\text{j}^{\text{th}}\text{ cluster summary}\right)$;
       end for
@@
     for $j\leftarrow 1$ to $k$ do
-       $\boldsymbol{\mu}_{j}\leftarrow\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}$;
+       $\bm{\mu}_{j}\leftarrow\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}$;
       end for
@@
  for $i\leftarrow 1$ to $n$ do
-    $j^{*}\leftarrow\arg\min_{j\in\{1,\dots,k\}}d(x_{i},\boldsymbol{\mu}_{j})$;
+    $j^{*}\leftarrow\arg\min_{j\in\{1,\dots,k\}}d(x_{i},\bm{\mu}_{j})$;
     // Assign $x_{i}$ to cluster $C_{j^{*}}$
@@
 end for
-return $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\},\{s_{1},\dots,s_{k}\}$
+return $\{\bm{\mu}_{1},\dots,\bm{\mu}_{k}\},\{s_{1},\dots,s_{k}\}$
 Algorithm 1 k-NLPmeans / k-LLMmeans
@@
 end for
-$\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\{{\bf 0},\dots,{\bf 0}\}$;
+$\{\bm{\mu}_{1},\dots,\bm{\mu}_{k}\}\leftarrow\{{\bf 0},\dots,{\bf 0}\}$;
 for $i\leftarrow 1$ to $b$ do
 // Compute k-NLPmeans / k-LLMmeans with documents in batch11 1Here k-NLPMmeans/k-LLMmeans is initialized with the final centroids of the previous batch
- $\{\boldsymbol{\mu}^{*}_{1},\dots,\boldsymbol{\mu}^{*}_{k}\},\{C^{*}_{1},\dots,C^{*}_{k}\},S_{b}\leftarrow\texttt{k-NLPmeans}(D_{i},k,I,m,l,T)\text{ or }\texttt{k-LLMmeans}(D_{i},k,I,m,l,T)$;
+ $\{\bm{\mu}^{*}_{1},\dots,\bm{\mu}^{*}_{k}\},\{C^{*}_{1},\dots,C^{*}_{k}\},S_{b}\leftarrow\texttt{k-NLPmeans}(D_{i},k,I,m,l,T)\text{ or }\texttt{k-LLMmeans}(D_{i},k,I,m,l,T)$;
  // Update centroids proportional to cluster and batch sizes
@@
     $\eta\leftarrow\frac{|C^{*}_{j}|}{|C_{j}|+|C^{*}_{j}|}$;
-    $\boldsymbol{\mu}_{j}\leftarrow\boldsymbol{\mu}_{j}(1-\eta)+\eta\boldsymbol{\mu}^{*}_{j}$;
+    $\bm{\mu}_{j}\leftarrow\bm{\mu}_{j}(1-\eta)+\eta\bm{\mu}^{*}_{j}$;
    end for
 end for
-return $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\},\{S_{1},\dots,S_{b}\}$
+return $\{\bm{\mu}_{1},\dots,\bm{\mu}_{k}\},\{S_{1},\dots,S_{b}\}$
 Algorithm 2 Mini-batch k-NLPmeans / k-LLMmeans